### v1

# FiDeLiS: Faithful Reasoning in Large Language Model for Knowledge Graph Question Answering

Abstract
While large language models (LLMs) have achieved significant success in various applications, they often struggle with hallucinations, especially in scenarios that require deep and responsible reasoning. These issues could be partially mitigated by integrating external knowledge graphs (KG) in LLM reasoning. However, the method of their incorporation is still largely unexplored. In this paper, we propose a retrieval-exploration interactive method, FiDeLiS to handle intermediate steps of reasoning grounded by KGs. Specifically, we propose Path-RAG module for retrieving useful intermediate knowledge from KG for LLM reasoning. We incorporate the logic and common-sense reasoning of LLMs and the topological connectivity of KGs into the knowledge retrieval process, which provides more accurate retrieving performance. Furthermore, we propose to leverage deductive reasoning capabilities of LLMs as a better criterion to automatically guide the reasoning process in a stepwise and generalizable manner. Deductive verification serves as a precise indicator for when to cease further reasoning, thus avoiding misleading the chains of reasoning and unnecessary computation. Extensive experiments show that our method, as a training-free method with lower computational cost and better generality outperforms the existing strong baselines in three benchmarks. The code of this paper will be released later.
1 Introduction
The emergence and application of large language models (LLMs) [1, 31, 27] have attracted widespread attention from researchers and the general public. It demonstrates remarkable reasoning capabilities, managing to solve complex reasoning problems through step-by-step thinking and planning [47, 49]. However, the reasoning of LLMs is not invariably reliable and may conflict with factual reality [32], which limits the application in areas requiring high reliability, such as healthcare [14] and science [41].
The knowledge graph (KG) stores high-quality domain-specific knowledge in structured triplets, such as Wikidata [43], YAGO [36], and NELL [2]. KGs offer structured, explicit, and updated factual knowledge, which provides a faithful knowledge source for reasoning. Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretive process transparent. Due to its reliability and interpretability, it is considered as a promising method to improve the reliability of LLM reasoning. Therefore, several attempts have been conducted to integrate KGs with language models [32, 29, 17, 38]. Among them, the knowledge graph question answering (KGQA) is a critical task to incorporate the knowledge of KG into reasoning models [12, 44, 55]. Concretely, KGQA faces the following main challenges:
(I) How to retrieve specific knowledge from KGs to help reasoning precisely? Existing solutions include direct retrieval [37, 21] and semantic parsing [39, 10]. Direct retrieval involves taking the questions as a query and the knowledge triplets in the KG as candidates, using either sparse or dense retrieval techniques to identify several candidates most relevant to the query. Semantic parsing, on the other hand, transforms the question into an executable structured query statement (e.g., SPARQL) and executes the query in KGs. However, in KGs, some triplet by itself lacks rich semantic information due to the difference of KG schema, such as some entities are represented as machine identifiers (MIDs), and direct retrieval make it difficult to understand how relevant it really is to the user query. This becomes especially tricky in multi-hop question answering, where the intermediate triplets seem not directly related to the query can actually be crucial links to finding the right answer. Furthermore, semantic parsing faces the issues of non-executable or incorrectly executed generated queries [55].
(II) How to make the reasoning model understand and utilize the retrieved structured knowledge in KGs? Existing solutions include (1) finetune LMs on KG structure to generate relation path grounded by KGs. These reasoning paths are then used for LLMs to conduct faithful reasoning and generate interpretable results [29, 55]; (2) leverage the reasoning model to iteratively retrieve and reason on the subgraphs from KG [38, 9, 21] (i.e., at each step, the LLMs retrieve neighbors of the current nodes and then decide to answer the question or continue the next search step.) However, for the first approach, the reasoning steps generated by language models offer few guarantees about feasibility (i.e., the reasoning step might not exist in the KG), especially when multiple consecutive steps are combined into a reasoning path. We conduct the error analysis regarding this issue in Section 4. The latter solution face the challenge regarding determining the optimal stopping point for the exploration process. For example, ToG [38] prompts the LLMs to evaluate the adequacy of current reasoning paths for answer generation. However, assessing the adequacy of reasoning paths is itself a complex task that demands a deep understanding of the domain. This lack of clarity can easily result in either premature stopping or excessive continuation, further complicating the decision-making process in LLMs.
To address these two challenges, we propose a retrieval-exploration interactive method, FiDeLiS. It is designed to iteratively retrieve and reason on the subgraphs from KG, simulating a dynamic searching process. Specifically, we propose a Path-RAG module for retrieving useful intermediate knowledge from KGs. Compared to vanilla retriever which easily returns outputs that are not useful for finding the correct answer based on relevance to the query, we incorporate LLMs to generate an exhaustive list of keywords from the query to maximize coverage and ensure that no potential reasoning paths will be overlooked. We also integrate the topological connectivity of KGs to enhance system accuracy by considering entity relationships in the knowledge retrieval process. Furthermore, we propose to leverage deductive reasoning capabilities of LLMs [26, 18] as a better criterion for automatically directing the reasoning process step by step. Deductive reasoning serves as a clear indicator for ceasing further reasoning when necessary, which helps prevent errors in the reasoning chain and reduce unnecessary computation. Our framework enables us to employ beam search, a technique tailored for intermediate reasoning steps within a controllable searching space. We can guide the searching process by ensuring the logical soundness of each reasoning step to prevent errors throughout the reasoning process. Overall, our main contributions include:
  • We propose a retrieval-exploration interactive method that is specifically tailored to handle intermediate steps of reasoning grounded by KGs.
  • We propose deductive verification as precise indicators for when to cease further reasoning, thus avoiding misleading the chains of reasoning and unnecessary computation.
  • We integrate the reasoning process and topological connectivity of KGs into the knowledge retrieval process, which aids in recalling useful intermediate knowledge for reasoning and maintains efficiency as KGs scale up.
  • Extensive experiments show that our method, as a training-free method with lower computational cost and better generality outperforms the existing strong baselines in three benchmarks.
2 Method
Overview. FiDeLiS is designed as a retrieval-exploration interactive method as shown in Figure 1. It is designed to iteratively retrieve and reason on the subgraphs from KG, simulating a dynamic searching process through KG. FiDeLiS incorporates two modules Reasoning Path Retrieval-Augmented Generation (Path-RAG) and Deductive-Verification Guided Beam Search (DVBS), where Path-RAG is utilized to retrieve a highly pertinent set of nodes/relations from KGs to the user query and determine the reasoning step based on the relevance score discussed in Section 2.1. On the other hand, DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) for finding the most promising reasoning paths. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path extension. To prevent either premature stopping or excessive continuation of reasoning path extension, DVBS enhances the clarity of termination signals using deductive reasoning discussed in Section 2.2.
2.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
Path-RAG comprises three main steps: initialization, retrieval, reasoning step candidates construction, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
Initialization. We initiate the Path-RAG by generating node and edge embeddings using a pre-trained Language Model (LM). These embeddings are then stored in a nearest neighbor data structure. To elaborate, consider $e_{i}\in\mathcal{E},r_{i}\in\mathcal{R}$, where $\mathcal{E}$ refers to the entities and $\mathcal{R}$ represents the relations in the relational graph. We leverage the pre-trained LM (e.g., SentenceBert [33] and E5 [46]) to $e_{i}$ and $r_{i}$, yielding the representation $z(e^{i})$ and $z(r^{i})$:
$z(e^{i})=\operatorname{LM}\left(e_{i}\right)\in\mathbb{R}^{d},\\ z(r^{i})=\operatorname{LM}\left(r_{i}\right)\in\mathbb{R}^{d}$ (1)
where $d$ denotes the dimension of the output vector.
Retrieval. To retrieve the entities and relations that are relevant to the user query, the embeddings of them are then utilized to populate a nearest neighbor index. Specifically, given a query $q$, we leverage the LM to generate an exhaustive list of keywords or relation names $k_{i}\in\mathcal{K}$ that would be used to find the reasoning path from the knowledge graph to answer the query (as referred to prompt list in Section B). The generated list of keywords aims to maximize the coverage of potential reasoning steps and to enhance the subsequent decision-making process of LLMs by concentrating on a smaller set of relevant reasoning step candidates.
$\mathcal{K}=\text{LM}(\text{prompt}_{p},q)$ (2)
Then each $k_{i}^{m}\in\mathcal{K}$ is concatenated into a string $k_{i}$ and embedded into the same latent space by the LM to obtain $z(k)\in\mathbb{R}^{d}$, the retrieval process involves identifying the node or sequence of nodes that best aligns with the semantic content of the keywords as follow:
$\mathcal{E}_{m}=\operatorname{argtopm}_{i\in\mathcal{E}}\cos\left(z(k),z(e)\right),\mathcal{R}_{m}=\operatorname{argtopm}_{i\in\mathcal{R}}\cos\left(z(k),z(r)\right)$ (3)
where $z(e)$ and $z(r)$ are the embeddings of nodes $e$ and edges $r$, respectively. We use the cosine similarity function, $\cos(\cdot,\cdot)$, to measure the similarity between the keyword representation and the node/edge embeddings. The $\operatorname{argtopm}$ operation retrieves the top-m elements based on this similarity, providing a set of nodes $e_{k}$ and edges $r_{k}$ considered most relevant to the query. We denote the score of each entity and each relation as $S_{e}^{i}=\cos\left(z(k),z(e)\right)$ and $S_{r}^{i}=\cos\left(z(k),z(r)\right)$.
Reasoning Step Candidates Construction. To construct reasoning steps candidates based on the finite set of nodes $\mathcal{E}_{m}$ and edges $\mathcal{R}_{m}$ relevant to the query and the relevance score $S_{e}^{i}$ and $S_{r}^{i}$, we denote the reasoning step can be scored and retrieved based on the following criteria as follows. The goal of this scoring function is to leverage the triplets to represent the connectivity of the KGs. It provides some hints on whether the current operation/decision would lead to a malicious operation several steps later by incorporating the next-hop neighbor’s information.
$S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}\left(S_{r}^{j}+S_{e}^{j}\right)$ (4)
It should be noted that the entity and relation mentioned in the equation should be logically connected in the corresponding KG. $S(p)$ denotes the score of the reasoning step; $S_{r}^{i}$, $S_{e}^{i}$ denote the relevance score of the next step relation and the corresponding neighbor; $S_{r}^{j}$, $S_{e}^{j}$ denote the relevance score of one further hop of next step. Parameter $\alpha$ is used to balance between immediate accuracy and long-term prospects within the graph. A higher $\alpha$ value encourages the model to consider less immediate but potentially vital outcomes, While a lower $\alpha$ makes it focus primarily on the immediate next steps without considerable regard for the long-term impacts. The detailed description can be found in Appendix Algorithm 1.
Figure 1: Overview of FiDeLiS. The goal of FiDeLiS is to extend the multiple-step reasoning paths grounded by the external knowledge graph. We instantiate the $T$ step reasoning path extension as shown in the figure. First, Path-RAG is used to retrieve a highly pertinent set of nodes/relations from KGs to the user query and construct the reasoning step candidates. Then, the constructed reasoning steps candidates are forwarded to DVBS. DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates for finding the most promising reasoning step. Specifically, at timestep $T$, DVBS leverage the LLMs to choose the top-$k$ (here we set the $k=2$ for better clarification) reasoning steps and decide whether to continue the next search step or cease the reasoning path extension based on the deductive verification.
2.2 DVBS: Deductive-Verification Guided Beam Search
DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) for finding the most promising reasoning paths. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path extension. It comprises three main steps: planning, beam-search, deductive-verification, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
Plan-and-Solve. Inspired by the recent works regarding planning capabilities of LLMs [35, 56], this step aims to provide more hints for subsequent LLM decision-making. Specifically, we prompt LLM to generate the planning steps for searching the most promising reasoning paths initially denoted as $p$. The detailed prompts can be found in Section B.
Beam Search. In multi-step reasoning, a reasoning path of $T$ steps is sequentially generated through several time steps as $R=\left[s^{1},s^{2},\ldots,s^{T}\right]=s^{1:T}$ as shown in Figure 1. However, different from the typical text decoding process where each step consists of a single token, we follow the [50] and view a sequence of pair $(r,e)$ as a single step, where $r$ is the relation and $e$ is the corresponding entity. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path. We present the beam search at each step is modeled as follows:
$\mathcal{H}_{t}=\text{Top}_{K}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},p):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\right\}$ (5)
where $\mathcal{H}_{t-1}$ denote the reasoning path at the previous time step $t-1$; $s^{t}$ refers to the current reasoning step; $\mathcal{S}$ refers to all possible next steps $s^{t}$ candidates produced from the Path-RAG; $\text{LM}(s^{t}|q,s^{1:t-1},p)$ refers to the language models prediction for the next step $s^{t}$ given the current sequence $s^{1:t-1}$, the query $q$, and the planning context $p$; $\oplus$ refers to appending $s^{t}$ to the current sequence $h$; $\text{Top}_{K}$ refers to selecting the top $K$ (denoted as beam width in the experiment setting) sets of reasoning steps.
Deductive Verification. To mitigate the reasoning path extension terminates at the wrong point, we propose to leverage deductive reasoning capabilities of LLMs [26, 18] as a better criterion $C(x,s^{t},s^{1:t-1})\in\left\{0,1\right\}$ to automatically guide the decision making of LLMs in Equation 5. Specially, we first leverage LLMs to convert the user query to a declarative statement (as shown in Appendix G.2) and then use LLMs to judge whether the declarative statement of query $x^{\prime}$ can be deduced from current reasoning step $s^{t}$ and the previous reasoning steps $s^{1:t-1}$. We utilize the same backend LLM with prompt referred in Section B and define the deductive-verification criteria as follows:
$C(q^{\prime},s^{t},s^{1:t-1})=\begin{cases}1&\text{if declarative statement ofquery }q^{\prime}\text{ can be deduced from }s^{t}\text{ and }s^{1:t-1},\\ 0&\text{otherwise.}\end{cases}$ (6)
The overall DVBS can be modeled as follows:
$\mathcal{H}_{t}=\text{Top}_{B}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},p):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\text{ and }C(q^{\prime},s^{t},s^{1:t-1})=1\right\}$ (7)
This criterion is an essential component of our method for leveraging LLMs to iteratively validate each step of reasoning to make sure it follows logically and matches the original query. It acts as a clear sign to stop further reasoning when necessary, which helps avoid mistakes in the reasoning process and reduces unnecessary computations.
3 Experiments
We aim to answer the following research questions during this section: RQ1: How does FiDeLiS compare with existing baselines in terms of accuracy and efficiency on KGQA tasks? RQ2: How does the design of each individual component contribute to the overall performance of FiDeLiS? RQ3: How robustness of FiDeLiS perform under varying conditions?
3.1 Experiment Settings
Datasets & Metrics. We adopt three benchmark KGQA datasets: WebQuestionSP (WebQSP) [54], Complex WebQuestions (CWQ) [40] and CR-LT-KGQA [11] in this work. We follow previous work [29] to use the same training and testing splits for fair comparison over WebQSP and CWQ. Both WebQSP and CWQ can be reasoned using Freebase KGs [^1] https://github.com/microsoft/FastRDFStore. To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we tested our methods on CR-LT-KGQA (further discussed in Section 5). We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples. Each of these can be reasoned based on the Wikidata KG [^2] https://www.wikidata.org/wiki/Wikidata:Main_Page. The statistics of the datasets are given in Table 8 and Table 9.
To streamline the KGs, we utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ [29]. Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA as well. We assess the performance of the method by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
Baselines. (1) RoG [29]: embed structure knowledge graph from KGs into the underlying neural networks during the pretraining and fine-tuning process to generate the reasoning path and explanation. (2) ToG [38]: ask LLM to iteratively explore multiple possible reasoning paths on KGs until the LLM determines that the question can be answered based on the current reasoning paths. Our method FiDeLiS follows this paradigm. (3) NSM [12] utilizes the sequential model to mimic the multi-hop reasoning process. (4) KD-CoT [44] retrieves relevant knowledge from KGs to generate faithful reasoning paths for LLMs. (5) DeCAF [55] combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance on KGQA tasks. The details of each baseline are further discussed in Section G.1.
Implementation Details. We set the default beam width as $4$ and depth as $4$ without specific annotation. We set the $\alpha$ in Eq 4 as $0.3$ to ensure reproducibility.
Backboned LLMs. We assess our approach on closed- and open-source LLMs. For closed-source LLMs, we choose GPT-4-turbo, GPT-3.5-turbo [^3] We use the recent gpt-4-turbo model that is released in 2024-04-09 from https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 and gpt-3.5-turbo model notated as “gpt-3.5-turbo-0125” from https://platform.openai.com/docs/models/gpt-3-5-turbo. to report and compare the results on all datasets. We use Llama-2-13B [42] and Mistral-7B [19] as our open-source LLMs to conduct cost–performance analysis on different datasets. The experiment results of open-source models can be found in Table 10.
3.2 RQ1: KGQA Performance Comparison
Comparison between baseline methods. Table 1 clearly showcases the comparative effectiveness of FiDeLiS, against different baselines. It shows that FiDeLiS outperforms all baselines integrated with gpt-4-turbo, even for strong baselines being trained, such as DeCAF [55] and RoG [29]. Across all prompting-based methods, gpt-4-turbo achieves higher performance compared to gpt-3.5-turbo, especially on the CWQ dataset. It indicates that gpt-4-turbo has a better understanding and processing capability for complex queries. The CR-LT dataset seems more demanding than WebQSP and CWQ, as implied by the consistently lower accuracy scores. However, FiDeLiS shows consistent enhancements compared to other baselines, demonstrating its capability to handle long-tail entities by referring to the knowledge from KGs and processing more complex queries effectively.
Table 1: Comparison of FiDeLiS with baseline methods and different backbone LLMs. We replicate the outcomes of ToG and RoG, and retrieve other baseline results directly from the original paper. We utilize 5 demonstrations as our default setting for FiDeLiS, ToG, Few-shot, and CoT. The experiment results of open-source models can be found in Table 10.
Backend Models Methods WebQSP CWQ CR-LT
Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
Prompting - LLM Only gpt-3.5-turbo Zero-shot 54.37 52.31 34.87 28.32 32.74
Few-shot 56.33 53.12 38.52 33.87 36.61
CoT 57.42 54.72 43.21 35.85 37.42
Prompting - LLM Only gpt-4-turbo Zero-shot 62.32 59.71 42.71 37.93 37.74
Few-shot 68.65 62.71 51.52 43.70 43.61
CoT 72.11 65.37 53.51 44.76 45.42
Finetuning - LLM + KG NSM [12] 74.31 - 53.92 - -
CBR-KBQA [6] - - 67.14 - -
DeCAF [55] 82.1 - 70.42 - -
KD-CoT [44] 73.7 50.2 50.5 - -
RoG [29] 83.15 69.81 61.39 56.17 60.32
Prompting - LLM + KG gpt-3.5-turbo ToG [38] 75.13 72.32 57.59 56.96 62.48
FiDeLiS 79.32 76.78 63.12 61.78 67.34
Prompting - LLM + KG gpt-4-turbo ToG [38] 81.84 75.97 68.51 60.20 67.24
FiDeLiS 84.39 78.32 71.47 64.32 72.12
3.3 RQ2: Ablation Study of FiDeLiS.
Table 2 provides a detailed view of how removing specific components from FiDeLiS impacts its performance. The results emphasize that some features, such as Path-RAG and beam search in DVBS, are particularly significant, especially for tackling complex reasoning tasks on datasets like CWQ. When Path-RAG and beam-search within DVBS are excluded, we observe substantial declines in performance. This underlines their importance in our model’s ability to handle intricate reasoning effectively. Furthermore, the integration of the reasoning and planning steps within DVBS plays a critical role in maintaining robust performance across varied datasets. One aspect highlighted by our study is the role of deductive verification during the beam search phase in DVBS. This feature enhances the model’s ability to delineate and reinforce decision boundaries, thus safeguarding against reasoning errors.
Table 2: Ablation Studies of FiDeLiS, we test the ablation study using model gpt-3.5-turbo-0125. DVBS refers to “Deductive-verification guided Beam Search’. “$\Delta$” refers to the performance gap between each component and the entire method. The top-3 performance gaps on each dataset are highlighted in green, with darker shades indicating significant performance gap.
- Components WebQSP CWQ CR-LT
Hits@1 (%) $\Delta$ Hits@1 (%) $\Delta$ Acc (%) $\Delta$
- FiDeLiS 79.32 0.00 63.12 0.00 67.34 0.00
w/o Path-RAG using vanilla retriever 72.35 6.97 57.11 6.01 59.78 7.56
using ToG 75.11 4.21 59.47 3.65 63.47 4.07
w/o DVBS w/o last step reasoning 75.68 3.64 59.45 3.67 63.72 3.62
w/o planning 76.23 3.09 60.14 2.98 64.13 3.21
w/o beam-search 60.35 18.97 49.78 13.34 61.87 5.47
w/o deductive-verifier 74.13 5.19 57.23 5.89 63.89 3.45
3.4 RQ3: Robustness Analysis
Comparison using different embedding backbones. Table 3 demonstrates the performance of the Path-RAG using various embedding models as backbones. It shows that the integration of the Openai-Embedding Model significantly enhances the performance across all datasets, indicating its superior capability in understanding and processing natural language queries, which leads to more effective retrieval. For instance, the Openai-Embedding Model increases the Hits@1 by 13.04% in the WebQSP dataset and 14.73% in the CWQ dataset compared to the base BM25 [34] model. Furthermore, this improvement is consistent and highlights the feasibility and potential of pairing more powerful language models with structured reasoning frameworks to address complex information retrieval and question answering challenges.
Table 3: Performance FiDeLiS using different Path-RAG embedding backbone models. ‘Openai-Embedding Model’ refers to text-embedding-3-small, the recent embedding model released from OpenAI.
Methods Backbones WebQSP CWQ CR-LT
Hits@1 (%) Hits@1 (%) Acc (%)
Vanilla Retriever w/ BM25 [34] 58.31 48.39 50.73
w/ SentenceBert [33] 62.74 50.14 51.80
w/ E5 [46]) 68.42 52.84 54.31
w/ Openai-Embedding-Model 72.35 57.11 59.78
Path-RAG w/ BM25 [34] 70.34 56.11 58.77
w/ SentenceBert [33] 73.45 58.41 60.45
w/ E5 [46]) 77.93 62.74 65.23
w/ Openai-Embedding-Model 79.32 63.12 67.34
Comparison using different searching widths and depths. To explore the influence of the search width and depth on FiDeLiS’s performance, we conduct experiments under settings with depths ranging from 1 to 4 and widths from 1 to 4. As shown in Figure 2, FiDeLiS’s performance improves with the search depth and width and when the searching depth is larger than 3, FiDeLiS’s performance start to decrease on both WebQSP and CWQ. This is because only a small part of questions in these two datasets have reasoning depths greater than 3, and also FiDeLiS implements stricter rules for terminating searches using deductive reasoning. However, for searching width, we find that the performance is consistent and implies the FiDeLiS’s performance could potentially be improved with the increment of the exploration breadth. Considering the computational cost (which increases linearly with the depth), we set both the default beam width and depth as 4 without specific annotations in all the experiments.
Figure 2: Performance of FiDeLiS with different beam search widths and depths.
Comparison between Path-RAG and Vanilla Retriever. To verify the validity of the Path-RAG compared to vanilla retrievers. We conduct the experiments as shown in Figure 3. For the vanilla retriever, we concatenate each entity with its corresponding relation to form a reasoning step. We then calculated the cosine similarity between the embeddings of each reasoning step and the query embeddings to select candidate paths.
Figure 3: Comparison between Path-RAG and Vanilla Retriever.
We calculate the coverage ratio (CR) of the retrieved reasoning paths and the ground-truth reasoning paths as follows: $\text{CR}=\left(\frac{N_{\text{retrieved}}\cap N_{\text{ground-truth}}}{N_{\text{ground-truth}}}\right)\times 100\%$, where $N_{\text{retrieved}}$ is the number of retrieved reasoning paths candidates, and $N_{\text{ground-truth}}$ is the total number of ground-truth reasoning paths. We observe that compared to the vanilla retriever, Path-RAG achieves a higher CR value and shows better alignment with the ground-truth paths. It demonstrates that our method can better leverage the connections that are often missed by simpler retrieval models. This feature is critical in subsequent LLM to focus on the information of interest and arrive at a correct and reasonable answer.
Table 4: Average depths of the returned reasoning paths. The average depth of ground-truth reasoning paths of WebQSP is 2.3, for CWQ is 3.2, and for CR-LT is 4.7.
Method WebQSP CWQ CR-LT
ToG 3.1 4.1 5.2
FiDeLiS 2.4 2.8 4.6
Effectiveness of deductive-verification. To verify the effectiveness of deductive-verification, we calculate the average depths of the grounded reasoning path from ToG and our methods in Table 4. It shows that FiDeLiS consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to ToG. This implies that our method may offer more precise termination signals and potentially more accurate reasoning paths.
Error Analysis regarding Whole Path Generation.
Figure 4: Validation of whole path generation from RoG [29] over WebQSP.
To verify the faithfulness of our stepwise method, we conduct the error analysis regarding the whole reasoning path generation. Specifically, we conduct an analysis of the validity of the whole reasoning path generation using the baseline methods RoG [29]. The result is shown in Figure 4. The definition of the metric validity ratio (VR) is the ratio of reasoning steps that existed in the knowledge graph to the total number of the reasoning steps in the output reasoning path: $\text{VR}=\left(\frac{N_{\text{valid-steps}}}{N_{\text{all-steps}}}\right)\times 100\%$, where $N_{\text{valid-steps}}$ is the number of reasoning steps that existed in the knowledge graph, and $N_{\text{all-steps}}$ is the number of all the reasoning steps in the reasoning path. We note that 33% of the generated reasoning paths are invalid, 18% exhibit formatting errors, and 15% include relations that do not exist in the knowledge graph. This illustrates the limitations of embedding knowledge graphs into base LLM models, particularly in controlling the generation process.
Case Study. To verify whether FiDeLiS can conduct faithful reasoning on real cases, we also present a case study in Table 5. We can find that FiDeLiS illustrates better performance in identifying multiple possible governmental forms—Theocracy, Unitary state, and Islamic republic established in Iran in 1979, as compared to existing models such as GPT + CoT, RoG, and ToG. This observation highlights the importance of recognizing the complexity of Iran’s government system, which combines elements of religion and democracy. Different models handle this complexity with varying levels of detail. For example, GPT + CoT and RoG mainly identified Iran as an ‘Islamic Republic’, a change that occurred after the 1979 Iranian Revolution, but they missed other possible descriptions. The ToG model even gives mixed results, such as simply naming it ‘Iran’, which shows confusion in its analysis. However, our method shows a better understanding of the context and provides a more detailed and accurate set of answers. This illustrates FiDeLiS strength and efficiency in dealing with intricate questions.
Table 5: Case study of FiDeLiS. The wrong answers are highlighted with red color, and the correct answers are highlighted with blue color. In this case, FiDeLiS scored an F1 of 0.857, outperforming the RoG method (0.4).
Question What form of government is in the country that uses the Iranian Rail and was established in 1979?
Answer ’Parliamentary system’, ’Presidential system’, ’Islamic republic’, ’Unitary state’, ’Theocracy’
GPT + CoT Think step by step. First, identify the country: the country that uses the Iranian Rail is Iran. Then, retrieve the historical Context of 1979: In 1979, Iran underwent a significant transformation with the Iranian Revolution, which led to the overthrow of the Pahlavi monarchy. Third, following the revolution, the country adopted a new form of government based on Islamic law. The result was the establishment of the Islamic Republic of Iran, which combines elements of a theocracy and a democracy, where religious leaders have significant influence over the state governance. So, the government form is the Islamic Republic of Iran, established in 1979 following the Iranian Revolution.
RoG location.country.form_of_government, location.location.containedby location.country.form_of_government’ location.location.containedby $\to$ government.form_of_government.countries The results are Islamic republic.
ToG Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic $\to$ government.form_of_government.countries $\to$ Iran Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy $\to$ government.form_of_government.countries $\to$ Iran Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state $\to$ government.form_of_government.countries $\to$ Iran The results are Iran.
FiDeLiS Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state The results are Theocracy, Unitary state, Islamic republic.
4 Related Work
KG-enhanced LLM. KGs have advantages in dynamic, explicit, and structured knowledge representation and techniques combining LLMs with KGs have been studied [32]. Early studies [29, 55] embed structured knowledge from KGs into the underlying neural networks during the pretraining or fine-tuning process. However, the reasoning steps generated by LMs are observed to be prone to errors, which could be non-existent in the knowledge graph, and can subsequently lead to incorrect reasoning during inference. Also, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating [17]. Another solutions are to keep the reasoning model as an agent to explore the external structure knowledge source [38, 9, 20, 44] ToG [38] directly employs LLMs to output scores for candidate selections from the KG. This model operates at the relation-entity level of the KG, aiming to identify relevant triples that aid the LLM in making accurate and responsible final answer predictions.
Knowledge Graph-based Question Answering. Conventional embedding-based methods represent the entities and relations in embedding space and design special model architectures (e.g., Key-Value memory networks, sequential models, and graph neural networks) to reason answers [30, 12, 52]. To integrate LLMs for KGQA, retrieval-augmented methods aim to retrieve the relative facts from the KGs to improve the reasoning performance [25, 22]. Recently, UniKGQA [21] which unifies the graph retrieval and reasoning process into a single model with LLMs, achieves STOA performance. Semantic parsing methods convert the question into a structural query (e.g., SPARQL) by LLMs, which can be executed by a query engine to reason the answers on KGs [39, 24]. However, these methods heavily rely on the quality of generated queries. If the query is not executable, no answers will be generated. DECAF [55] combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance.
5 Conclusion
This paper proposes a retrieval-exploration interactive method specifically designed to enhance intermediate steps of LLM reasoning grounded by KGs. The Path-RAG module and the use of deductive reasoning as a calibration tool effectively guide the reasoning process, leading to more accurate knowledge retrieval and prevention of misleading reasoning chains. Extensive experiments demonstrate that our method, being training-free, not only reduces computational costs but also offers superior generality, consistently outperforming established strong baselines across three distinct benchmarks. We believe this study will significantly benefit the integration of LLMs and KGs, or serve as an auxiliary tool to enhance the interpretability and factual reliability of LLM outputs.
Appendix A Defintions
A.1 Defining a Valid Reasoning Trace/Path
Definition 1. A reasoning step is a pair $(r,e)$, where $r$ is the relation and $e$ is the corresponding entity.
Definition 2. A reasoning path $\mathcal{P}$ is a pair $(s,\mathcal{T})$, where $s$ is the starting entity for the reasoning path, and $\mathcal{T}$ is a sequence of reasoning steps $\mathcal{T}=\left\{t_{0},t_{1},\ldots,t_{n}\right\}$ and $t_{k}=(r_{k},e_{k})$. The $t_{k}\in\mathcal{R}$ denotes the $k$-th reasoning step in the path and $n$ denotes the length of the path.
[…]
A.2 Knowledge Graph-based Question Answering (KGQA)
In this work, we mainly focus on the question answering task based on the knowledge graph.
Knowledge Graphs (KGs) contain abundant factual knowledge in the form of a set of triples: $\mathcal{G}=$$\left\{\left(e,r,e^{\prime}\right)\mid e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\right\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations, respectively.
Knowledge Graph Question Answering (KGQA) is a typical reasoning task based on KGs. Given a natural language question $q$ and a $\mathrm{KG}$ $\mathcal{G}$, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ based on knowledge from $\mathcal{G}$, i.e., $a=f(q,\mathcal{G})$. Following previous works [29, 28], we assume the entities $e_{q}\in\mathcal{T}_{q}$ mentioned in $q$ and answers $a\in\mathcal{A}_{q}$ are labeled and linked to the corresponding entities in $\mathcal{G}$, i.e., $\mathcal{T}_{q},\mathcal{A}_{q}\subseteq\mathcal{E}$.
Appendix B Prompt List
In this section, we show all the prompts that need to be used in the main experiments. The In-Context Few-shot refers to the few-shot examples we used for in-context learning.
B.1 Plan-and-solve
You are a helpful assistant designed to output JSON that aids in navigating a knowledge graph to answer a provided question. The response should include the following keys:
(1) ’keywords’: an exhaustive list of keywords or relation names that you would use to find the reasoning path from the knowledge graph to answer the question. Aim for maximum coverage to ensure no potential reasoning paths will be overlooked;
(2) ’planning_steps’: a list of detailed steps required to trace the reasoning path with. Each step should be a string instead of a dict.
(3) ’declarative_statement’: a string of declarative statement that can be transformed from the given query, For example, convert the question ’What do Jamaican people speak?’ into the statement ’Jamaican people speak *placeholder*.’ leave the *placeholder* unchanged; Ensure the JSON object clearly separates these components.
In-Context Few-shot
Q: {Query}
A:
B.2 Deductive-verification
You are asked to verify whether the reasoning step follows deductively from the question and the current reasoning path in a deductive manner. If yes return yes, if no, return no".
In-Context Few-shot
Whether the conclusion ’{declarative_statement}’ can be deduced from ’{parsed_reasoning_path}’, if yes, return yes, if no, return no.
A:
B.3 Adequacy-verification
You are asked to verify whether it’s sufficient for you to answer the question with the following reasoning path. For each reasoning path, respond with ’Yes’ if it is sufficient, and ’No’ if it is not. Your response should be either ’Yes’ or ’No’.
In-Context Few-shot
Whether the reasoning path ’{reasoning_path}’ be sufficient to answer the query ‘{Query}’, if yes, return yes, if no, return no.
A:
B.4 Beam Search
Given a question and the starting entity from a knowledge graph, you are asked to retrieve reasoning paths from the given reasoning paths that are useful for answering the question.
In-Context Few-shot
Considering the planning context {plan_context} and the given question {Query}, you are asked to choose the best {beam_width} reasoning paths from the following candidates with the highest probability to lead to a useful reasoning path for answering the question. {reasoning_paths}. Only return the index of the {beam_width} selected reasoning paths in a list.
A:
B.5 Reasoning
Given a question and the associated retrieved reasoning path from a knowledge graph, you are asked to answer the following question based on the reasoning path and your knowledge. Only return the answer to the question.
In-Context Few-shot
Question: {Query}
Reasoning path: {reasoning_path}
Only return the answer to the question.
A:
Appendix C Potential Impacts and Limitations
The proposed retrieval-exploration interactive method, FiDeLiS, holds the potential to significantly enhance the performance of LLMs by addressing the issue of hallucinations and enabling deep, responsible reasoning. By integrating KGs with LLMs, the method offers more accurate knowledge retrieval, leveraging deductive reasoning capabilities to guide the reasoning process and avoid logical fallacies. The stepwise and generalizable approach, along with the use of deductive verification as a stop criterion, may reduce unnecessary computation and misleading reasoning chains. Additionally, the training-free nature of this method combined with its lower computational cost could be easily acted as a plugin-in for other methods/frameworks.
However, there are limitations to consider. As the method is still in its early stages of development, there may be unforeseen challenges or drawbacks when applied to real-world scenarios. The reliance on external knowledge graphs means that the quality and coverage of these resources can impact the overall effectiveness of the method. There may also be computational constraints when working with very large or complex graphs. Additionally, while the method has shown promise in benchmarks, its performance in more diverse or specialized tasks is yet to be evaluated. Despite these challenges, the research presents a promising direction for improving the reasoning capabilities of LLMs.
Appendix D Efficiency of FiDeLiS
The bottleneck of computation is the beam search process, which contributes to $N*D$ times LLM calling, where $D$ is the depth (or equivalently length) of the reasoning path, and $N$ is the width of the beam-search (how many paths are remained in the pool in each iteration). Specifically, we need to call $ND+D+C$ times LLM for each sample question, where $C$ is a constant (equals to $1$ if there is no error occurs when calling the API). [38] illustrates that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling. In this way, we can reduce the number of LLM calling from $ND+D+C$ to $D+C$. However, [38] shows that this optimization sacrifices the accuracy due to the weaker scoring model in decision making.
Table 6: Average number of LLM calls per question.
Dataset WebQSP CWQ CR-LT
Average LLM Calls 10.7 15.2 17.3
We noted that $ND+D+C$ is the maximal computational complexity. In most cases, FiDeLiS does not need $ND+D+C$ LLM calls for a question because the whole reasoning process might be early stopped before the maximum reasoning depth $D$ is reached if LLM determines the query can be deductive reasoning by the current retrieved reasoning paths. As an illustration, Table 6 shows the average numbers of LLM calls per question needed by FiDeLiS on different datasets. It can be seen that in three KGQA datasets, the average numbers of LLM calls (ranging from ) are smaller than $21$, which is the theoretical maximum number of LLM calls calculated from $ND+D+C$ when $N=4$ and $D=4$. We can also see that this average number gets even smaller for dataset covering a lot of single-hop reasoning questions, such as WebQSP.
Table 7: Average token usage of LLM calls per question.
Dataset WebQSP CWQ CR-LT
Average Token Usage 2,452 2,741 1,824
Cost efficiency. Another benefit of FiDeLiS is the cost efficiency. We control the token usage for each LLM calling through Path-RAG module. Specifically, Path-RAG only retrieves the top-n reasoning steps as the candidates for each beam search decision making, before making calls to the LLM. This approach ensures token expenditure without compromising the model’s performance, especially in some cases where hundreds of neighbors exist around a entity node. Table 7 shows the average token usage per question needed by FiDeLiS on different datasets. It can be seen that in three KGQA datasets, the numbers of token usages are ranging from 1,824 to 2,751, which could significantly reduce the cost for running the method.
Appendix E Variants of Path-RAG
Variant-1: To enhance the scoring function for reasoning in a knowledge graph and ensure each entity and relation are consistently linked, we can modify the scoring function to more explicitly account for the connectivity and interdependences between entities and relations throughout the path. This can be done by incorporating an accumulative sum that keeps track of previous scores and emphasizes path completeness. The revised formula introduces variables for the accumulated scores of entities and relations, which will help ensure all links in the path are considered:
$S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}\left(S_{r}^{j}+S_{e}^{j}\right)+\beta\sum_{k=0}^{i-1}\left(S_{r}^{k}+S_{e}^{k}\right)$ (8)
where: (1) $S_{r}^{i}$ and $S_{e}^{i}$ denote the relevance score of the current step relation and the corresponding neighboring entity, respectively. (2) $S_{r}^{j}$ and $S_{e}^{j}$ denote the relevance score of the 1-hop relation and the corresponding neighboring entity from node $N_{i}$. (3) The new term, $\sum_{k=0}^{i-1}\left(S_{r}^{k}+S_{e}^{k}\right)$, accumulates the scores from all previous steps from the start up to the step before the current, enhancing the influence of the entire path on the decision of the next step. (4) $\alpha$ and $\beta$ are weight factors that control the influence of the maximum score of neighboring nodes and the accumulated score to adjust the balance in prioritizing extended paths versus immediate gains. This adjustment helps to: firstly, reward paths that maintain strong links throughout, discouraging breaks or weakly connected steps in the reasoning process; secondly, enhance the depth of consideration in the decision-making process, looking both at immediate and long-term implications of each step.
Variant-2: We also consider to incorporate the randomness into the scoring function to balance the quality-diversity trade-off in searching for better reasoning paths as follows:
$S(p)=\lambda_{1}\left(S_{r}^{i}+S_{e}^{i}+\alpha\sum_{j\in N_{i}}(S_{r}^{j}+S_{e}^{j})\right)+\lambda_{2}\sum_{k=0}^{i-1}\left(D(p,P_{k})\right)$ (9)
where $D(p,P_{\text{sel}})$ refers to the distance between the selected path with the grounded reasoning paths. while $\lambda_{1}$ and $\lambda_{2}$ are weighting factors that balance the importance of internal quality metrics against diversity metrics.
Appendix F Additional Related Works: Reasoning with LLM Prompting
With the development of LLMs, many creative ways to leverage LLMs are proposed: End-to-End, Chain-of-Thought, and Semantic Parsing/Code Generation.
End-to-End methods [4, 15] aims to leverage LLMs to generate final answers directly, often done by providing a task description and/or a few examples for in-context learning. While it offers convenience, however, it suffers from the un-interpretability of the generation, and the lack of explicit steps and reliance solely on the LLMs’ training data may result in un-robustness and exhibit sensitivity to slight input variations.
Chain-of-Thought methods [48, 23, 8] emphasize breaking down complex reasoning into a series of intermediate steps to support complex reasoning. However, CoT suffers from unreliability and uncontrollability since the generated reasoning steps may not always align with the intended logical progression or may produce incorrect or inconsistent answers when multiple questions are posed consecutively.
Plan-and-solve methods [45] prompts LLMs to generate a plan and conduct reasoning based on it. DecomP [12] prompts LLMs to decompose the reasoning task into a series of sub-tasks and solve them step by step.
Semantic Parsing/Code Generation methods [5, 53, 7] leverages LLMs to convert natural language queries into executable code or structured representations. This approach enables more precise control over the output and facilitates better interpretability. However, it has its limitations, such as the limited coverage of programming language grammar and semantics. This restricts the model’s ability to handle complex programming tasks and may require additional techniques to handle out-of-domain queries effectively. [5] propose to leverage LLMs to bridge the out-of-domain queries, however, the form of knowledge injection still lacks a thorough exploration and the trigger functions are still naive for complex queries. However, the problem of hallucinations and lack of knowledge affect the faithfulness of LLMs’ reasoning. ReACT [51] treats LLMs as agents, which interact with the environment to get the latest knowledge for reasoning. To explore faithful reasoning, FAME [16] introduces the Monte-Carlo planning to generate faithful reasoning steps. RR [13] and KD-CoT [44] further retrieve relevant knowledge from KGs to produce faithful reasoning plans for LLMs.
Appendix G Experiment Details
G.1 Baselines
RoG [29]: It aims to embed structure knowledge from KGs into the underling neural networks during the pretraining and fine-tuning process. However, the controllablity of the reasoning paths generation is hard to achieve since the decoding process can be shortsighted, i.e., it is difficult to tell whether the current operation/decision would lead to a malicious operation several steps later. Also, the reasoning steps generated by LLMs are observed to be prone to errors, which could be non-exist in the knowledge graph, and can subsequently lead to incorrect reasoning during inference. Additionally, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating [17]. Compared to RoG, FiDeLiS propose to employ beam search tailored for intermediate reasoning steps and guide the beam searching process by controlling the error of each reasoning step to prevent accumulation thought the chaining.
ToG [38]: It aims step-wise grounding to elicit more accurate results. However, it directly employs LLMs to output scores for candidate selections from the KG, which require an LLM/LM to iteratively score candidate plans, which are resource-consuming in terms of both time and computing. Also, the performance of the proposed ToG [38] seems rely heavily on the use of very strong LLMs (including Codex [3], GPT-4 [31]) to do the final reasoning; and the true potential of enabling few-shot grounded reasoning path understanding has yet to be fully realized.
G.2 Datasets
Table 8: Statistics of the number of answers for questions in WebQSP and CWQ.
Dataset #Ans = 1 2 $\geq$ #Ans $\leq$ 4 5 $\geq$ #Ans $\leq$ 9 #Ans $\geq$ 10
WebQSP 51.2% 27.4% 8.3% 12.1%
CWQ 70.6% 19.4% 6% 4%
Table 9: Statistics of the question hops in WebQSP, CWQ and CR-LT-KGQA.
Dataset 1 hop 2 hop $\geq$ 3 hop
WebQSP 65.49 % 34.51% 0.00%
CWQ 40.91 % 38.34% 20.75%
CR-LT 5.31 % 43.22% 51.57%
Figure 5: Distribution of CR-LT-KGQA dataset.
Motivation of CR-LT-KGQA The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities. These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs. Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering. Against this backdrop, we specifically chose CR-LT-KGQA for our evaluation as it includes queries related to obscure or long-tail entities, the distribution of the frequency and entity popularity of CR-LT is shown as Figure 5. Such scenarios are where KGs play a critical role because they provide a reliable source of verifiable information, crucial when LLMs encounter entities not well covered in their training data. By testing our methods on CR-LT-KGQA, we aim to explore how effective LLMs can operate when combined with KGs, especially in contexts involving less common knowledge domains where LLM performance typically declines. This evaluation helps us understand the extent to which KGs remain necessary for supporting LLMs in a diverse range of query scenarios.
Table 10: Comparison of FiDeLiS using different backbone LLMs.
Backend Models Methods WebQSP CWQ CR-LT
Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
Llama-2-13B FiDeLiS 72.34 69.78 58.41 54.78 60.87
Mistral-7B FiDeLiS 74.11 70.23 60.71 56.87 63.12
Example of Deductive Verification Question: “Who is the ex-wife of Justin Bieber’s father?" After one round of beam searching, the current reasoning path is “Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber", and the next step candidates are “1. people.married_to.person $\to$ Erin Wagner"; 2. people.person.place_of_birth $\to$ US, $\ldots$". The deductive reasoning can be formulated as: Premises: Justin_bieber $\to$ people.person.father —> Jeremy_bieber (from the current reasoning path) Jeremy_bieber $\to$ people.married_to.person $\to$ Erin Wagner (from the next step candidates) Conclusion: Erin Wagner is the ex-wife of Justin Bieber’s father. (Using LLM zero-shot to reformat the question to a cloze filling, and use the last entity from the next step candidate"Erin Wagner” to fill the cloze) The prompt will be whether the conclusion can be deduced from the given premises, if yes, return “yes”, it no, return “no".
Appendix H Algorithm Description
Algorithm 1 Path-RAG Initialization and Retrieval Process
1:Initialization:
2:for all $e_{i}\in\mathcal{E},r_{i}\in\mathcal{R}$ do
3:     $z_{e}^{i}=\operatorname{LM}(e_{i})$triangleright Embed entities
4:     $z_{r}^{i}=\operatorname{LM}(r_{i})$triangleright Embed relations
5:end for
6:Populate nearest neighbor index with $\{z_{e}^{i}\}$ and $\{z_{r}^{i}\}$triangleright Facilitate retrieval
7:procedure Retrieve(query $q$)
8:     $\mathcal{K}_{i}=\operatorname{LM}(\text{`prompt'},q)$triangleright Generate keywords
9:     for all $k_{i}^{m}\in\mathcal{K}_{i}$ do
10:         $k_{i}\leftarrow\text{concatenate}(k_{i}^{m})$
11:         $z_{k}=\operatorname{LM}(k_{i})$triangleright Embed concatenated keywords
12:         $\mathcal{E}_{k}=\operatorname{argtopk}_{i\in\mathcal{E}}\cos(z_{k},z_{e}^{i})$triangleright Retrieve top-k entities
13:         $\mathcal{R}_{k}=\operatorname{argtopk}_{i\in\mathcal{R}}\cos(z_{k},z_{r}^{i})$triangleright Retrieve top-k relations
14:     end for
15:     return $\mathcal{E}_{k},\mathcal{R}_{k}$
16:end procedure
17:procedure ScorePath($\mathcal{E}_{k},\mathcal{R}_{k}$)
18:     Initialize $\text{Score}\leftarrow 0$
19:     for each $e_{k}\in\mathcal{E}_{k}$ and $r_{k}\in\mathcal{R}_{k}$ do
20:         Calculate $S_{e}^{i},S_{r}^{i}\leftarrow\cos(z_{k},z_{e}^{i}),\cos(z_{k},z_{r}^{i})$triangleright Compute similarity scores
21:         $S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}(S_{r}^{j}+S_{e}^{j})$triangleright Score path using Eq. 4
22:         $\text{Score}\leftarrow\max(\text{Score},S(p))$triangleright Update max score
23:     end for
24:     return $\text{Score},p$
25:end procedure
Algorithm 2 Deductive-Verification Guided Beam Search
1:User query $x$, Beam width $B$
2:Reasoning path $s^{1:T}$
3:Initialize $\mathcal{H}_{0}=\{\emptyset\}$
4:Utilize LLM to generate from $x$:
5:      Planning steps.
6:      Declarative statement $x^{\prime}$.
7:for $t=1$ to $T$ do
8:     for each $h\in\mathcal{H}_{t-1}$ do
9:         Generate possible next steps $s^{t}\in\mathcal{S}$ using Path-RAG.
10:         for each $s^{t}$ do
11:              Compute $C(x^{\prime},s^{t},s^{1:t-1})$ using LLM:
12:                    $C(x^{\prime},s^{t},s^{1:t-1})=\begin{cases}1&\text{if }x^{\prime}\text{ can bededuced from }s^{t}\text{ and }s^{1:t-1},\\ 0&\text{otherwise.}\end{cases}$
13:              if $C(x^{\prime},s^{t},s^{1:t-1})=1$ then
14:                  Append $s^{t}$ to $h$ to form new hypothesis $h^{\prime}$.
15:                  Add $h^{\prime}$ to $\mathcal{H}_{t}$.
16:              end if
17:         end for
18:     end for
19:     $\mathcal{H}_{t}=\text{Top}_{B}(\mathcal{H}_{t})$ based on scoring function (like plausibility or likelihood).
20:end for
21:return the best hypothesis from $\mathcal{H}_{T}$.

### v2

@@
-FiDeLiS: Faithful Reasoning in Large Language Model for Knowledge Graph Question Answering
+FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering
 Abstract
-While large language models (LLMs) have achieved significant success in various applications, they often struggle with hallucinations, especially in scenarios that require deep and responsible reasoning.
-These issues could be partially mitigated by integrating external knowledge graphs (KG) in LLM reasoning.
-However, the method of their incorporation is still largely unexplored.
-In this paper, we propose a retrieval-exploration interactive method, FiDeLiS to handle intermediate steps of reasoning grounded by KGs.
-Specifically, we propose Path-RAG module for retrieving useful intermediate knowledge from KG for LLM reasoning.
-We incorporate the logic and common-sense reasoning of LLMs and the topological connectivity of KGs into the knowledge retrieval process, which provides more accurate retrieving performance.
-Furthermore, we propose to leverage deductive reasoning capabilities of LLMs as a better criterion to automatically guide the reasoning process in a stepwise and generalizable manner.
-Deductive verification serves as a precise indicator for when to cease further reasoning, thus avoiding misleading the chains of reasoning and unnecessary computation.
-Extensive experiments show that our method, as a training-free method with lower computational cost and better generality outperforms the existing strong baselines in three benchmarks.
-The code of this paper will be released later.
+Large language models are often challenged by generating erroneous or ‘hallucinated’ responses, especially in complex reasoning tasks.
+To mitigate this, we propose a retrieval augmented reasoning method, FiDeLiS, which enhances knowledge graph question answering by anchoring responses to structured, verifiable reasoning paths.
+FiDeLiS uses a keyword-enhanced retrieval mechanism that fetches relevant entities and relations from a vector-based index of KGs to ensure high-recall retrieval.
+Once these entities and relations are retrieved, our method constructs candidate reasoning paths which are then refined using a stepwise beam search.
+This ensures that all the paths we create can be confidently linked back to KGs, ensuring they are accurate and reliable.
+A distinctive feature of our approach is its blend of natural language planning with beam search to optimize the selection of reasoning paths.
+Moreover, we redesign the way reasoning paths are scored by transforming this process into a deductive reasoning task, allowing the LLM to assess the validity of the paths through deductive reasoning rather than traditional logit-based scoring.
+This helps avoid misleading reasoning chains and reduces unnecessary computational demand.
+Extensive experiments demonstrate that our method, even as a training-free method which has lower computational costs and superior generality, outperforms established strong baselines across three datasets.
+The code of this paper will be released at https://anonymous.4open.science/r/FiDELIS-E7FC.
 1 Introduction
-The emergence and application of large language models (LLMs) [1, 31, 27] have attracted widespread attention from researchers and the general public.
-It demonstrates remarkable reasoning capabilities, managing to solve complex reasoning problems through step-by-step thinking and planning [47, 49].
-However, the reasoning of LLMs is not invariably reliable and may conflict with factual reality [32], which limits the application in areas requiring high reliability, such as healthcare [14] and science [41].
-The knowledge graph (KG) stores high-quality domain-specific knowledge in structured triplets, such as Wikidata [43], YAGO [36], and NELL [2].
-KGs offer structured, explicit, and updated factual knowledge, which provides a faithful knowledge source for reasoning.
+The emergence and application of large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; LLaMa-v3, 2024) have attracted widespread attention from researchers and the general public.
+They demonstrate remarkable reasoning capabilities, managing to solve complex reasoning problems through step-by-step thinking and planning (Wang et al., 2022b; Wei et al., 2022b).
+However, the reasoning of LLMs is not always reliable and may conflict with factual reality (Pan et al., 2024), limiting their application in areas requiring high reliability, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
+Knowledge graphs (KGs) store high-quality knowledge in structured triplets, such as Wikidata (Vrandečić & Krötzsch, 2014), YAGO (Suchanek et al., 2007), and NELL (Carlson et al., 2010).
+KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning.
 Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance.
-This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretive process transparent.
+This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
 Due to its reliability and interpretability, it is considered as a promising method to improve the reliability of LLM reasoning.
-Therefore, several attempts have been conducted to integrate KGs with language models [32, 29, 17, 38].
-Among them, the knowledge graph question answering (KGQA) is a critical task to incorporate the knowledge of KG into reasoning models [12, 44, 55].
-Concretely, KGQA faces the following main challenges:
-(I) How to retrieve specific knowledge from KGs to help reasoning precisely?
-Existing solutions include direct retrieval [37, 21] and semantic parsing [39, 10].
-Direct retrieval involves taking the questions as a query and the knowledge triplets in the KG as candidates, using either sparse or dense retrieval techniques to identify several candidates most relevant to the query.
-Semantic parsing, on the other hand, transforms the question into an executable structured query statement (e.g., SPARQL) and executes the query in KGs.
-However, in KGs, some triplet by itself lacks rich semantic information due to the difference of KG schema, such as some entities are represented as machine identifiers (MIDs), and direct retrieval make it difficult to understand how relevant it really is to the user query.
-This becomes especially tricky in multi-hop question answering, where the intermediate triplets seem not directly related to the query can actually be crucial links to finding the right answer.
-Furthermore, semantic parsing faces the issues of non-executable or incorrectly executed generated queries [55].
+Therefore, several attempts have been conducted to integrate KGs with language models (Pan et al., 2024; Luo et al., 2024; Hu et al., 2023; Sun et al., 2023).
+Among them, knowledge graph question answering (KGQA) is a critical task to verify the effectiveness of incorporating the knowledge from KGs into reasoning models (He et al., 2021; Wang et al., 2023a; Yu et al., 2022).
+However, effectively integrating KGs into reasoning models for KGQA presents several challenges, including data sparsity, the complexity of query interpretation, and the need for advanced inference capabilities.
+More specifically, we focus on two main questions concerning the integration of KGs with reasoning models as follows:
+(I) How to retrieve specific knowledge from KGs to allow more precise reasoning?
+Existing solutions include direct retrieval (Sun et al., 2019; Jiang et al., 2022) and semantic parsing (Sun et al., 2020; Gu & Su, 2022).
+Direct retrieval utilize the query to find relevant knowledge triplets within the KGs.
+However, these triplets sometimes lack comprehensive semantic information due to variations in KG schema, such as entities being represented as machine identifiers (MIDs), rather than descriptive labels.
+This lack of semantic richness can obscure the relevance of the retrieved information with the user query, and becomes especially tricky in multi-hop question answering, where seemingly unrelated intermediate triplets may be pivotal in deriving the correct answer.
+On the other hand, semantic parsing aim to transform user queries into executable structured queries (e.g., SPARQL), which are then run on KGs.
+However, these methods often grapple with issues related to the non-executability or incorrect execution of the generated queries (Yu et al., 2022), which undermines the relaibility of these methods.
 (II) How to make the reasoning model understand and utilize the retrieved structured knowledge in KGs?
-Existing solutions include (1) finetune LMs on KG structure to generate relation path grounded by KGs.
-These reasoning paths are then used for LLMs to conduct faithful reasoning and generate interpretable results [29, 55]; (2) leverage the reasoning model to iteratively retrieve and reason on the subgraphs from KG [38, 9, 21] (i.e., at each step, the LLMs retrieve neighbors of the current nodes and then decide to answer the question or continue the next search step.) However, for the first approach, the reasoning steps generated by language models offer few guarantees about feasibility (i.e., the reasoning step might not exist in the KG), especially when multiple consecutive steps are combined into a reasoning path.
-We conduct the error analysis regarding this issue in Section 4.
-The latter solution face the challenge regarding determining the optimal stopping point for the exploration process.
-For example, ToG [38] prompts the LLMs to evaluate the adequacy of current reasoning paths for answer generation.
+Existing solutions include (1) finetuning LMs to generate relation paths grounded by KGs.
+These reasoning paths are then used for LLMs to conduct faithful reasoning and generate interpretable results (Luo et al., 2024; Yu et al., 2022); (2) leverage the reasoning model to iteratively retrieve and reason over subgraphs from the KG, deciding at each step to either answer the question, or to continue the searching step (Sun et al., 2023; Gu et al., 2023; Jiang et al., 2022).
+However, for the first approach, the reasoning steps generated by language models are not guaranteed to exist in the KG, especially when multiple consecutive steps are combined into a reasoning path.
+We conduct error analysis in Section 3 and show that only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG.
+The latter solution faces the challenge of determining the optimal stopping point for the exploration process.
+For example, ToG (Sun et al., 2023) prompts LLMs to assess the adequacy of current reasoning paths for answer generation.
 However, assessing the adequacy of reasoning paths is itself a complex task that demands a deep understanding of the domain.
-This lack of clarity can easily result in either premature stopping or excessive continuation, further complicating the decision-making process in LLMs.
-To address these two challenges, we propose a retrieval-exploration interactive method, FiDeLiS.
-It is designed to iteratively retrieve and reason on the subgraphs from KG, simulating a dynamic searching process.
-Specifically, we propose a Path-RAG module for retrieving useful intermediate knowledge from KGs.
-Compared to vanilla retriever which easily returns outputs that are not useful for finding the correct answer based on relevance to the query, we incorporate LLMs to generate an exhaustive list of keywords from the query to maximize coverage and ensure that no potential reasoning paths will be overlooked.
-We also integrate the topological connectivity of KGs to enhance system accuracy by considering entity relationships in the knowledge retrieval process.
-Furthermore, we propose to leverage deductive reasoning capabilities of LLMs [26, 18] as a better criterion for automatically directing the reasoning process step by step.
-Deductive reasoning serves as a clear indicator for ceasing further reasoning when necessary, which helps prevent errors in the reasoning chain and reduce unnecessary computation.
-Our framework enables us to employ beam search, a technique tailored for intermediate reasoning steps within a controllable searching space.
-We can guide the searching process by ensuring the logical soundness of each reasoning step to prevent errors throughout the reasoning process.
+This lack of clarity can easily result in premature stopping or excessive continuation, further complicating the decision-making process in LLMs.
+To address these two challenges, we propose a retrieval augmented reasoning method, FiDeLiS.
+It is designed to enhance KGQA by anchoring responses to structured, verifiable reasoning paths.
+FiDeLiS is composed of two major components, Path-RAG, which retrieves chain of entities and relations from KGs in a effective manner (discussed in Section 2.1), and Deductive-verification Beam Search (DVBS), which conducts deductive-reasoning-based beam search to generate multiple reasoning paths leading to final answers (discussed in Section 2.2).
+The Path-RAG module uses a keyword-enhanced retrieval mechanism that fetches relevant entities and relations from a vector-based index of KGs to ensure high-recall retrieval.
+Compared to a vanilla retriever which easily returns outputs that are not useful for finding the correct answer based on relevance to the query, we incorporate LLMs to generate an exhaustive list of keywords from the query to maximize coverage and ensure that no potential reasoning paths will be overlooked.
+Once these entities and relations are retrieved, the module constructs candidate reasoning steps which are then refined using a stepwise beam search discussed below.
+Next, DVBS leverages the deductive reasoning capabilities of LLMs (Ling et al., 2024; Huang & Chang, 2023) as a clear-defined criterion for automatically directing the beam search process step by step to create the complete reasoning paths.
+A key characteristic of DVBS is its integration of natural language planning (Song et al., 2023; Zhou et al., 2022) and beam search.
+This combination enhances the process of choosing the best reasoning paths by providing additional hints for decision making in LLMs.
+In addition, we redesign how reasoning paths are evaluated by converting the scoring process into a deductive reasoning task.
+This allows the LLM to validate the reasoning paths through deductive verification rather than traditional logit-based scoring.
+This proposed deductive verification serves as presie indicators for when to cease further reasoning, thus avoiding misleading reasoning chains and reducing the computational resources required.
 Overall, our main contributions include:
-  • We propose a retrieval-exploration interactive method that is specifically tailored to handle intermediate steps of reasoning grounded by KGs.
-  • We propose deductive verification as precise indicators for when to cease further reasoning, thus avoiding misleading the chains of reasoning and unnecessary computation.
-  • We integrate the reasoning process and topological connectivity of KGs into the knowledge retrieval process, which aids in recalling useful intermediate knowledge for reasoning and maintains efficiency as KGs scale up.
-  • Extensive experiments show that our method, as a training-free method with lower computational cost and better generality outperforms the existing strong baselines in three benchmarks.
+  • We propose a retrieval augmented reasoning method, which enhances knowledge graph question answering by anchoring responses to structured, verifiable reasoning paths grounded by KGs.
+  • We propose a step-wise keyword retrieval method that enhances the recall of relevant intermediate knowledge from KGs.
+This ensures that all the paths we create can be confidently linked back to KGs, ensuring that they are faithful and reliable.
+  • We propose deductive verification as precise indicators for when to cease further reasoning, thus avoiding misleading the chains of reasoning and reducing unnecessary computation required.
+  • Extensive experiments show that our method, as a training-free method with lower computational cost and better generality, outperforms existing strong baselines in three datasets.
 2 Method
+[…]
 Overview.
-FiDeLiS is designed as a retrieval-exploration interactive method as shown in Figure 1.
-It is designed to iteratively retrieve and reason on the subgraphs from KG, simulating a dynamic searching process through KG.
-FiDeLiS incorporates two modules Reasoning Path Retrieval-Augmented Generation (Path-RAG) and Deductive-Verification Guided Beam Search (DVBS), where Path-RAG is utilized to retrieve a highly pertinent set of nodes/relations from KGs to the user query and determine the reasoning step based on the relevance score discussed in Section 2.1.
-On the other hand, DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) for finding the most promising reasoning paths.
-At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path extension.
-To prevent either premature stopping or excessive continuation of reasoning path extension, DVBS enhances the clarity of termination signals using deductive reasoning discussed in Section 2.2.
+The framework of FiDeLiS is as shown in Figure 1.
+Given a query, it first retrieves the reasoning step candidates from KGs (as discussed in Section 2.1).
+To ensure high recall, our method first generates a list of keywords from the query.
+And then based on the generated keywords, it retrieves entities and relations from an index (where the entities and relations are also embedded by vector representation obtained from an LM).
+Next, following the retrieved entities, n-hop reasoning step candidates are created and scored based on semantics similarity and neighborhoods aggregation.
+Then, a deductive verification-based beam search (discussed in Section 2.2) is employed to generate the top-k reasoning paths leading to the answer.
+First, a natural language plan is generated from the query to provide more hints for LLMs decision making and then a beam search procedure is conducted to score each reasoning step.
+We redesign the scoring function by converting the process into a deductive reasoning task.
+This allows the LLM to validate connections between paths through simple verification rather than traditional logit-based scoring.
+This change helps prevent easily misleading reasoning chains and reduces the computational resources required.
+Figure 1: Overview of FiDeLiS.
+The goal of FiDeLiS is to extend the multiple-step reasoning paths $T$ grounded by KG.
+The figure shows given a query, Path-RAG first retrieve entities and relations that are relevant to the query, and construct the reasoning path candidates.
+Then DVBS constructs reasoning paths by iterative extending these path candidates using beam search.
+Specifically, at timestep $t$, DVBS leverage LLMs to choose the top-$k$ (here we set $k=2$ as an example) reasoning steps and decide whether to continue the next search step or cease the reasoning path extension based on deductive verification.
+At last, FiDeLiS return the top-$k$ reasoning paths as the final output leading to the answers of the query.
 2.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
-Path-RAG comprises three main steps: initialization, retrieval, reasoning step candidates construction, as depicted in Figure 1.
+Path-RAG aims to retrieve reasoning path candidates from KGs and comprises three main steps: initialization, retrieval, reasoning step candidates construction, as depicted in Figure 1.
 The implementation details of each step are elaborated in the following paragraphs.
 Initialization.
-We initiate the Path-RAG by generating node and edge embeddings using a pre-trained Language Model (LM).
-These embeddings are then stored in a nearest neighbor data structure.
-To elaborate, consider $e_{i}\in\mathcal{E},r_{i}\in\mathcal{R}$, where $\mathcal{E}$ refers to the entities and $\mathcal{R}$ represents the relations in the relational graph.
-We leverage the pre-trained LM (e.g., SentenceBert [33] and E5 [46]) to $e_{i}$ and $r_{i}$, yielding the representation $z(e^{i})$ and $z(r^{i})$:
-$z(e^{i})=\operatorname{LM}\left(e_{i}\right)\in\mathbb{R}^{d},\\ z(r^{i})=\operatorname{LM}\left(r_{i}\right)\in\mathbb{R}^{d}$ (1)
+We initiate the Path-RAG by generating embeddings for entities (nodes) and relations (edges) using a pre-trained language model (LM).
+We begin by extracting entities ($e_{i}$) and relations ($r_{i}$) from the KG, denoted as $\mathcal{E}$ and $\mathcal{R}$, respectively.
+Each entity and relation is then encoded using a pre-trained language model such as SentenceBert (Reimers & Gurevych, 2019) or E5 (Wang et al., 2022a), yielding dense vector representations:
+$z(e^{i})=\operatorname{LM}(e_{i})\in\mathbb{R}^{d},\ z(r^{i})=\operatorname{LM}(r_{i})\in\mathbb{R}^{d}$ (1)
 where $d$ denotes the dimension of the output vector.
+These embeddings are stored in a nearest neighbor data structure, which facilitates quick retrieval of similar entities or relations based on their vector representations.
 Retrieval.
-To retrieve the entities and relations that are relevant to the user query, the embeddings of them are then utilized to populate a nearest neighbor index.
-Specifically, given a query $q$, we leverage the LM to generate an exhaustive list of keywords or relation names $k_{i}\in\mathcal{K}$ that would be used to find the reasoning path from the knowledge graph to answer the query (as referred to prompt list in Section B).
-The generated list of keywords aims to maximize the coverage of potential reasoning steps and to enhance the subsequent decision-making process of LLMs by concentrating on a smaller set of relevant reasoning step candidates.
-$\mathcal{K}=\text{LM}(\text{prompt}_{p},q)$ (2)
-Then each $k_{i}^{m}\in\mathcal{K}$ is concatenated into a string $k_{i}$ and embedded into the same latent space by the LM to obtain $z(k)\in\mathbb{R}^{d}$, the retrieval process involves identifying the node or sequence of nodes that best aligns with the semantic content of the keywords as follow:
+To retrieve the entities and relations that are relevant to the user query, the embeddings of them are utilized to populate a nearest neighbor index.
+Specifically, for a given query $q$, we use LLM to generate an exhaustive list of keywords or relation names referred to as $\mathcal{K}$ (the prompt is given in Section F).
+These keywords are generated to maximize the coverage of potential reasoning steps essential for answering the query.
+The list is derived as $\mathcal{K}=\text{LM}(\text{prompt}_{p},q)$.
+Then each keyword $\in\mathcal{K}$ is concatenated into a string denoted as $K$ and then encoded into the same latent space by the same LM used in initialization step to obtain $z(K)=\text{LM}(K)\in\mathbb{R}^{d}$.
+To find the entities and relations that best match the keyword embeddings, we compute the cosine similarity $\cos(\cdot,\cdot)$ between $z(K)$ and the embeddings of entities ($z(e)$) and relations ($z(r)$) in the KG.
+We then select the top-$m$ elements that exhibit the highest similarity:
 $\mathcal{E}_{m}=\operatorname{argtopm}_{i\in\mathcal{E}}\cos\left(z(K),z(e)\right),\mathcal{R}_{m}=\operatorname{argtopm}_{i\in\mathcal{R}}\cos\left(z(K),z(r)\right)$ (2)
-where $z(e)$ and $z(r)$ are the embeddings of nodes $e$ and edges $r$, respectively.
-We use the cosine similarity function, $\cos(\cdot,\cdot)$, to measure the similarity between the keyword representation and the node/edge embeddings.
-The $\operatorname{argtopm}$ operation retrieves the top-m elements based on this similarity, providing a set of nodes $e_{k}$ and edges $r_{k}$ considered most relevant to the query.
-We denote the score of each entity and each relation as $S_{e}^{i}=\cos\left(z(k),z(e)\right)$ and $S_{r}^{i}=\cos\left(z(k),z(r)\right)$.
+Each entity and relation in the retrieved set is then assigned a score, reflective of its similarity to the keyword embedding as $S_{\text{ent}}(e)=\cos\left(z(K),z(e)\right)$ and $S_{\text{rel}}(r)=\cos\left(z(k),z(r)\right)$.
 Reasoning Step Candidates Construction.
-To construct reasoning steps candidates based on the finite set of nodes $\mathcal{E}_{m}$ and edges $\mathcal{R}_{m}$ relevant to the query and the relevance score $S_{e}^{i}$ and $S_{r}^{i}$, we denote the reasoning step can be scored and retrieved based on the following criteria as follows.
-The goal of this scoring function is to leverage the triplets to represent the connectivity of the KGs.
-It provides some hints on whether the current operation/decision would lead to a malicious operation several steps later by incorporating the next-hop neighbor’s information.
-$S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}\left(S_{r}^{j}+S_{e}^{j}\right)$ (4)
-It should be noted that the entity and relation mentioned in the equation should be logically connected in the corresponding KG.
-$S(p)$ denotes the score of the reasoning step; $S_{r}^{i}$, $S_{e}^{i}$ denote the relevance score of the next step relation and the corresponding neighbor; $S_{r}^{j}$, $S_{e}^{j}$ denote the relevance score of one further hop of next step.
-Parameter $\alpha$ is used to balance between immediate accuracy and long-term prospects within the graph.
-A higher $\alpha$ value encourages the model to consider less immediate but potentially vital outcomes, While a lower $\alpha$ makes it focus primarily on the immediate next steps without considerable regard for the long-term impacts.
-The detailed description can be found in Appendix Algorithm 1.
-Figure 1: Overview of FiDeLiS.
-The goal of FiDeLiS is to extend the multiple-step reasoning paths grounded by the external knowledge graph.
-We instantiate the $T$ step reasoning path extension as shown in the figure.
-First, Path-RAG is used to retrieve a highly pertinent set of nodes/relations from KGs to the user query and construct the reasoning step candidates.
-Then, the constructed reasoning steps candidates are forwarded to DVBS.
-DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates for finding the most promising reasoning step.
-Specifically, at timestep $T$, DVBS leverage the LLMs to choose the top-$k$ (here we set the $k=2$ for better clarification) reasoning steps and decide whether to continue the next search step or cease the reasoning path extension based on the deductive verification.
+After the initial retrieval process, which runs once at the start of the algorithm, we iteratively construct reasoning step candidates to extend the reasoning paths based based on entities $\mathcal{E}_{m}$ and relations $\mathcal{R}_{m}$ that are relevant to the given query.
+To guide the selection of these candidates, we propose another scoring function based on the derived relevance score $S_{\text{ent}}(e)$ and $S_{\text{rel}}(r)$.
+Our scoring function leverages triplets to represent the connectivity of KGs and assess whether extending a path in certain directions might lead to negative outcomes over multiple steps by incorporating next-hop neighbor’s information.
+$S((r,e))=S_{\text{rel}}(r)+S_{\text{ent}}(e)+\alpha\max_{\forall(r_{j},e_{j})\in N(e)}\left(S_{\text{rel}}(r_{j})+S_{\text{ent}}(e_{j})\right)$ (3)
+Here, $(r,e)$ refers to the reasoning step defined as a relation-entity pair.
+$S_{\text{rel}}(r_{j})$ and $S_{\text{ent}}(e_{j})$ are the similarity scores for relations and entities at the next step.
+The component $N(e)$ corresponds to entities reachable from $e$ within one hop.
+The factor $\alpha$ is used to balance short-term outcomes and long-term potential in reasoning paths: a higher $\alpha$ prioritizes paths with long-term benefits, even if they seem sub-optimal initially, while a lower $\alpha$ emphasizes immediate steps, potentially overlooking future impacts.
+Details of Path-RAG are provided in Algorithm 1.
 2.2 DVBS: Deductive-Verification Guided Beam Search
-DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) for finding the most promising reasoning paths.
-At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path extension.
+DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) to find the top-$k$ most promising reasoning paths that lead to the answer to the user query.
+At each timestep, we leverage the LLMs to choose the top-$k$ reasoning steps and decide whether to continue extending the reasoning paths or cease the extension process.
 It comprises three main steps: planning, beam-search, deductive-verification, as depicted in Figure 1.
@@
 Plan-and-Solve.
-Inspired by the recent works regarding planning capabilities of LLMs [35, 56], this step aims to provide more hints for subsequent LLM decision-making.
-Specifically, we prompt LLM to generate the planning steps for searching the most promising reasoning paths initially denoted as $p$.
-The detailed prompts can be found in Section B.
+Inspired by the recent works regarding planning capabilities of LLMs (Song et al., 2023; Zhou et al., 2022), we prompt LLM to generate the planning steps for answering the user query, denoted as $w$.
+This step aims to provide more hints for subsequent LLM decision making process.
+The detailed prompt can be found in Section F.
 Beam Search.
-In multi-step reasoning, a reasoning path of $T$ steps is sequentially generated through several time steps as $R=\left[s^{1},s^{2},\ldots,s^{T}\right]=s^{1:T}$ as shown in Figure 1.
-However, different from the typical text decoding process where each step consists of a single token, we follow the [50] and view a sequence of pair $(r,e)$ as a single step, where $r$ is the relation and $e$ is the corresponding entity.
-At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path.
-We present the beam search at each step is modeled as follows:
-$\mathcal{H}_{t}=\text{Top}_{K}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},p):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\right\}$ (5)
-where $\mathcal{H}_{t-1}$ denote the reasoning path at the previous time step $t-1$; $s^{t}$ refers to the current reasoning step; $\mathcal{S}$ refers to all possible next steps $s^{t}$ candidates produced from the Path-RAG; $\text{LM}(s^{t}|q,s^{1:t-1},p)$ refers to the language models prediction for the next step $s^{t}$ given the current sequence $s^{1:t-1}$, the query $q$, and the planning context $p$; $\oplus$ refers to appending $s^{t}$ to the current sequence $h$; $\text{Top}_{K}$ refers to selecting the top $K$ (denoted as beam width in the experiment setting) sets of reasoning steps.
+To enable multi-step reasoning, a reasoning path of $T$ steps is sequentially generated through several time steps as $\left[s^{1},s^{2},\ldots,s^{t},s^{T}\right]=s^{1:T}$, where $s^{t}$ represents the reasoning step at timestamp $t$.
+We constrain each reasoning step should within a set of candidates $\mathcal{S}^{t}$ to control the computation efficiency.
+We construct the reasoning step candidates $\mathcal{S}^{t}$ based on the scoring function defined in Eq.
+3 by selecting the top-$m$ reasoning steps with the highest scores.
+At each timestamp $t$, we leverage the LLMs to select the top-$k$ reasoning steps from the reasoning step candidates $\mathcal{S}^{t}$.
+The beam search process at timestamp $t$ is modeled as follows:
+$\mathcal{H}_{t}=\text{Top}_{k}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},w):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}^{t}\right\}$ (4)
+where $\mathcal{H}_{t-1}$ denote the reasoning path up to the previous timestamp $t-1$ and $s^{t}$ refers to the current reasoning step.
+$\mathcal{S}^{t}$ refers to the set of possible next-step candidates retrieved by the Path-RAG at timestamp $t$.
+$\text{LM}(s^{t}|q,s^{1:t-1},w)$ refers to the language models prediction for the next step $s^{t}$ given the previous sequence $s^{1:t-1}$, the query $q$, and the planning context $w$.
+The operator $\oplus$ appends $s^{t}$ to the current path $h$, and $\text{Top}_{k}$ selects the top-$k$ reasoning paths (with $k$ controlling the beam search width, where a larger $k$ typically yield better performance (see the analysis in Figure 2)).
 Deductive Verification.
-To mitigate the reasoning path extension terminates at the wrong point, we propose to leverage deductive reasoning capabilities of LLMs [26, 18] as a better criterion $C(x,s^{t},s^{1:t-1})\in\left\{0,1\right\}$ to automatically guide the decision making of LLMs in Equation 5.
-Specially, we first leverage LLMs to convert the user query to a declarative statement (as shown in Appendix G.2) and then use LLMs to judge whether the declarative statement of query $x^{\prime}$ can be deduced from current reasoning step $s^{t}$ and the previous reasoning steps $s^{1:t-1}$.
-We utilize the same backend LLM with prompt referred in Section B and define the deductive-verification criteria as follows:
-$C(q^{\prime},s^{t},s^{1:t-1})=\begin{cases}1&\text{if declarative statement ofquery }q^{\prime}\text{ can be deduced from }s^{t}\text{ and }s^{1:t-1},\\ 0&\text{otherwise.}\end{cases}$ (6)
-The overall DVBS can be modeled as follows:
-$\mathcal{H}_{t}=\text{Top}_{B}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},p):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\text{ and }C(q^{\prime},s^{t},s^{1:t-1})=1\right\}$ (7)
-This criterion is an essential component of our method for leveraging LLMs to iteratively validate each step of reasoning to make sure it follows logically and matches the original query.
-It acts as a clear sign to stop further reasoning when necessary, which helps avoid mistakes in the reasoning process and reduces unnecessary computations.
+To terminate the reasoning path extension process at the right point, we propose to leverage the deductive reasoning capabilities of LLMs (Ling et al., 2024; Huang & Chang, 2023) as a criterion $C(x,s^{t},s^{1:t-1})\in\left\{0,1\right\}$ to automatically guide the decision making of LLMs in Equation 4.
+Specially, we first leverage LLMs to convert the user query to a declarative statement $q^{\prime}$ (as shown in Appendix F.6) and then use LLMs to judge whether it can be deduced from current reasoning step $s^{t}$ and the previous reasoning steps $s^{1:t-1}$.
+We utilize the same backend LLM with prompt referred in Section F and define the deductive-verification criteria as follows:
+$C(q^{\prime},s^{t},s^{1:t-1})=\begin{cases}1,&\text{if }q^{\prime}\text{ can be deduced from }s^{t}\text{ and }s^{1:t-1},\\ 0,&\text{otherwise.}\end{cases}$ (5)
+The overall goal of DVBS model can be represented as:
+$\mathcal{H}_{t}=\text{Top}_{B}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},w):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\text{ and }C(q^{\prime},s^{t},s^{1:t-1})=1\right\}$ (6)
+This criterion is an essential component of our method for leveraging LLMs to iteratively validate each step of reasoning, ensuring that each step logically follows from the preceding steps and aligns with the original query.
+The function effectively signals when to terminate further reasoning, enhancing accuracy and minimizing unnecessary computational efforts.
 3 Experiments
-We aim to answer the following research questions during this section: RQ1: How does FiDeLiS compare with existing baselines in terms of accuracy and efficiency on KGQA tasks?
+During the experiment section, we aim to answer the following research questions: RQ1: How does FiDeLiS perform compare with existing baselines on KGQA tasks?
 RQ2: How does the design of each individual component contribute to the overall performance of FiDeLiS?
 RQ3: How robustness of FiDeLiS perform under varying conditions?
-3.1 Experiment Settings
-Datasets & Metrics.
-We adopt three benchmark KGQA datasets: WebQuestionSP (WebQSP) [54], Complex WebQuestions (CWQ) [40] and CR-LT-KGQA [11] in this work.
-We follow previous work [29] to use the same training and testing splits for fair comparison over WebQSP and CWQ.
-Both WebQSP and CWQ can be reasoned using Freebase KGs [^1] https://github.com/microsoft/FastRDFStore.
-To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we tested our methods on CR-LT-KGQA (further discussed in Section 5).
-We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples.
-Each of these can be reasoned based on the Wikidata KG [^2] https://www.wikidata.org/wiki/Wikidata:Main_Page.
-The statistics of the datasets are given in Table 8 and Table 9.
-To streamline the KGs, we utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ [29].
-Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA as well.
-We assess the performance of the method by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
-Baselines.
-(1) RoG [29]: embed structure knowledge graph from KGs into the underlying neural networks during the pretraining and fine-tuning process to generate the reasoning path and explanation.
-(2) ToG [38]: ask LLM to iteratively explore multiple possible reasoning paths on KGs until the LLM determines that the question can be answered based on the current reasoning paths.
-Our method FiDeLiS follows this paradigm.
-(3) NSM [12] utilizes the sequential model to mimic the multi-hop reasoning process.
-(4) KD-CoT [44] retrieves relevant knowledge from KGs to generate faithful reasoning paths for LLMs.
-(5) DeCAF [55] combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance on KGQA tasks.
-The details of each baseline are further discussed in Section G.1.
-Implementation Details. We set the default beam width as $4$ and depth as $4$ without specific annotation. We set the $\alpha$ in Eq 4 as $0.3$ to ensure reproducibility.
-Backboned LLMs.
-We assess our approach on closed- and open-source LLMs.
-For closed-source LLMs, we choose GPT-4-turbo, GPT-3.5-turbo [^3] We use the recent gpt-4-turbo model that is released in 2024-04-09 from https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 and gpt-3.5-turbo model notated as “gpt-3.5-turbo-0125” from https://platform.openai.com/docs/models/gpt-3-5-turbo.
-to report and compare the results on all datasets.
-We use Llama-2-13B [42] and Mistral-7B [19] as our open-source LLMs to conduct cost–performance analysis on different datasets.
-The experiment results of open-source models can be found in Table 10.
-3.2 RQ1: KGQA Performance Comparison
-Comparison between baseline methods.
-Table 1 clearly showcases the comparative effectiveness of FiDeLiS, against different baselines.
-It shows that FiDeLiS outperforms all baselines integrated with gpt-4-turbo, even for strong baselines being trained, such as DeCAF [55] and RoG [29].
-Across all prompting-based methods, gpt-4-turbo achieves higher performance compared to gpt-3.5-turbo, especially on the CWQ dataset.
-It indicates that gpt-4-turbo has a better understanding and processing capability for complex queries.
-The CR-LT dataset seems more demanding than WebQSP and CWQ, as implied by the consistently lower accuracy scores.
-However, FiDeLiS shows consistent enhancements compared to other baselines, demonstrating its capability to handle long-tail entities by referring to the knowledge from KGs and processing more complex queries effectively.
+RQ4: How about the efficiency of FiDeLiS?
+All experiment settings are detailed in Appendix D due to page constraints.
 Table 1: Comparison of FiDeLiS with baseline methods and different backbone LLMs.
@@
 CoT 72.11 65.37 53.51 44.76 45.42
-Finetuning - LLM + KG NSM [12] 74.31 - 53.92 - -
-CBR-KBQA [6] - - 67.14 - -
-DeCAF [55] 82.1 - 70.42 - -
-KD-CoT [44] 73.7 50.2 50.5 - -
-RoG [29] 83.15 69.81 61.39 56.17 60.32
-Prompting - LLM + KG gpt-3.5-turbo ToG [38] 75.13 72.32 57.59 56.96 62.48
+Finetuning - LLM + KG NSM (He et al., 2021) 74.31 - 53.92 - -
+CBR-KBQA (Das et al., 2021) - - 67.14 - -
+DeCAF (Yu et al., 2022) 82.1 - 70.42 - -
+KD-CoT (Wang et al., 2023a) 73.7 50.2 50.5 - -
+RoG (Luo et al., 2024) 83.15 69.81 61.39 56.17 60.32
+Prompting - LLM + KG gpt-3.5-turbo ToG (Sun et al., 2023) 75.13 72.32 57.59 56.96 62.48
 FiDeLiS 79.32 76.78 63.12 61.78 67.34
-Prompting - LLM + KG gpt-4-turbo ToG [38] 81.84 75.97 68.51 60.20 67.24
+Prompting - LLM + KG gpt-4-turbo ToG (Sun et al., 2023) 81.84 75.97 68.51 60.20 67.24
 FiDeLiS 84.39 78.32 71.47 64.32 72.12
+3.1 RQ1: KGQA Performance Comparison
+Table 1 clearly showcases the comparative effectiveness of FiDeLiS, against different baselines.
+It shows that FiDeLiS outperforms all baselines using gpt-4-turbo model, even for strong baselines being fine-tuned, such as DeCAF (Yu et al., 2022) and RoG (Luo et al., 2024).
+Across all prompting-based methods, gpt-4-turbo achieves higher performance compared to gpt-3.5-turbo, especially on the CWQ dataset.
+It indicates that gpt-4-turbo has a better understanding and processing capability for complex queries.
+The CR-LT dataset seems more challenging compared with WebQSP and CWQ, as implied by the consistently lower accuracy scores.
+However, FiDeLiS shows consistent enhancements compared to other baselines, demonstrating its capability to handle long-tail entities by referring to the knowledge from KGs and processing more complex queries effectively.
 3.2 RQ2: Ablation Study of FiDeLiS.
-Table 2 provides a detailed view of how removing specific components from FiDeLiS impacts its performance.
-The results emphasize that some features, such as Path-RAG and beam search in DVBS, are particularly significant, especially for tackling complex reasoning tasks on datasets like CWQ.
-When Path-RAG and beam-search within DVBS are excluded, we observe substantial declines in performance.
-This underlines their importance in our model’s ability to handle intricate reasoning effectively.
-Furthermore, the integration of the reasoning and planning steps within DVBS plays a critical role in maintaining robust performance across varied datasets.
-One aspect highlighted by our study is the role of deductive verification during the beam search phase in DVBS.
-This feature enhances the model’s ability to delineate and reinforce decision boundaries, thus safeguarding against reasoning errors.
-Table 2: Ablation Studies of FiDeLiS, we test the ablation study using model gpt-3.5-turbo-0125.
-DVBS refers to “Deductive-verification guided Beam Search’.
-“$\Delta$” refers to the performance gap between each component and the entire method.
-The top-3 performance gaps on each dataset are highlighted in green, with darker shades indicating significant performance gap.
-- Components WebQSP CWQ CR-LT
+Table 2 demonstrates the ablation study of FiDeLiS using the gpt-3.5-turbo-0125 model across the WebQSP, CWQ, and CR-LT datasets.
+It highlights the critical importance of both the Path-RAG and DVBS components.
+Removing Path-RAG, whether by employing a vanilla retriever or the think-on-graph (ToG) approach (Sun et al., 2023), results in substantial performance declines, particularly a 6.97% drop in Hits@1 on WebQSP and a 6.01% decrease on CWQ, underscoring the necessity of an effective retrieval mechanism.
+Similarly, ablating parts of DVBS, especially the beam search component, leads to significant reductions, with an 18.97% decrease in Hits@1 on WebQSP and a 13.34% drop on CWQ, indicating that beam search is vital for maintaining high accuracy.
+Other DVBS subcomponents, such as the deductive verifier and planning, also contribute notably to performance, albeit to a lesser degree.
+Overall, the findings demonstrate that the integrated Path-RAG and DVBS frameworks are essential for FiDeLiS’s robust performance, with their removal causing marked decreases in accuracy and Hits@1 scores across all evaluated datasets.
+Table 2: Ablation Studies of FiDeLiS using model gpt-3.5-turbo-0125.
+$\Delta$ refers to the performance gap between each component and the entire method.
+The three largest performance gaps on each dataset are highlighted in green, with darker shades denoting more significant differences.
+Ablation Setting Components WebQSP CWQ CR-LT
 Hits@1 (%) $\Delta$ Hits@1 (%) $\Delta$ Acc (%) $\Delta$
-- FiDeLiS 79.32 0.00 63.12 0.00 67.34 0.00
+No ablation FiDeLiS 79.32 0.00 63.12 0.00 67.34 0.00
 w/o Path-RAG using vanilla retriever 72.35 6.97 57.11 6.01 59.78 7.56
@@
 It shows that the integration of the Openai-Embedding Model significantly enhances the performance across all datasets, indicating its superior capability in understanding and processing natural language queries, which leads to more effective retrieval.
-For instance, the Openai-Embedding Model increases the Hits@1 by 13.04% in the WebQSP dataset and 14.73% in the CWQ dataset compared to the base BM25 [34] model.
+For instance, the Openai-Embedding Model increases the Hits@1 by 13.04% in the WebQSP dataset and 14.73% in the CWQ dataset compared to the base BM25 (Robertson et al., 2009) model.
 Furthermore, this improvement is consistent and highlights the feasibility and potential of pairing more powerful language models with structured reasoning frameworks to address complex information retrieval and question answering challenges.
@@
 Hits@1 (%) Hits@1 (%) Acc (%)
-Vanilla Retriever w/ BM25 [34] 58.31 48.39 50.73
-w/ SentenceBert [33] 62.74 50.14 51.80
-w/ E5 [46]) 68.42 52.84 54.31
+Vanilla Retriever w/ BM25 (Robertson et al., 2009) 58.31 48.39 50.73
+w/ SentenceBert (Reimers & Gurevych, 2019) 62.74 50.14 51.80
+w/ E5 (Wang et al., 2022a) 68.42 52.84 54.31
 w/ Openai-Embedding-Model 72.35 57.11 59.78
-Path-RAG w/ BM25 [34] 70.34 56.11 58.77
-w/ SentenceBert [33] 73.45 58.41 60.45
-w/ E5 [46]) 77.93 62.74 65.23
+Path-RAG w/ BM25 (Robertson et al., 2009) 70.34 56.11 58.77
+w/ SentenceBert (Reimers & Gurevych, 2019) 73.45 58.41 60.45
+w/ E5 (Wang et al., 2022a) 77.93 62.74 65.23
 w/ Openai-Embedding-Model 79.32 63.12 67.34
 Comparison using different searching widths and depths.
-To explore the influence of the search width and depth on FiDeLiS’s performance, we conduct experiments under settings with depths ranging from 1 to 4 and widths from 1 to 4.
-As shown in Figure 2, FiDeLiS’s performance improves with the search depth and width and when the searching depth is larger than 3, FiDeLiS’s performance start to decrease on both WebQSP and CWQ.
-This is because only a small part of questions in these two datasets have reasoning depths greater than 3, and also FiDeLiS implements stricter rules for terminating searches using deductive reasoning.
-However, for searching width, we find that the performance is consistent and implies the FiDeLiS’s performance could potentially be improved with the increment of the exploration breadth.
-Considering the computational cost (which increases linearly with the depth), we set both the default beam width and depth as 4 without specific annotations in all the experiments.
-Figure 2: Performance of FiDeLiS with different beam search widths and depths.
+To investigate how the search width and depth affect the performance of FiDeLiS, we have performed experiments with settings ranging from depths of 1 to 4 and widths of 1 to 4.
+The results, presented in Figure 2, show that performance generally improves as search depth and width increase.
+However, performance begins to decline when the search depth exceeds 3 for both the WebQSP and CWQ datasets.
+This decline is attributed to the fact that only a small fraction of questions in these datasets require reasoning at depths greater than 3.
+Additionally, FiDeLiS utilizes stricter rules for ending the search process, which can preemptively terminate misguided searches.
+Contrastingly, increasing the search width consistently shows potential for improved performance, suggesting benefits from broader exploration.
+However, considering the computational costs, which rise linearly with depth, we have chosen to set both the default beam width and depth to 4 for all experiments unless stated otherwise.
+This default setting aims to balance performance gains with computational efficiency.
+(a) RD over CWQ
+(b) RD over WebQSP
+(c) BW over CWQ
+(d) BW over WebQSP
+Figure 2: Performance (Hits@1) of FiDeLiS with different beam search widths (BW) and reasoning depths (RD) over WebQSP and CWQ.
 Comparison between Path-RAG and Vanilla Retriever.
-To verify the validity of the Path-RAG compared to vanilla retrievers.
-We conduct the experiments as shown in Figure 3.
-For the vanilla retriever, we concatenate each entity with its corresponding relation to form a reasoning step.
-We then calculated the cosine similarity between the embeddings of each reasoning step and the query embeddings to select candidate paths.
-Figure 3: Comparison between Path-RAG and Vanilla Retriever.
+To verify the robustness of the Path-RAG compared to vanilla retrievers.
+We conduct the experiments as shown in Figure 3(a) and 3(b).
 We calculate the coverage ratio (CR) of the retrieved reasoning paths and the ground-truth reasoning paths as follows: $\text{CR}=\left(\frac{N_{\text{retrieved}}\cap N_{\text{ground-truth}}}{N_{\text{ground-truth}}}\right)\times 100\%$, where $N_{\text{retrieved}}$ is the number of retrieved reasoning paths candidates, and $N_{\text{ground-truth}}$ is the total number of ground-truth reasoning paths.
+We set the vanilla retriever as the baseline, specifically, we concatenate each entity with its corresponding relation to form a reasoning step.
+We then calculated the cosine similarity between the embeddings of each reasoning steps and the query embeddings to select candidate paths.
 We observe that compared to the vanilla retriever, Path-RAG achieves a higher CR value and shows better alignment with the ground-truth paths.
@@
 This feature is critical in subsequent LLM to focus on the information of interest and arrive at a correct and reasonable answer.
-Table 4: Average depths of the returned reasoning paths. The average depth of ground-truth reasoning paths of WebQSP is 2.3, for CWQ is 3.2, and for CR-LT is 4.7.
+(a) Coverage Ratio over CWQ
+(b) Coverage Ratio over WebQSP
+(c) Error Analysis over WebQSP
+Figure 3: (a)-(b): Empirical study regarding the coverage ratio of the retrieved reasoning paths.
+(c): Error analysis of the validation of whole paths generated from RoG (Luo et al., 2024) over WebQSP.
+Error analysis regarding whole path generation.
+To verify the faithfulness of our stepwise method, we conduct the error analysis regarding the whole reasoning path generation.
+Specifically, we conduct an analysis of the validity of the whole reasoning path generation using the baseline methods RoG (Luo et al., 2024).
+The result is shown in Figure 3(c).
+The definition of the metric validity ratio (VR) is the ratio of reasoning steps that existed in the knowledge graph to the total number of the reasoning steps in the output reasoning path: $\text{VR}=\left(\frac{N_{\text{valid-steps}}}{N_{\text{all-steps}}}\right)\times 100\%$, where $N_{\text{valid-steps}}$ is the number of reasoning steps that existed in the knowledge graph, and $N_{\text{all-steps}}$ is the number of all the reasoning steps in the reasoning path.
+It show that only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG.
+This illustrates that the reasoning steps generated by language models offer few guarantees about feasibility especially when multiple consecutive steps are combined into a reasoning path.
+Table 4: Average depths of the returned reasoning paths. GT refers to ground-truth reasoning paths.
 Method WebQSP CWQ CR-LT
+GT 2.3 3.2 4.7
 ToG 3.1 4.1 5.2
@@
 It shows that FiDeLiS consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to ToG.
-This implies that our method may offer more precise termination signals and potentially more accurate reasoning paths.
-Error Analysis regarding Whole Path Generation.
-Figure 4: Validation of whole path generation from RoG [29] over WebQSP.
-To verify the faithfulness of our stepwise method, we conduct the error analysis regarding the whole reasoning path generation.
-Specifically, we conduct an analysis of the validity of the whole reasoning path generation using the baseline methods RoG [29].
-The result is shown in Figure 4.
-The definition of the metric validity ratio (VR) is the ratio of reasoning steps that existed in the knowledge graph to the total number of the reasoning steps in the output reasoning path: $\text{VR}=\left(\frac{N_{\text{valid-steps}}}{N_{\text{all-steps}}}\right)\times 100\%$, where $N_{\text{valid-steps}}$ is the number of reasoning steps that existed in the knowledge graph, and $N_{\text{all-steps}}$ is the number of all the reasoning steps in the reasoning path.
-We note that 33% of the generated reasoning paths are invalid, 18% exhibit formatting errors, and 15% include relations that do not exist in the knowledge graph.
-This illustrates the limitations of embedding knowledge graphs into base LLM models, particularly in controlling the generation process.
+This implies that our method may offer more precise termination signals and potentially more accurate reasoning paths compared to baselines.
 Case Study.
@@
 RoG location.country.form_of_government, location.location.containedby location.country.form_of_government’ location.location.containedby $\to$ government.form_of_government.countries The results are Islamic republic.
-ToG Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic $\to$ government.form_of_government.countries $\to$ Iran Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy $\to$ government.form_of_government.countries $\to$ Iran Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state $\to$ government.form_of_government.countries $\to$ Iran The results are Iran.
-FiDeLiS Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state The results are Theocracy, Unitary state, Islamic republic.
+ToG Path-1: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic $\to$ government.form_of_government.countries $\to$ Iran Path-2: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy $\to$ government.form_of_government.countries $\to$ Iran Path-3: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state $\to$ government.form_of_government.countries $\to$ Iran Based on the reasoning paths, the result is Iran.
+FiDeLiS Path-1: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic Path-2: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy Path-3: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state Based on the reasoning paths, the results are Theocracy, Unitary state, Islamic republic.
+3.4 RQ4: Efficiency of FiDeLiS
+To investigate the runtime efficiency and cost efficiency of FiDeLiS, we present a comparison regarding the average runtime, average token usage, average times of LLM calling per question in Table 6.
+We found that (1) our method shows superior efficiency compared to the ToG (which is also training-free), by reducing approximately 1.7x runtime costs.
+(2) Path-Rag component is critical in enhancing both the accuracy and efficiency of the model.
+Its ability to constrain potential path candidates effectively reduces unnecessary computational overhead, leading to quicker and more accurate results.
+Table 6: Runtime efficiency of FiDeLiS per question.
+Dataset Method Hits@1 (%) Avg Runtime (s) Avg Token Usage Avg LLM calling
+WebQSP FiDeLiS (ours) 79.32 43.83 2,452 10.7
+   w/o Path-RAG using vanilla retriever 72.35 48.37 2,873 10.7
+   w/o Path-RAG using ToG 75.11 74.26 6,437 10.7
+FiDeLiS (ours) - GPT-4o 81.17 37.82 2,452 10.7
+FiDeLiS (ours) - GPT-4o-mini 76.48 24.31 2,452 10.7
+CWQ FiDeLiS (ours) 63.12 74.59 2,741 15.2
+   w/o Path-RAG using vanilla retriever 57.11 78.41 3,093 15.2
+   w/o Path-RAG using ToG 59.47 132.59 5,372 15.2
+FiDeLiS (ours) - GPT-4o 65.33 50.12 2,741 15.2
+FiDeLiS (ours) - GPT-4o-mini 58.34 42.54 2,741 15.2
+To address concern regarding our method’s potential application in real-time scenarios, we also tested our method using faster and more advanced LLMs.
+Table 6 shows that our method could be further accelerated with newer, faster models like GPT-4o or GPT-4-mini.
+The potential of the ongoing advancements in LLMs are expected to further enhance the scalability and efficiency of FiDeLiS, making it a practical development in challenging environments.
+More detailed analysis of bottleneck of computation of FiDeLiS can be further found in Appendix C.
 4 Related Work
 KG-enhanced LLM.
-KGs have advantages in dynamic, explicit, and structured knowledge representation and techniques combining LLMs with KGs have been studied [32].
-Early studies [29, 55] embed structured knowledge from KGs into the underlying neural networks during the pretraining or fine-tuning process.
+KGs have advantages in dynamic, explicit, and structured knowledge representation and techniques combining LLMs with KGs have been studied (Pan et al., 2024).
+Early studies (Luo et al., 2024; Yu et al., 2022) embed structured knowledge from KGs into the underlying neural networks during the pretraining or fine-tuning process.
 However, the reasoning steps generated by LMs are observed to be prone to errors, which could be non-existent in the knowledge graph, and can subsequently lead to incorrect reasoning during inference.
-Also, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating [17].
-Another solutions are to keep the reasoning model as an agent to explore the external structure knowledge source [38, 9, 20, 44] ToG [38] directly employs LLMs to output scores for candidate selections from the KG.
+Also, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating (Hu et al., 2023).
+Another solutions are to keep the reasoning model as an agent to explore the external structure knowledge source (Sun et al., 2023; Gu et al., 2023; Jiang et al., 2023b; Wang et al., 2023a) ToG (Sun et al., 2023) directly employs LLMs to output scores for candidate selections from the KG.
 This model operates at the relation-entity level of the KG, aiming to identify relevant triples that aid the LLM in making accurate and responsible final answer predictions.
 Knowledge Graph-based Question Answering.
-Conventional embedding-based methods represent the entities and relations in embedding space and design special model architectures (e.g., Key-Value memory networks, sequential models, and graph neural networks) to reason answers [30, 12, 52].
-To integrate LLMs for KGQA, retrieval-augmented methods aim to retrieve the relative facts from the KGs to improve the reasoning performance [25, 22].
-Recently, UniKGQA [21] which unifies the graph retrieval and reasoning process into a single model with LLMs, achieves STOA performance.
-Semantic parsing methods convert the question into a structural query (e.g., SPARQL) by LLMs, which can be executed by a query engine to reason the answers on KGs [39, 24].
+To integrate LLMs for KGQA, retrieval-augmented methods aim to retrieve the relative facts from the KGs to improve the reasoning performance (Li et al., 2023; Karpukhin et al., 2020).
+Recently, UniKGQA (Jiang et al., 2022) which unifies the graph retrieval and reasoning process into a single model with LLMs, achieves STOA performance.
+Semantic parsing methods convert the question into a structural query (e.g., SPARQL) by LLMs, which can be executed by a query engine to reason the answers on KGs (Sun et al., 2020; Lan & Jiang, 2020).
 However, these methods heavily rely on the quality of generated queries.
 If the query is not executable, no answers will be generated.
-DECAF [55] combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance.
+DECAF (Yu et al., 2022) combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance.
 5 Conclusion
@@
 We believe this study will significantly benefit the integration of LLMs and KGs, or serve as an auxiliary tool to enhance the interpretability and factual reliability of LLM outputs.
-Appendix A Defintions
-A.1 Defining a Valid Reasoning Trace/Path
-Definition 1. A reasoning step is a pair $(r,e)$, where $r$ is the relation and $e$ is the corresponding entity.
-Definition 2.
-A reasoning path $\mathcal{P}$ is a pair $(s,\mathcal{T})$, where $s$ is the starting entity for the reasoning path, and $\mathcal{T}$ is a sequence of reasoning steps $\mathcal{T}=\left\{t_{0},t_{1},\ldots,t_{n}\right\}$ and $t_{k}=(r_{k},e_{k})$.
-The $t_{k}\in\mathcal{R}$ denotes the $k$-th reasoning step in the path and $n$ denotes the length of the path.
-[…]
-A.2 Knowledge Graph-based Question Answering (KGQA)
+Appendix A Definitions
+A.1 Knowledge Graph-based Question Answering (KGQA)
 In this work, we mainly focus on the question answering task based on the knowledge graph.
@@
 Given a natural language question $q$ and a $\mathrm{KG}$ $\mathcal{G}$, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ based on knowledge from $\mathcal{G}$, i.e., $a=f(q,\mathcal{G})$.
-Following previous works [29, 28], we assume the entities $e_{q}\in\mathcal{T}_{q}$ mentioned in $q$ and answers $a\in\mathcal{A}_{q}$ are labeled and linked to the corresponding entities in $\mathcal{G}$, i.e., $\mathcal{T}_{q},\mathcal{A}_{q}\subseteq\mathcal{E}$.
-Appendix B Prompt List
-In this section, we show all the prompts that need to be used in the main experiments. The In-Context Few-shot refers to the few-shot examples we used for in-context learning.
-B.1 Plan-and-solve
-You are a helpful assistant designed to output JSON that aids in navigating a knowledge graph to answer a provided question. The response should include the following keys:
-(1) ’keywords’: an exhaustive list of keywords or relation names that you would use to find the reasoning path from the knowledge graph to answer the question.
-Aim for maximum coverage to ensure no potential reasoning paths will be overlooked;
-(2) ’planning_steps’: a list of detailed steps required to trace the reasoning path with. Each step should be a string instead of a dict.
-(3) ’declarative_statement’: a string of declarative statement that can be transformed from the given query, For example, convert the question ’What do Jamaican people speak?’ into the statement ’Jamaican people speak *placeholder*.’ leave the *placeholder* unchanged; Ensure the JSON object clearly separates these components.
-In-Context Few-shot
-Q: {Query}
-A:
-B.2 Deductive-verification
-You are asked to verify whether the reasoning step follows deductively from the question and the current reasoning path in a deductive manner. If yes return yes, if no, return no".
-In-Context Few-shot
-Whether the conclusion ’{declarative_statement}’ can be deduced from ’{parsed_reasoning_path}’, if yes, return yes, if no, return no.
-A:
-B.3 Adequacy-verification
-You are asked to verify whether it’s sufficient for you to answer the question with the following reasoning path.
-For each reasoning path, respond with ’Yes’ if it is sufficient, and ’No’ if it is not.
-Your response should be either ’Yes’ or ’No’.
-In-Context Few-shot
-Whether the reasoning path ’{reasoning_path}’ be sufficient to answer the query ‘{Query}’, if yes, return yes, if no, return no.
-A:
-B.4 Beam Search
-Given a question and the starting entity from a knowledge graph, you are asked to retrieve reasoning paths from the given reasoning paths that are useful for answering the question.
-In-Context Few-shot
-Considering the planning context {plan_context} and the given question {Query}, you are asked to choose the best {beam_width} reasoning paths from the following candidates with the highest probability to lead to a useful reasoning path for answering the question.
-{reasoning_paths}.
-Only return the index of the {beam_width} selected reasoning paths in a list.
-A:
-B.5 Reasoning
-Given a question and the associated retrieved reasoning path from a knowledge graph, you are asked to answer the following question based on the reasoning path and your knowledge.
-Only return the answer to the question.
-In-Context Few-shot
-Question: {Query}
-Reasoning path: {reasoning_path}
-Only return the answer to the question.
-A:
-Appendix C Potential Impacts and Limitations
-The proposed retrieval-exploration interactive method, FiDeLiS, holds the potential to significantly enhance the performance of LLMs by addressing the issue of hallucinations and enabling deep, responsible reasoning.
-By integrating KGs with LLMs, the method offers more accurate knowledge retrieval, leveraging deductive reasoning capabilities to guide the reasoning process and avoid logical fallacies.
-The stepwise and generalizable approach, along with the use of deductive verification as a stop criterion, may reduce unnecessary computation and misleading reasoning chains.
-Additionally, the training-free nature of this method combined with its lower computational cost could be easily acted as a plugin-in for other methods/frameworks.
-However, there are limitations to consider.
-As the method is still in its early stages of development, there may be unforeseen challenges or drawbacks when applied to real-world scenarios.
-The reliance on external knowledge graphs means that the quality and coverage of these resources can impact the overall effectiveness of the method.
-There may also be computational constraints when working with very large or complex graphs.
-Additionally, while the method has shown promise in benchmarks, its performance in more diverse or specialized tasks is yet to be evaluated.
-Despite these challenges, the research presents a promising direction for improving the reasoning capabilities of LLMs.
-Appendix D Efficiency of FiDeLiS
+Following previous works Luo et al.
+(2024; 2023), we assume the entities $e_{q}\in\mathcal{T}_{q}$ mentioned in $q$ and answers $a\in\mathcal{A}_{q}$ are labeled and linked to the corresponding entities in $\mathcal{G}$, i.e., $\mathcal{T}_{q},\mathcal{A}_{q}\subseteq\mathcal{E}$.
+Appendix B Potential Impacts and Limitations
+The proposed method holds the potential to significantly enhance the performance of large language models (LLMs) by tackling the issue of hallucinations, thereby fostering deep, responsible reasoning.
+By integrating KGs with LLMs, the approach not only facilitates more accurate knowledge retrieval but also leverages deductive reasoning capabilities to steer the reasoning process and circumvent logical fallacies.
+The method is characterized by its stepwise, generalizable approach, and the use of deductive verification as a stopping criterion, which together may reduce superfluous computation and curb misleading reasoning chains.
+Moreover, due to its training-free nature and lower computational demands, this method could seamlessly serve as a plug-in for other existing frameworks.
+However, it’s important to recognize certain limitations.
+Since the method is still in its early stages of development, it might encounter unanticipated challenges or drawbacks in real-world applications.
+Its dependency on external KGs means that the quality and comprehensiveness of these resources can affect its overall effectiveness.
+Computational constraints might also arise when dealing with very large or intricate graphs.
+Furthermore, although the method has demonstrated promise in benchmark tests, its performance across more diverse or specialized tasks remains untested.
+Appendix C Bottleneck of Beam Search Efficiency
 The bottleneck of computation is the beam search process, which contributes to $N*D$ times LLM calling, where $D$ is the depth (or equivalently length) of the reasoning path, and $N$ is the width of the beam-search (how many paths are remained in the pool in each iteration).
 Specifically, we need to call $ND+D+C$ times LLM for each sample question, where $C$ is a constant (equals to $1$ if there is no error occurs when calling the API).
-[38] illustrates that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling.
+Sun et al.
+(2023) illustrates that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling.
 In this way, we can reduce the number of LLM calling from $ND+D+C$ to $D+C$.
-However, [38] shows that this optimization sacrifices the accuracy due to the weaker scoring model in decision making.
-Table 6: Average number of LLM calls per question.
-Dataset WebQSP CWQ CR-LT
-Average LLM Calls 10.7 15.2 17.3
+However, Sun et al.
+(2023) shows that this optimization sacrifices the accuracy due to the weaker scoring model in decision making.
 We noted that $ND+D+C$ is the maximal computational complexity.
@@
 We can also see that this average number gets even smaller for dataset covering a lot of single-hop reasoning questions, such as WebQSP.
-Table 7: Average token usage of LLM calls per question.
-Dataset WebQSP CWQ CR-LT
-Average Token Usage 2,452 2,741 1,824
-Cost efficiency.
-Another benefit of FiDeLiS is the cost efficiency.
-We control the token usage for each LLM calling through Path-RAG module.
-Specifically, Path-RAG only retrieves the top-n reasoning steps as the candidates for each beam search decision making, before making calls to the LLM.
-This approach ensures token expenditure without compromising the model’s performance, especially in some cases where hundreds of neighbors exist around a entity node.
-Table 7 shows the average token usage per question needed by FiDeLiS on different datasets.
-It can be seen that in three KGQA datasets, the numbers of token usages are ranging from 1,824 to 2,751, which could significantly reduce the cost for running the method.
-Appendix E Variants of Path-RAG
-Variant-1: To enhance the scoring function for reasoning in a knowledge graph and ensure each entity and relation are consistently linked, we can modify the scoring function to more explicitly account for the connectivity and interdependences between entities and relations throughout the path.
-This can be done by incorporating an accumulative sum that keeps track of previous scores and emphasizes path completeness.
-The revised formula introduces variables for the accumulated scores of entities and relations, which will help ensure all links in the path are considered:
-$S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}\left(S_{r}^{j}+S_{e}^{j}\right)+\beta\sum_{k=0}^{i-1}\left(S_{r}^{k}+S_{e}^{k}\right)$ (8)
-where: (1) $S_{r}^{i}$ and $S_{e}^{i}$ denote the relevance score of the current step relation and the corresponding neighboring entity, respectively.
-(2) $S_{r}^{j}$ and $S_{e}^{j}$ denote the relevance score of the 1-hop relation and the corresponding neighboring entity from node $N_{i}$.
-(3) The new term, $\sum_{k=0}^{i-1}\left(S_{r}^{k}+S_{e}^{k}\right)$, accumulates the scores from all previous steps from the start up to the step before the current, enhancing the influence of the entire path on the decision of the next step.
-(4) $\alpha$ and $\beta$ are weight factors that control the influence of the maximum score of neighboring nodes and the accumulated score to adjust the balance in prioritizing extended paths versus immediate gains.
-This adjustment helps to: firstly, reward paths that maintain strong links throughout, discouraging breaks or weakly connected steps in the reasoning process; secondly, enhance the depth of consideration in the decision-making process, looking both at immediate and long-term implications of each step.
-Variant-2: We also consider to incorporate the randomness into the scoring function to balance the quality-diversity trade-off in searching for better reasoning paths as follows:
-$S(p)=\lambda_{1}\left(S_{r}^{i}+S_{e}^{i}+\alpha\sum_{j\in N_{i}}(S_{r}^{j}+S_{e}^{j})\right)+\lambda_{2}\sum_{k=0}^{i-1}\left(D(p,P_{k})\right)$ (9)
-where $D(p,P_{\text{sel}})$ refers to the distance between the selected path with the grounded reasoning paths.
-while $\lambda_{1}$ and $\lambda_{2}$ are weighting factors that balance the importance of internal quality metrics against diversity metrics.
-Appendix F Additional Related Works: Reasoning with LLM Prompting
-With the development of LLMs, many creative ways to leverage LLMs are proposed: End-to-End, Chain-of-Thought, and Semantic Parsing/Code Generation.
-End-to-End methods [4, 15] aims to leverage LLMs to generate final answers directly, often done by providing a task description and/or a few examples for in-context learning.
-While it offers convenience, however, it suffers from the un-interpretability of the generation, and the lack of explicit steps and reliance solely on the LLMs’ training data may result in un-robustness and exhibit sensitivity to slight input variations.
-Chain-of-Thought methods [48, 23, 8] emphasize breaking down complex reasoning into a series of intermediate steps to support complex reasoning.
-However, CoT suffers from unreliability and uncontrollability since the generated reasoning steps may not always align with the intended logical progression or may produce incorrect or inconsistent answers when multiple questions are posed consecutively.
-Plan-and-solve methods [45] prompts LLMs to generate a plan and conduct reasoning based on it.
-DecomP [12] prompts LLMs to decompose the reasoning task into a series of sub-tasks and solve them step by step.
-Semantic Parsing/Code Generation methods [5, 53, 7] leverages LLMs to convert natural language queries into executable code or structured representations.
-This approach enables more precise control over the output and facilitates better interpretability.
-However, it has its limitations, such as the limited coverage of programming language grammar and semantics.
-This restricts the model’s ability to handle complex programming tasks and may require additional techniques to handle out-of-domain queries effectively.
-[5] propose to leverage LLMs to bridge the out-of-domain queries, however, the form of knowledge injection still lacks a thorough exploration and the trigger functions are still naive for complex queries.
-However, the problem of hallucinations and lack of knowledge affect the faithfulness of LLMs’ reasoning.
-ReACT [51] treats LLMs as agents, which interact with the environment to get the latest knowledge for reasoning.
-To explore faithful reasoning, FAME [16] introduces the Monte-Carlo planning to generate faithful reasoning steps.
-RR [13] and KD-CoT [44] further retrieve relevant knowledge from KGs to produce faithful reasoning plans for LLMs.
-Appendix G Experiment Details
-G.1 Baselines
-RoG [29]: It aims to embed structure knowledge from KGs into the underling neural networks during the pretraining and fine-tuning process.
-However, the controllablity of the reasoning paths generation is hard to achieve since the decoding process can be shortsighted, i.e., it is difficult to tell whether the current operation/decision would lead to a malicious operation several steps later.
-Also, the reasoning steps generated by LLMs are observed to be prone to errors, which could be non-exist in the knowledge graph, and can subsequently lead to incorrect reasoning during inference.
-Additionally, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating [17].
-Compared to RoG, FiDeLiS propose to employ beam search tailored for intermediate reasoning steps and guide the beam searching process by controlling the error of each reasoning step to prevent accumulation thought the chaining.
-ToG [38]: It aims step-wise grounding to elicit more accurate results.
-However, it directly employs LLMs to output scores for candidate selections from the KG, which require an LLM/LM to iteratively score candidate plans, which are resource-consuming in terms of both time and computing.
-Also, the performance of the proposed ToG [38] seems rely heavily on the use of very strong LLMs (including Codex [3], GPT-4 [31]) to do the final reasoning; and the true potential of enabling few-shot grounded reasoning path understanding has yet to be fully realized.
-G.2 Datasets
+Appendix D Experiment Details
+D.1 Baselines
+Baselines.
+(1) RoG (Luo et al., 2024): embed structure knowledge graph from KGs into the underlying neural networks during the pretraining and fine-tuning process to generate the reasoning path and explanation.
+(2) ToG (Sun et al., 2023): ask LLM to iteratively explore multiple possible reasoning paths on KGs until the LLM determines that the question can be answered based on the current reasoning paths.
+Our method FiDeLiS follows the similar paradigm.
+(3) NSM (He et al., 2021) utilizes the sequential model to mimic the multi-hop reasoning process.
+(4) KD-CoT (Wang et al., 2023a) retrieves relevant knowledge from KGs to generate faithful reasoning paths for LLMs.
+(5) DeCAF (Yu et al., 2022) combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance on KGQA tasks.
+Implementation Details.
+We set the default beam width as $4$ and depth as $4$ without specific annotation.
+We set the $\alpha$ in Eq 3 as $0.3$ to ensure reproducibility.
+For LLMs usage, we set all the inference using temperature $T=0.3$ and $p=1.0$.
+The hyperparameter tuning experiments for beam search and $\alpha$ can be found in Figure 2 and Table 9.
+Backboned LLMs.
+We assess our approach on closed- and open-source LLMs.
+For closed-source LLMs, we choose GPT-4-turbo, GPT-3.5-turbo [^1] We use the recent gpt-4-turbo model that is released in 2024-04-09 from https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 and gpt-3.5-turbo model notated as “gpt-3.5-turbo-0125” from https://platform.openai.com/docs/models/gpt-3-5-turbo.
+to report and compare the results on all datasets.
+We use Llama-2-13B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023a) as our open-source LLMs to conduct cost–performance analysis on different datasets.
+The experiment results of open-source models can be found in Table 10.
 Table 7: Statistics of the number of answers for questions in WebQSP and CWQ.
@@
 CWQ 70.6% 19.4% 6% 4%
+D.2 Datasets
+Datasets & Metrics.
+We adopt three benchmark KGQA datasets: WebQuestionSP (WebQSP) (Yih et al., 2016), Complex WebQuestions (CWQ) (Talmor & Berant, 2018) and CR-LT-KGQA (Guo et al., 2024) in this work.
+We follow previous work (Luo et al., 2024) to use the same training and testing splits for fair comparison over WebQSP and CWQ.
+The questins from both WebQSP and CWQ can be reasoned using Freebase KGs [^2] https://github.com/microsoft/FastRDFStore.
+To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we further test our method on CR-LT-KGQA (further discussed in Appendix Section 8).
+We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples.
+Each of the question can be reasoned based on the Wikidata [^3] https://www.wikidata.org/wiki/Wikidata:Main_Page.
+The statistics of the datasets are given in Table 7 and Table 8.
+To streamline the KGs, we utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ followsing RoG (Luo et al., 2024).
+Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA as well.
+We assess the performance of the methods by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
 Table 8: Statistics of the question hops in WebQSP, CWQ and CR-LT-KGQA.
@@
 CR-LT 5.31 % 43.22% 51.57%
-Figure 5: Distribution of CR-LT-KGQA dataset.
-Motivation of CR-LT-KGQA The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities.
+Motivation of CR-LT-KGQA.
+The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities.
 These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs.
 Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering.
+Figure 4: Distribution of CR-LT-KGQA dataset.
 Against this backdrop, we specifically chose CR-LT-KGQA for our evaluation as it includes queries related to obscure or long-tail entities, the distribution of the frequency and entity popularity of CR-LT is shown as Figure 4.
@@
 This evaluation helps us understand the extent to which KGs remain necessary for supporting LLMs in a diverse range of query scenarios.
-Table 10: Comparison of FiDeLiS using different backbone LLMs.
-Backend Models Methods WebQSP CWQ CR-LT
-Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
-Llama-2-13B FiDeLiS 72.34 69.78 58.41 54.78 60.87
-Mistral-7B FiDeLiS 74.11 70.23 60.71 56.87 63.12
-Example of Deductive Verification Question: “Who is the ex-wife of Justin Bieber’s father?" After one round of beam searching, the current reasoning path is “Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber", and the next step candidates are “1.
-people.married_to.person $\to$ Erin Wagner"; 2.
-people.person.place_of_birth $\to$ US, $\ldots$".
-The deductive reasoning can be formulated as: Premises: Justin_bieber $\to$ people.person.father —> Jeremy_bieber (from the current reasoning path) Jeremy_bieber $\to$ people.married_to.person $\to$ Erin Wagner (from the next step candidates) Conclusion: Erin Wagner is the ex-wife of Justin Bieber’s father.
-(Using LLM zero-shot to reformat the question to a cloze filling, and use the last entity from the next step candidate"Erin Wagner” to fill the cloze) The prompt will be whether the conclusion can be deduced from the given premises, if yes, return “yes”, it no, return “no".
-Appendix H Algorithm Description
+D.3 Parameter Tuning for $\alpha$ for Scoring Function
+For hyperparameter tuning regarding $\alpha$ for Eq 3, we added an extra comparison in Table 9 which shows that $\alpha$ is actually not a very impact parameter for the whole system, however for the reproducibility of our method, we set the $\alpha$ as 0.3.
+Table 9: Parameter tuning for $\alpha$ for scoring function over WebQSP
+$\alpha$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
+Hits@1 83.76 84.15 84.39 83.98 83.55 84.12 83.79 82.25 83.60 83.73
+Appendix E Algorithm Description
 Algorithm 1 Path-RAG Initialization and Retrieval Process
@@
 21:return the best hypothesis from $\mathcal{H}_{T}$.
+Appendix F Prompt List
+In this section, we show all the prompts that need to be used in the main experiments. The In-Context Few-shot refers to the few-shot examples we used for in-context learning.
+F.1 Plan-and-solve
+You are a helpful assistant designed to output JSON that aids in navigating a knowledge graph to answer a provided question. The response should include the following keys:
+(1) ’keywords’: an exhaustive list of keywords or relation names that you would use to find the reasoning path from the knowledge graph to answer the question.
+Aim for maximum coverage to ensure no potential reasoning paths will be overlooked;
+(2) ’planning_steps’: a list of detailed steps required to trace the reasoning path with. Each step should be a string instead of a dict.
+(3) ’declarative_statement’: a string of declarative statement that can be transformed from the given query, For example, convert the question ’What do Jamaican people speak?’ into the statement ’Jamaican people speak *placeholder*.’ leave the *placeholder* unchanged; Ensure the JSON object clearly separates these components.
+In-Context Few-shot
+Q: {Query}
+A:
+F.2 Deductive-verification
+You are asked to verify whether the reasoning step follows deductively from the question and the current reasoning path in a deductive manner. If yes return yes, if no, return no".
+In-Context Few-shot
+Whether the conclusion ’{declarative_statement}’ can be deduced from ’{parsed_reasoning_path}’, if yes, return yes, if no, return no.
+A:
+F.3 Adequacy-verification
+You are asked to verify whether it’s sufficient for you to answer the question with the following reasoning path.
+For each reasoning path, respond with ’Yes’ if it is sufficient, and ’No’ if it is not.
+Your response should be either ’Yes’ or ’No’.
+In-Context Few-shot
+Whether the reasoning path ’{reasoning_path}’ be sufficient to answer the query ‘{Query}’, if yes, return yes, if no, return no.
+A:
+F.4 Beam Search
+Given a question and the starting entity from a knowledge graph, you are asked to retrieve reasoning paths from the given reasoning paths that are useful for answering the question.
+In-Context Few-shot
+Considering the planning context {plan_context} and the given question {Query}, you are asked to choose the best {beam_width} reasoning paths from the following candidates with the highest probability to lead to a useful reasoning path for answering the question.
+{reasoning_paths}.
+Only return the index of the {beam_width} selected reasoning paths in a list.
+A:
+F.5 Reasoning
+Given a question and the associated retrieved reasoning path from a knowledge graph, you are asked to answer the following question based on the reasoning path and your knowledge.
+Only return the answer to the question.
+In-Context Few-shot
+Question: {Query}
+Reasoning path: {reasoning_path}
+Only return the answer to the question.
+A:
+F.6 Demonstration of Deductive Verification
+Deductive Verification Example Question: Who is the ex-wife of Justin Bieber’s father?
+After one round of beam searching, the current reasoning path is: Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber.
+The next step candidates are: 1.
+people.married_to.person $\to$ Erin Wagner 2.
+people.person.place_of_birth $\to$ US, $\ldots$ The deductive reasoning can be formulated as follows: Premises: - Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber (from the current reasoning path) - Jeremy_bieber $\to$ people.married_to.person $\to$ Erin Wagner (from the next step candidates) Conclusion: Erin Wagner is the ex-wife of Justin Bieber’s father.
+(Using a large language model (LLM) zero-shot approach to reformat the question into a cloze filling task, we use the last entity from the next step candidates, "Erin Wagner", to fill the cloze.) The prompt will ask whether the conclusion can be deduced from the given premises.
+If the answer is "yes", return “yes”, otherwise return “no.”
+Table 10: Comparison of FiDeLiS using different backbone LLMs.
+Backend Models Methods WebQSP CWQ CR-LT
+Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
+Llama-2-13B FiDeLiS 72.34 69.78 58.41 54.78 60.87
+Mistral-7B FiDeLiS 74.11 70.23 60.71 56.87 63.12
+Appendix G Related Works: Reasoning with LLM Prompting
+With the development of LLMs, many creative ways to leverage LLMs are proposed: End-to-End, Chain-of-Thought, and Semantic Parsing/Code Generation.
+End-to-End methods (Chen, 2022; Hoffmann et al., 2022) aims to leverage LLMs to generate final answers directly, often done by providing a task description and/or a few examples for in-context learning.
+While it offers convenience, however, it suffers from the un-interpretability of the generation, and the lack of explicit steps and reliance solely on the LLMs’ training data may result in un-robustness and exhibit sensitivity to slight input variations.
+Chain-of-Thought methods (Wei et al., 2022a; Kojima et al., 2022; Gong et al., 2020) emphasize breaking down complex reasoning into a series of intermediate steps to support complex reasoning.
+However, CoT suffers from unreliability and uncontrollability since the generated reasoning steps may not always align with the intended logical progression or may produce incorrect or inconsistent answers when multiple questions are posed consecutively.
+Plan-and-solve methods (Wang et al., 2023b) prompts LLMs to generate a plan and conduct reasoning based on it.
+DecomP He et al.
+(2021) prompts LLMs to decompose the reasoning task into a series of sub-tasks and solve them step by step.
+Semantic Parsing/Code Generation methods (Cheng et al., 2023; Ye et al., 2023; Gemmell & Dalton, 2023) leverages LLMs to convert natural language queries into executable code or structured representations.
+This approach enables more precise control over the output and facilitates better interpretability.
+However, it has its limitations, such as the limited coverage of programming language grammar and semantics.
+This restricts the model’s ability to handle complex programming tasks and may require additional techniques to handle out-of-domain queries effectively.
+Cheng et al.
+(2023) propose to leverage LLMs to bridge the out-of-domain queries, however, the form of knowledge injection still lacks a thorough exploration and the trigger functions are still naive for complex queries.
+However, the problem of hallucinations and lack of knowledge affect the faithfulness of LLMs’ reasoning.
+ReACT Yao et al.
+(2022) treats LLMs as agents, which interact with the environment to get the latest knowledge for reasoning.
+To explore faithful reasoning, FAME Hong et al.
+(2023) introduces the Monte-Carlo planning to generate faithful reasoning steps.
+RR He et al.
+(2022) and KD-CoT Wang et al.
+(2023a) further retrieve relevant knowledge from KGs to produce faithful reasoning plans for LLMs.

### v3

@@
 Abstract
-Large language models are often challenged by generating erroneous or ‘hallucinated’ responses, especially in complex reasoning tasks.
-To mitigate this, we propose a retrieval augmented reasoning method, FiDeLiS, which enhances knowledge graph question answering by anchoring responses to structured, verifiable reasoning paths.
-FiDeLiS uses a keyword-enhanced retrieval mechanism that fetches relevant entities and relations from a vector-based index of KGs to ensure high-recall retrieval.
-Once these entities and relations are retrieved, our method constructs candidate reasoning paths which are then refined using a stepwise beam search.
-This ensures that all the paths we create can be confidently linked back to KGs, ensuring they are accurate and reliable.
-A distinctive feature of our approach is its blend of natural language planning with beam search to optimize the selection of reasoning paths.
-Moreover, we redesign the way reasoning paths are scored by transforming this process into a deductive reasoning task, allowing the LLM to assess the validity of the paths through deductive reasoning rather than traditional logit-based scoring.
-This helps avoid misleading reasoning chains and reduces unnecessary computational demand.
-Extensive experiments demonstrate that our method, even as a training-free method which has lower computational costs and superior generality, outperforms established strong baselines across three datasets.
-The code of this paper will be released at https://anonymous.4open.science/r/FiDELIS-E7FC.
+Large language models (LLMs) are often challenged by generating erroneous or hallucinated responses, especially in complex reasoning tasks.
+Leveraging knowledge graphs (KGs) as external knowledge sources has emerged as a viable solution.
+However, existing KG-enhanced methods, either retrieval-based or agent-based, encounter difficulties in accurately retrieving knowledge and efficiently traversing KGs at scale.
+In this paper, we propose a unified framework, FiDeLiS, designed to improve the factuality of LLM responses by anchoring answers to verifiable reasoning steps retrieved from a KG.
+To achieve this, we leverage step-wise beam search with a deductive scoring function, allowing the LLM to validate each reasoning step and halt the search once the question is deducible.
+In addition, our Path-rag module pre-selects a smaller candidate set for each beam search step, reducing computational costs by narrowing the search space.
+Extensive experiments show that our training-free and efficient approach outperforms strong baselines, enhancing both factuality and interpretability.
+Code is released at https://anonymous.4open.science/r/FiDELIS-E7FC.
+FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering
+Yuan Sui1, Yufei He1, Nian Liu1, Xiaoxin He1, Kun Wang1,2, Bryan Hooi1 1National University of Singapore 2University of Science and Technology of China {yuansui, yufei.he, nianliu, xiaoxin, bhooi}@comp.nus.edu.sg, wk520529@mail.ustc.edu.cn
 1 Introduction
-The emergence and application of large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; LLaMa-v3, 2024) have attracted widespread attention from researchers and the general public.
-They demonstrate remarkable reasoning capabilities, managing to solve complex reasoning problems through step-by-step thinking and planning (Wang et al., 2022b; Wei et al., 2022b).
-However, the reasoning of LLMs is not always reliable and may conflict with factual reality (Pan et al., 2024), limiting their application in areas requiring high reliability, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
-Knowledge graphs (KGs) store high-quality knowledge in structured triplets, such as Wikidata (Vrandečić & Krötzsch, 2014), YAGO (Suchanek et al., 2007), and NELL (Carlson et al., 2010).
-KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning.
-Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance.
-This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
-Due to its reliability and interpretability, it is considered as a promising method to improve the reliability of LLM reasoning.
-Therefore, several attempts have been conducted to integrate KGs with language models (Pan et al., 2024; Luo et al., 2024; Hu et al., 2023; Sun et al., 2023).
-Among them, knowledge graph question answering (KGQA) is a critical task to verify the effectiveness of incorporating the knowledge from KGs into reasoning models (He et al., 2021; Wang et al., 2023a; Yu et al., 2022).
-However, effectively integrating KGs into reasoning models for KGQA presents several challenges, including data sparsity, the complexity of query interpretation, and the need for advanced inference capabilities.
-More specifically, we focus on two main questions concerning the integration of KGs with reasoning models as follows:
-(I) How to retrieve specific knowledge from KGs to allow more precise reasoning?
-Existing solutions include direct retrieval (Sun et al., 2019; Jiang et al., 2022) and semantic parsing (Sun et al., 2020; Gu & Su, 2022).
-Direct retrieval utilize the query to find relevant knowledge triplets within the KGs.
-However, these triplets sometimes lack comprehensive semantic information due to variations in KG schema, such as entities being represented as machine identifiers (MIDs), rather than descriptive labels.
-This lack of semantic richness can obscure the relevance of the retrieved information with the user query, and becomes especially tricky in multi-hop question answering, where seemingly unrelated intermediate triplets may be pivotal in deriving the correct answer.
-On the other hand, semantic parsing aim to transform user queries into executable structured queries (e.g., SPARQL), which are then run on KGs.
-However, these methods often grapple with issues related to the non-executability or incorrect execution of the generated queries (Yu et al., 2022), which undermines the relaibility of these methods.
-(II) How to make the reasoning model understand and utilize the retrieved structured knowledge in KGs?
-Existing solutions include (1) finetuning LMs to generate relation paths grounded by KGs.
-These reasoning paths are then used for LLMs to conduct faithful reasoning and generate interpretable results (Luo et al., 2024; Yu et al., 2022); (2) leverage the reasoning model to iteratively retrieve and reason over subgraphs from the KG, deciding at each step to either answer the question, or to continue the searching step (Sun et al., 2023; Gu et al., 2023; Jiang et al., 2022).
-However, for the first approach, the reasoning steps generated by language models are not guaranteed to exist in the KG, especially when multiple consecutive steps are combined into a reasoning path.
-We conduct error analysis in Section 3 and show that only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG.
-The latter solution faces the challenge of determining the optimal stopping point for the exploration process.
-For example, ToG (Sun et al., 2023) prompts LLMs to assess the adequacy of current reasoning paths for answer generation.
-However, assessing the adequacy of reasoning paths is itself a complex task that demands a deep understanding of the domain.
-This lack of clarity can easily result in premature stopping or excessive continuation, further complicating the decision-making process in LLMs.
-To address these two challenges, we propose a retrieval augmented reasoning method, FiDeLiS.
-It is designed to enhance KGQA by anchoring responses to structured, verifiable reasoning paths.
-FiDeLiS is composed of two major components, Path-RAG, which retrieves chain of entities and relations from KGs in a effective manner (discussed in Section 2.1), and Deductive-verification Beam Search (DVBS), which conducts deductive-reasoning-based beam search to generate multiple reasoning paths leading to final answers (discussed in Section 2.2).
-The Path-RAG module uses a keyword-enhanced retrieval mechanism that fetches relevant entities and relations from a vector-based index of KGs to ensure high-recall retrieval.
-Compared to a vanilla retriever which easily returns outputs that are not useful for finding the correct answer based on relevance to the query, we incorporate LLMs to generate an exhaustive list of keywords from the query to maximize coverage and ensure that no potential reasoning paths will be overlooked.
-Once these entities and relations are retrieved, the module constructs candidate reasoning steps which are then refined using a stepwise beam search discussed below.
-Next, DVBS leverages the deductive reasoning capabilities of LLMs (Ling et al., 2024; Huang & Chang, 2023) as a clear-defined criterion for automatically directing the beam search process step by step to create the complete reasoning paths.
-A key characteristic of DVBS is its integration of natural language planning (Song et al., 2023; Zhou et al., 2022) and beam search.
-This combination enhances the process of choosing the best reasoning paths by providing additional hints for decision making in LLMs.
-In addition, we redesign how reasoning paths are evaluated by converting the scoring process into a deductive reasoning task.
-This allows the LLM to validate the reasoning paths through deductive verification rather than traditional logit-based scoring.
-This proposed deductive verification serves as presie indicators for when to cease further reasoning, thus avoiding misleading reasoning chains and reducing the computational resources required.
+Large language models (LLMs) have shown impressive reasoning capabilities in tackling complex tasks Yu et al.
+(2024).
+However, the reasoning of LLMs is not always reliable and can be prone to generating outputs that are either inconsistent with real-world facts Xu et al.
+(2024); Huang et al.
+(2025) or show flawed reasoning process Li et al.
+(2024); Sui et al.
+(2024), which greatly undermine the reliability of LLMs in real-world applications Kung et al.
+(2023); Zhang et al.
+(2024).
+Figure 1: Challenges for existing KG-enhanced methods: How to balance faithfulness and efficiency?
+To address this issue, leveraging knowledge graphs (KGs) as external knowledge sources has emerged as a viable solution Sun et al.
+(2023); Ma et al.
+(2024); Luo et al.
+(2024a).
+Unlike traditional retrieval-augmented generation (RAG) that relies on web pages or documents Liu et al.
+(2024); Qian et al.
+(2024); Bayarri-Planas et al.
+(2024), KGs represent information in a structured and interconnected format, where each fact is stored as entities and relations.
+This format supports explicit, traceable reasoning processes Pan et al.
+(2023) and facilitates multi-hop reasoning through graph traversal.
+Moreover, each fact in a KG can be traced back to its source Sui et al.
+(2024); Agrawal et al.
+(2024), providing both context and original details, which further enhances the information authenticity and reliability of the reasoning processes.
+Existing KG-enhanced LLM reasoning methods face notable challenges and can be roughly categorized into two primary approaches: retrieval-based and agent-based paradigms Luo et al.
+(2024b).
+Retrieval-based methods Wang et al.
+(2023); Luo et al.
+(2024a); Baek et al.
+(2023) retrieve relevant KG facts to support LLM reasoning by either prompting Baek et al.
+(2023) or fine-tuning LLMs to learn the underlying structure of KG Luo et al.
+(2024a, b).
+These methods often suffer from incomplete or imprecise information extraction due to a lack of contextual understanding or an inability to fully capture the graph structure Luo et al.
+(2024b).Our error analysis of a strong retrieval-based method (i.e., Luo et al.
+(2024a)) in §4.3 reveals that only 67% of the generated reasoning steps are valid, with 33% containing format errors or referencing non-existent KG facts.
+In contrast, agent-based methods Sun et al.
+(2023); Ma et al.
+(2024) treat LLMs as interactive agents that explore KGs iteratively to construct reasoning paths and generate answers.
+While this approach by nature can enhance reasoning accuracy, it is computationally expensive, resulting in high latency and scalability limitations.
+As illustrated in Figure 1, balancing faithfulness and efficiency remains a critical challenge for existing KG-enhanced reasoning methods.
+To this end, we propose FiDeLiS, a unified framework designed to improve the factual accuracy and reasoning efficiency of LLMs on KGQA task.
+FiDeLiS anchors LLM responses to verifiable reasoning steps derived from a KG by employing two core components: (1) Deductive-Verification Beam Search (DVBS) which systematically constructs and validates reasoning paths step-by-step, ensuring logical consistency and factual correctness (discussed in §3.2).
+This module also prevents premature reasoning termination and incorrect path extension to ensure the validity of the generated reasoning paths.
+(2) Path-RAG, a retrieval-augmented mechanism that pre-selects a constrained set of candidate entities and relations for each step to mitigate computational inefficiencies.
+It combines semantic similarity measures with graph-based connectivity analysis to optimize the search space, significantly reducing latency without sacrificing recall or accuracy (discussed in §3.1).
+Extensive experiments show that our method outperform strong baselines in both accuracy and efficiency, offering a scalable, training-free solution for KG-enhanced LLM reasoning.
 Overall, our main contributions include:
-  • We propose a retrieval augmented reasoning method, which enhances knowledge graph question answering by anchoring responses to structured, verifiable reasoning paths grounded by KGs.
-  • We propose a step-wise keyword retrieval method that enhances the recall of relevant intermediate knowledge from KGs.
-This ensures that all the paths we create can be confidently linked back to KGs, ensuring that they are faithful and reliable.
-  • We propose deductive verification as precise indicators for when to cease further reasoning, thus avoiding misleading the chains of reasoning and reducing unnecessary computation required.
-  • Extensive experiments show that our method, as a training-free method with lower computational cost and better generality, outperforms existing strong baselines in three datasets.
+  • We propose FiDeLiS, a unified framework designed to improve the factual accuracy of LLMs by grounding reasoning paths in structured KG efficiently.
+  • We enable efficient, verifiable reasoning by deductively validating reasoning steps and narrowing the search space with high-quality retrieval mechanism.
+  • FiDeLiS performs robustly across different experiments without the need for model fine-tuning, demonstrating adaptability and scalability with improved performance on multiple benchmarks.
+  • By anchoring responses in verifiable reasoning paths, FiDeLiS enhances interpretability, enabling users to verify and understand each reasoning step.
+2 Preliminary
+Notation. To facilitate the demonstration of our method, we define the necessary notation below:
+  • Definition 1. A reasoning step is a pair $(r,e)$, where $r$ is the relation and $e$ is the corresponding entity.
+  • Definition 2.
+A reasoning path $\mathcal{P}$ is a pair $(s,\mathcal{T})$, where $s$ is the starting entity for the reasoning path, and $\mathcal{T}$ is a sequence of reasoning steps $\mathcal{T}=\left\{t_{1},\ldots,t_{n}\right\}$ and $t_{k}=(r_{k},e_{k})$ denotes the $k$-th reasoning step in the path and $n$ denotes the length of the path.
+  • Definition 3.
+The next-hop candidates given path $\mathcal{P}$, denoted $\mathcal{N}_{1}(e_{n})$, is defined as the 1-hop neighborhood of $e_{n}$, the last node in the reasoning path $\mathcal{P}$.
+[…]
+Task definition.
+In this work, we focus on the task of knowledge graph-based question answering (KGQA), a common reasoning task involving KGs.
+It is defined as: given a user query $q$ and a KG $\mathcal{G}=\{(e,r,e^{\prime})\mid e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations in KG, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ conditioned on $q$ and $\mathcal{G}$.
+Following existing KG-enhanced LLMs methods Sun et al.
+(2023); Ma et al.
+(2024), the function $f$ can be generally expressed as finding valid reasoning path(s) $\mathcal{P}$ on KGs that connects the entities mentioned in the query and the answer as: $P(a|q,\mathcal{G})=\sum_{\mathcal{P}}P_{\theta}(a|q,\mathcal{P})P_{\phi}(\mathcal{P}|q,\mathcal{G})$, where $P_{\theta}(a|q,\mathcal{P})$ denotes the probability of generating answer $a$ conditioned on $q$ and reasoning path(s) $\mathcal{P}$ by a function parameterized by $\theta$, and $P_{\phi}(\mathcal{P}|q,\mathcal{G})$ denotes the probability of discovering reasoning path(s) $\mathcal{P}$ by a function parameterized by $\phi$.
+As reasoning path $\mathcal{P}$ is defined as a sequence of reasoning steps, we factorize the reasoning path probability using the chain rule as Eq 1:
+$P(a|q,\mathcal{G})=\sum_{\mathcal{P}}P_{\theta}(a|q,\mathcal{P})\prod_{k=1}^{n}P_{\phi}(t_{k}|q,t_{<k},\mathcal{G})$ (1)
+To acquire valid reasoning paths, most prior studies follow the retrieval-based Li et al.
+(2023); Luo et al.
+(2024a) or agent-based Sun et al.
+(2023) paradigm.
+As indicated in Luo et al.
+(2024b), retrieval-based methods rely on precise additional retrievers, while agent-based methods are computationally intensive and lead to high latency.
+To address these issues, we propose our method, FiDeLiS, to enable both efficient and faithful reasoning over KGs.
+Figure 2: An illustration of FiDeLiS.
+Top: The workflow of Path-RAG.
+An LLM first extracts key terms and generates dense embeddings that feed into the Path-RAG module.
+Path-RAG rapidly retrieves relevant entities and relations from a pre-embedded KG and constructs candidate reasoning steps by combining semantic similarity with graph connectivity.
+Bottom: The workflow of DVBS.
+Then, the DVBS module uses an LLM-generated planning outline to guide a beam search that builds reasoning paths step-by-step over candidates constructed by Path-RAG, with deductive verification ensuring each step logically follows the previous steps and support the user question.
 3 Method
-[…]
-Overview.
-The framework of FiDeLiS is as shown in Figure 1.
-Given a query, it first retrieves the reasoning step candidates from KGs (as discussed in Section 2.1).
-To ensure high recall, our method first generates a list of keywords from the query.
-And then based on the generated keywords, it retrieves entities and relations from an index (where the entities and relations are also embedded by vector representation obtained from an LM).
-Next, following the retrieved entities, n-hop reasoning step candidates are created and scored based on semantics similarity and neighborhoods aggregation.
-Then, a deductive verification-based beam search (discussed in Section 2.2) is employed to generate the top-k reasoning paths leading to the answer.
-First, a natural language plan is generated from the query to provide more hints for LLMs decision making and then a beam search procedure is conducted to score each reasoning step.
-We redesign the scoring function by converting the process into a deductive reasoning task.
-This allows the LLM to validate connections between paths through simple verification rather than traditional logit-based scoring.
-This change helps prevent easily misleading reasoning chains and reduces the computational resources required.
-Figure 1: Overview of FiDeLiS.
-The goal of FiDeLiS is to extend the multiple-step reasoning paths $T$ grounded by KG.
-The figure shows given a query, Path-RAG first retrieve entities and relations that are relevant to the query, and construct the reasoning path candidates.
-Then DVBS constructs reasoning paths by iterative extending these path candidates using beam search.
-Specifically, at timestep $t$, DVBS leverage LLMs to choose the top-$k$ (here we set $k=2$ as an example) reasoning steps and decide whether to continue the next search step or cease the reasoning path extension based on deductive verification.
-At last, FiDeLiS return the top-$k$ reasoning paths as the final output leading to the answers of the query.
+Motivated by the insight that integrating KGs with LLMs can mitigate hallucinations and enable verifiable reasoning, we propose FiDeLiS to improve the factuality of LLM responses by anchoring answers to verifiable reasoning steps retrieved from a KG.
+The overall framework of FiDeLiS is illustrated in Figure 2, which consists of two main components: (1) Reasoning Path Retrieval-Augmented Generation (Path-RAG, Algorithm 1) and (1) Deductive-verification Beam Search (DVBS, Algorithm 2).
+Given a complex question $q$, we first use an LLM to extract key terms from $q$ and generate dense embeddings that capture the question’s core concepts.
+These embeddings are input into Path-RAG module, which rapidly retrieves relevant entities and relations from a pre-embedded KG to select a smaller candidate sets for further beam search step, addressing the latency and computational burden of traditional agent-based methods.
+Path-RAG then constructs candidate reasoning steps by combining immediate semantic similarity with the structural connectivity of the graph, overcoming the dependence on highly precise retrievers in standard retrieval-based approaches.
+Next, the DVBS module employs an LLM-generated planning outline to guide a beam search that builds reasoning paths step-by-step.
+At each step, deductive verification checks that the accumulated reasoning steps logically supports the user question, ensuring the final reasoning path is both verifiable and accurate.
 3.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
-Path-RAG aims to retrieve reasoning path candidates from KGs and comprises three main steps: initialization, retrieval, reasoning step candidates construction, as depicted in Figure 1.
-The implementation details of each step are elaborated in the following paragraphs.
+Previous agent-based methods Ma et al.
+(2024); Sun et al.
+(2023) treat LLMs as agents that iteratively interact with KGs to find reasoning paths and answer, which necessitate multiple rounds of interaction between agents and KGs and lead to high computational costs and latency.
+We instead propose a module, Path-RAG which iteratively pre-select a smaller candidate set to reduces the search space for exploring the potential reasoning paths from KGs.
+It consists of three steps and we detail the workflow as follows:
 Initialization.
-We initiate the Path-RAG by generating embeddings for entities (nodes) and relations (edges) using a pre-trained language model (LM).
-We begin by extracting entities ($e_{i}$) and relations ($r_{i}$) from the KG, denoted as $\mathcal{E}$ and $\mathcal{R}$, respectively.
-Each entity and relation is then encoded using a pre-trained language model such as SentenceBert (Reimers & Gurevych, 2019) or E5 (Wang et al., 2022a), yielding dense vector representations:
-$z(e^{i})=\operatorname{LM}(e_{i})\in\mathbb{R}^{d},\ z(r^{i})=\operatorname{LM}(r_{i})\in\mathbb{R}^{d}$ (1)
-where $d$ denotes the dimension of the output vector.
-These embeddings are stored in a nearest neighbor data structure, which facilitates quick retrieval of similar entities or relations based on their vector representations.
-Retrieval.
-To retrieve the entities and relations that are relevant to the user query, the embeddings of them are utilized to populate a nearest neighbor index.
-Specifically, for a given query $q$, we use LLM to generate an exhaustive list of keywords or relation names referred to as $\mathcal{K}$ (the prompt is given in Section F).
-These keywords are generated to maximize the coverage of potential reasoning steps essential for answering the query.
-The list is derived as $\mathcal{K}=\text{LM}(\text{prompt}_{p},q)$.
-Then each keyword $\in\mathcal{K}$ is concatenated into a string denoted as $K$ and then encoded into the same latent space by the same LM used in initialization step to obtain $z(K)=\text{LM}(K)\in\mathbb{R}^{d}$.
-To find the entities and relations that best match the keyword embeddings, we compute the cosine similarity $\cos(\cdot,\cdot)$ between $z(K)$ and the embeddings of entities ($z(e)$) and relations ($z(r)$) in the KG.
-We then select the top-$m$ elements that exhibit the highest similarity:
-$\mathcal{E}_{m}=\operatorname{argtopm}_{i\in\mathcal{E}}\cos\left(z(K),z(e)\right),\mathcal{R}_{m}=\operatorname{argtopm}_{i\in\mathcal{R}}\cos\left(z(K),z(r)\right)$ (2)
-Each entity and relation in the retrieved set is then assigned a score, reflective of its similarity to the keyword embedding as $S_{\text{ent}}(e)=\cos\left(z(K),z(e)\right)$ and $S_{\text{rel}}(r)=\cos\left(z(k),z(r)\right)$.
+We initiate the Path-RAG by encoding each entity $e_{i}\in\mathcal{E}$, and relation $r_{i}\in\mathcal{R}$ in the KG using a pre-trained language model (LM), which produces dense vectors $z(e^{i})=\mathrm{LM}(e_{i})\in\mathbb{R}^{d}$ and $z(r^{i})=\mathrm{LM}(r_{i})\in\mathbb{R}^{d}$, where $d$ denotes the embedding dimension.
+These embeddings are stored in a nearest neighbor structure to facilitate rapid similarity search.
+Keyword-Driven Retrieval.
+We then populate a nearest neighbor index to retrieve relevant entities and relations for the user query.
+We first use an LLM to analysis the user query and generate exhaustive keywords/relation names that could be useful for finding the reasoning path to answer the query (See the prompt in §C.1).
+This step is designed to maximize coverage of potential reasoning steps, ensuring that no potential reasoning paths are overlooked during the retrieval process.
+The extracted keywords are then encoded using the same LM in initialization, yielding $z(K)=\mathrm{LM}(K)\in\mathbb{R}^{d}$.
+We subsequently compute the cosine similarity between $z(K)$ and the pre-stored embeddings, retrieving the top-$m$ entities and relations: $\mathcal{E}_{m}=\operatorname{argtopm}_{i\in|\mathcal{E}|}\cos\left(z(K),z(e_{i})\right)$ and $\mathcal{R}_{m}=\operatorname{argtopm}_{i\in|\mathcal{R}|}\cos\left(z(K),z(r_{i})\right)$.
 Reasoning Step Candidates Construction.
-After the initial retrieval process, which runs once at the start of the algorithm, we iteratively construct reasoning step candidates to extend the reasoning paths based based on entities $\mathcal{E}_{m}$ and relations $\mathcal{R}_{m}$ that are relevant to the given query.
-To guide the selection of these candidates, we propose another scoring function based on the derived relevance score $S_{\text{ent}}(e)$ and $S_{\text{rel}}(r)$.
-Our scoring function leverages triplets to represent the connectivity of KGs and assess whether extending a path in certain directions might lead to negative outcomes over multiple steps by incorporating next-hop neighbor’s information.
-$S((r,e))=S_{\text{rel}}(r)+S_{\text{ent}}(e)+\alpha\max_{\forall(r_{j},e_{j})\in N(e)}\left(S_{\text{rel}}(r_{j})+S_{\text{ent}}(e_{j})\right)$ (3)
-Here, $(r,e)$ refers to the reasoning step defined as a relation-entity pair.
-$S_{\text{rel}}(r_{j})$ and $S_{\text{ent}}(e_{j})$ are the similarity scores for relations and entities at the next step.
-The component $N(e)$ corresponds to entities reachable from $e$ within one hop.
-The factor $\alpha$ is used to balance short-term outcomes and long-term potential in reasoning paths: a higher $\alpha$ prioritizes paths with long-term benefits, even if they seem sub-optimal initially, while a lower $\alpha$ emphasizes immediate steps, potentially overlooking future impacts.
-Details of Path-RAG are provided in Algorithm 1.
-2.2 DVBS: Deductive-Verification Guided Beam Search
-DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) to find the top-$k$ most promising reasoning paths that lead to the answer to the user query.
-At each timestep, we leverage the LLMs to choose the top-$k$ reasoning steps and decide whether to continue extending the reasoning paths or cease the extension process.
-It comprises three main steps: planning, beam-search, deductive-verification, as depicted in Figure 1.
-The implementation details of each step are elaborated in the following paragraphs.
-Plan-and-Solve.
-Inspired by the recent works regarding planning capabilities of LLMs (Song et al., 2023; Zhou et al., 2022), we prompt LLM to generate the planning steps for answering the user query, denoted as $w$.
-This step aims to provide more hints for subsequent LLM decision making process.
-The detailed prompt can be found in Section F.
-Beam Search.
-To enable multi-step reasoning, a reasoning path of $T$ steps is sequentially generated through several time steps as $\left[s^{1},s^{2},\ldots,s^{t},s^{T}\right]=s^{1:T}$, where $s^{t}$ represents the reasoning step at timestamp $t$.
-We constrain each reasoning step should within a set of candidates $\mathcal{S}^{t}$ to control the computation efficiency.
-We construct the reasoning step candidates $\mathcal{S}^{t}$ based on the scoring function defined in Eq.
-3 by selecting the top-$m$ reasoning steps with the highest scores.
-At each timestamp $t$, we leverage the LLMs to select the top-$k$ reasoning steps from the reasoning step candidates $\mathcal{S}^{t}$.
-The beam search process at timestamp $t$ is modeled as follows:
-$\mathcal{H}_{t}=\text{Top}_{k}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},w):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}^{t}\right\}$ (4)
-where $\mathcal{H}_{t-1}$ denote the reasoning path up to the previous timestamp $t-1$ and $s^{t}$ refers to the current reasoning step.
-$\mathcal{S}^{t}$ refers to the set of possible next-step candidates retrieved by the Path-RAG at timestamp $t$.
-$\text{LM}(s^{t}|q,s^{1:t-1},w)$ refers to the language models prediction for the next step $s^{t}$ given the previous sequence $s^{1:t-1}$, the query $q$, and the planning context $w$.
-The operator $\oplus$ appends $s^{t}$ to the current path $h$, and $\text{Top}_{k}$ selects the top-$k$ reasoning paths (with $k$ controlling the beam search width, where a larger $k$ typically yield better performance (see the analysis in Figure 2)).
-Deductive Verification.
-To terminate the reasoning path extension process at the right point, we propose to leverage the deductive reasoning capabilities of LLMs (Ling et al., 2024; Huang & Chang, 2023) as a criterion $C(x,s^{t},s^{1:t-1})\in\left\{0,1\right\}$ to automatically guide the decision making of LLMs in Equation 4.
-Specially, we first leverage LLMs to convert the user query to a declarative statement $q^{\prime}$ (as shown in Appendix F.6) and then use LLMs to judge whether it can be deduced from current reasoning step $s^{t}$ and the previous reasoning steps $s^{1:t-1}$.
-We utilize the same backend LLM with prompt referred in Section F and define the deductive-verification criteria as follows:
-$C(q^{\prime},s^{t},s^{1:t-1})=\begin{cases}1,&\text{if }q^{\prime}\text{ can be deduced from }s^{t}\text{ and }s^{1:t-1},\\ 0,&\text{otherwise.}\end{cases}$ (5)
-The overall goal of DVBS model can be represented as:
-$\mathcal{H}_{t}=\text{Top}_{B}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},w):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\text{ and }C(q^{\prime},s^{t},s^{1:t-1})=1\right\}$ (6)
-This criterion is an essential component of our method for leveraging LLMs to iteratively validate each step of reasoning, ensuring that each step logically follows from the preceding steps and aligns with the original query.
-The function effectively signals when to terminate further reasoning, enhancing accuracy and minimizing unnecessary computational efforts.
-3 Experiments
-During the experiment section, we aim to answer the following research questions: RQ1: How does FiDeLiS perform compare with existing baselines on KGQA tasks?
-RQ2: How does the design of each individual component contribute to the overall performance of FiDeLiS?
-RQ3: How robustness of FiDeLiS perform under varying conditions?
-RQ4: How about the efficiency of FiDeLiS?
-All experiment settings are detailed in Appendix D due to page constraints.
-Table 1: Comparison of FiDeLiS with baseline methods and different backbone LLMs.
-We replicate the outcomes of ToG and RoG, and retrieve other baseline results directly from the original paper.
-We utilize 5 demonstrations as our default setting for FiDeLiS, ToG, Few-shot, and CoT.
-The experiment results of open-source models can be found in Table 10.
+Next, we construct candidate reasoning steps defined in §2 using the retrieved candidate entities $\mathcal{E}_{m}$ and relations $\mathcal{R}_{m}$.
+To guide the selection of potential candidate, we propose a scoring function that combines semantic similarity with the KG’s structural connectivity.
+First we define the base score function $S_{0}$ that captures only the semantic alignment of the candidate with the query as: $S_{0}((r,e))=S_{\mathrm{rel}}(r)+S_{\mathrm{ent}}(e)$, where $S_{\mathrm{ent}}(e)$ and $S_{\mathrm{rel}}(r)$ represents the cosine similarity between the entity/relation to the query respectfully.
+To account for the KG’s structural connectivity, i.e., the potential for a candidate to lead to fruitful next steps, we incorporate information from the next-hop candidates and define the overall scoring function as Eq 2:
+$S((r,e))=S_{0}((r,e))+\alpha\max_{\forall(r_{j},e_{j})\in N(e)}S_{0}((r_{j},e_{j}))$ (2)
+Where $N(e)$ denotes the set of candidate relation-entity pairs reachable from entity $e$ within one hop in the KG.
+$\alpha$ is a hyper-parameter that balances the immediate semantic relevance (captured by $S_{0}((r,e))$ with the candidate’s potential for future connectivity (captured by the maximum next-hop score).
+A higher $\alpha$ favors candidates with long-term benefits, even if they seem sub-optimal initially, while a lower $\alpha$ emphasizes immediate rewards, potentially overlooking future impacts.
+We verify the effectiveness of this new scoring function in Table 3 and append the tuning results of hyper-parameter $\alpha$ in Figure 5.
+3.2 Deductive-Verification Beam Search
+The objective of DVBS is two-fold: (1) to provide a step-wise beam search for exploring verifiable reasoning paths from KG based on candidates constructed by Path-RAG (§3.1), and (2) to verify each reasoning step based on deductive reasoning Ling et al.
+(2023) to ensure each step logically follows the previous steps and supports the user query.
+Compared with existing methods like Sun et al.
+(2023) and Ma et al.
+(2024), while we both consider treat LLMs as agents that iteratively interact with KGs to find reasoning paths and answers, our method leverage deductive reasoning to ensure each reasoning steps are logically connected and only halt the search if the question can be deduced based on the reasoning paths.
+Based on our robustness analysis in §4.3, DVBS demonstrate higher ratio of valid reasoning paths and can prevent issues of either premature stopping Huang et al.
+(2017) or excessive continuation of reasoning path extension.
+The DVBS consists of three steps and we detail the workflow as follows:
+Plan Generation.
+Inspired by the recent works regarding planning capabilities of LLMs Zhang et al.
+(2023); Kagaya et al.
+(2024), we prompt an LLM to generate the planning steps for answering the user query, denoted as $w$.
+This step is designed to provide more hints for subsequent LLM decision making process.
+Even this step is more like an engineering trick, we find that it may unlock some of the capabilities of LLM to do “higher-order” thinking.
+By including more hints in the prompt, the LLM tends to make more accurate and deterministic decisions during beam search, thus improving the quality of the traversed reasoning paths.
+(See Table 2 for ablation analysis).
 Backend Models Methods WebQSP CWQ CR-LT
@@
 CoT 57.42 54.72 43.21 35.85 37.42
-Prompting - LLM Only gpt-4-turbo Zero-shot 62.32 59.71 42.71 37.93 37.74
+[…]
+Prompting - LLM Only gpt-4-turbo IO 62.32 59.71 42.71 37.93 37.74
 Few-shot 68.65 62.71 51.52 43.70 43.61
 CoT 72.11 65.37 53.51 44.76 45.42
-Finetuning - LLM + KG NSM (He et al., 2021) 74.31 - 53.92 - -
-CBR-KBQA (Das et al., 2021) - - 67.14 - -
-DeCAF (Yu et al., 2022) 82.1 - 70.42 - -
-KD-CoT (Wang et al., 2023a) 73.7 50.2 50.5 - -
-RoG (Luo et al., 2024) 83.15 69.81 61.39 56.17 60.32
-Prompting - LLM + KG gpt-3.5-turbo ToG (Sun et al., 2023) 75.13 72.32 57.59 56.96 62.48
+[…]
+Finetuning - LLM + KG NSM He et al. (2021) 74.31 - 53.92 - -
+CBR-KBQA Das et al. (2021) - - 67.14 - -
+DeCAF Yu et al. (2023) 82.1 - 70.42 - -
+KD-CoT Wang et al. (2023) 73.7 50.2 50.5 - -
+RoG Luo et al. (2024a) 83.15 69.81 61.39 56.17 60.32
+[…]
+Prompting - LLM + KG gpt-3.5-turbo ToG Sun et al. (2023) 75.13 72.32 57.59 56.96 62.48
+KAPING Baek et al. (2023) 72.42 65.12 53.42 50.32 -
 FiDeLiS 79.32 76.78 63.12 61.78 67.34
-Prompting - LLM + KG gpt-4-turbo ToG (Sun et al., 2023) 81.84 75.97 68.51 60.20 67.24
+[…]
+Prompting - LLM + KG gpt-4-turbo ToG Sun et al. (2023) 81.84 75.97 68.51 60.20 67.24
 FiDeLiS 84.39 78.32 71.47 64.32 72.12
-3.1 RQ1: KGQA Performance Comparison
-Table 1 clearly showcases the comparative effectiveness of FiDeLiS, against different baselines.
-It shows that FiDeLiS outperforms all baselines using gpt-4-turbo model, even for strong baselines being fine-tuned, such as DeCAF (Yu et al., 2022) and RoG (Luo et al., 2024).
-Across all prompting-based methods, gpt-4-turbo achieves higher performance compared to gpt-3.5-turbo, especially on the CWQ dataset.
-It indicates that gpt-4-turbo has a better understanding and processing capability for complex queries.
-The CR-LT dataset seems more challenging compared with WebQSP and CWQ, as implied by the consistently lower accuracy scores.
-However, FiDeLiS shows consistent enhancements compared to other baselines, demonstrating its capability to handle long-tail entities by referring to the knowledge from KGs and processing more complex queries effectively.
-3.2 RQ2: Ablation Study of FiDeLiS.
-Table 2 demonstrates the ablation study of FiDeLiS using the gpt-3.5-turbo-0125 model across the WebQSP, CWQ, and CR-LT datasets.
-It highlights the critical importance of both the Path-RAG and DVBS components.
-Removing Path-RAG, whether by employing a vanilla retriever or the think-on-graph (ToG) approach (Sun et al., 2023), results in substantial performance declines, particularly a 6.97% drop in Hits@1 on WebQSP and a 6.01% decrease on CWQ, underscoring the necessity of an effective retrieval mechanism.
-Similarly, ablating parts of DVBS, especially the beam search component, leads to significant reductions, with an 18.97% decrease in Hits@1 on WebQSP and a 13.34% drop on CWQ, indicating that beam search is vital for maintaining high accuracy.
-Other DVBS subcomponents, such as the deductive verifier and planning, also contribute notably to performance, albeit to a lesser degree.
-Overall, the findings demonstrate that the integrated Path-RAG and DVBS frameworks are essential for FiDeLiS’s robust performance, with their removal causing marked decreases in accuracy and Hits@1 scores across all evaluated datasets.
-Table 2: Ablation Studies of FiDeLiS using model gpt-3.5-turbo-0125.
-$\Delta$ refers to the performance gap between each component and the entire method.
-The three largest performance gaps on each dataset are highlighted in green, with darker shades denoting more significant differences.
+Table 1: Comparison of FiDeLiS with baseline methods and different backbone LLMs.
+We replicate the outcomes of ToG and RoG, and retrieve other baseline results directly from the original paper.
+We utilize 5 demonstrations as our default setting for FiDeLiS, ToG, Few-shot, and CoT.
+The experiment results of open-source models can be found in Table 11.
+Beam Search.
+We then construct the multi-step reasoning paths by iteratively extending partial paths using a beam search strategy.
+At each time step $t$, we use an LLM as agent to select one reasoning step $s^{t}$ from a candidates set $\mathcal{S}^{t}$3.1) conditioned on (1) the likelihood of each candidate $s_{i}\in\mathcal{S}^{t}$, (2) the user query $q$, (3) the history of previous steps $s^{1:{t-1}}$, and (4) planning context $w$3.2), denoted as $\mathrm{LM}(s^{t}|q,s^{1:t-1},w,\mathcal{S}^{t})$.
+Instead of exploring every possibility, we retain only the top-$k$ scoring paths from the previous beam $\mathcal{H}_{t-1}$ and extend them by appending candidate steps.
+The overall process can be expressed as Eq 3:
+$\mathcal{H}_{t}=\mathrm{Top}_{k}\big{\{}h\oplus\mathrm{LM}(s^{t}|q,s^{1:t-1},w,\mathcal{S}^{t}):h\in\mathcal{H}_{t-1}\big{\}}$ (3)
+where $\oplus$ denotes the concatenation of the current path $h$ with the selected candidate step $s^{t}$.
+The beam search strategy enable efficiently navigate the vast space of potential reasoning paths while concentrating on the most promising ones.
+While beam search, by its nature, can incur high computational costs and latency due to multiple rounds of LLM interactions.
+Our retrieval module Path-RAG mitigate this issue by constraining candidate set $\mathcal{S}^{t}$ at each time step $t$ to a narrow, high-quality subset rather than requiring the LLM to consider all available options.
+This targeted retrieval not only reduces the number of candidates to evaluate at each step but also increases the likelihood of selecting relevant reasoning steps, thereby enabling efficient traversal of KGs at scale.
+Find more discussion regarding efficiency of FiDeLiS in §4.4 and Appendix §B.
+Deductive Verification.
+To ensure that each reasoning step logically follows from its predecessors and adequately supports the original query, we leverage the deductive reasoning capabilities of LLMs as a verification criterion Ling et al.
+(2023) for the beam search process.
+We first convert the user query $q$ into a clear declarative statement $q^{\prime}$, which encapsulates its logical intent and allows the LLM to operate on a well-defined logical target (See the concrete example in §C.6).
+Next, during the beam search, candidate reasoning step $s^{t}$ are appended to the history $s^{1:t-1}$ to form potential reasoning paths.
+For each candidate, we then invoke two deductive verification checks, $C_{\mathrm{global}}$ and $C_{\mathrm{local}}$ (the prompts are given in §C.2).
+Only those candidates that pass local verification, indicating that the new step maintains logical consistency with the established context, are retained in the beam search process.
+Once the candidates pass both verification indicate that the user query $q$ can be deduced based on the retained reasoning paths $s^{1:t}$ and the beam search progress should be halted.
+Global Verification: $C_{\mathrm{global}}(s^{1:t-1},s^{t})$ returns 1 if $(s^{t}\land s^{1:t-1})\models q^{\prime}$, and 0 otherwise.
+Local Verification: $C_{\mathrm{local}}(s^{1:t-1},s^{t})$ returns 1 if $s^{t}$ logically follows from $s^{1:t-1}$, and 0 otherwise.
+By integrating this verification into the beam search offers several benefits: it (1) enhances the robustness and validity of the final answer by enforcing logical coherence at every step, (2) reduces computational overhead by pruning unpromising paths early, and (3) mitigates risks such as premature termination or excessive extension of the reasoning process.
+We provide a concrete example of the deductive verification process in §C.6 and the complete DVBS algorithm in Algorithm 2.
+4 Experiments
+In this section, we focus on verifying FiDeLiS from four perspectives as follows: (1) comparison results with other baselines over KGQA; (2) ablation study; (3) robustness analysis and (4) efficiency analysis.
+We provide all the experiment settings in Appendix A due to page constraints.
+The prompts for plan generation, beam search and deductive verification can be found in §C.
+4.1 Main Results
+In Table 1, we compare the performance of different methods with various backbend LLMs across three datasets.
+We found that LLM + KG approaches generally outperform LLM-only methods (Zero-shot, Few-shot, and CoT) by a wide margin, indicating the significant benefit of incorporating KGs into LLM reasoning.
+In the LLM + KG category, FiDeLiS stands out as the best-performing method across all datasets, particularly when paired with GPT-4-turbo.
+For example, on WebQSP, FiDeLiS achieves 84.39% Hits@1 and 78.32% F1, surpassing ToG (81.84% Hits@1, 75.97% F1) and RoG (83.15% Hits@1, 69.81% F1).
+This improvement is consistent across other datasets, and even compared with some finetuning methods like DeCAF and RoG, FiDeLiS as a training-free method still demonstrate better performance.
+The consistent performance of FiDeLiS highlights its effective use of both the KG and LLM, as well as its optimization of hyper-parameters like beam width and depth.
+Overall, the results illustrate that FiDeLiS, leveraging advanced LLMs like GPT-4-turbo and KG-based reasoning, sets a new standard for performance in KG-related tasks.
 Ablation Setting Components WebQSP CWQ CR-LT
-Hits@1 (%) $\Delta$ Hits@1 (%) $\Delta$ Acc (%) $\Delta$
-No ablation FiDeLiS 79.32 0.00 63.12 0.00 67.34 0.00
-w/o Path-RAG using vanilla retriever 72.35 6.97 57.11 6.01 59.78 7.56
-using ToG 75.11 4.21 59.47 3.65 63.47 4.07
-w/o DVBS w/o last step reasoning 75.68 3.64 59.45 3.67 63.72 3.62
-w/o planning 76.23 3.09 60.14 2.98 64.13 3.21
-w/o beam-search 60.35 18.97 49.78 13.34 61.87 5.47
-w/o deductive-verifier 74.13 5.19 57.23 5.89 63.89 3.45
-3.3 RQ3: Robustness Analysis
-Comparison using different embedding backbones.
-Table 3 demonstrates the performance of the Path-RAG using various embedding models as backbones.
-It shows that the integration of the Openai-Embedding Model significantly enhances the performance across all datasets, indicating its superior capability in understanding and processing natural language queries, which leads to more effective retrieval.
-For instance, the Openai-Embedding Model increases the Hits@1 by 13.04% in the WebQSP dataset and 14.73% in the CWQ dataset compared to the base BM25 (Robertson et al., 2009) model.
-Furthermore, this improvement is consistent and highlights the feasibility and potential of pairing more powerful language models with structured reasoning frameworks to address complex information retrieval and question answering challenges.
-Table 3: Performance FiDeLiS using different Path-RAG embedding backbone models. ‘Openai-Embedding Model’ refers to text-embedding-3-small, the recent embedding model released from OpenAI.
+Hits@1 (%) Hits@1 (%) Acc (%)
+No ablation FiDeLiS 79.32 63.12 67.34
+[…]
+w/o Path-RAG using vanilla retriever 72.35 57.11 59.78
+using ToG 75.11 59.47 63.47
+[…]
+w/o DVBS w/o last step reasoning 75.68 59.45 63.72
+w/o planning 76.23 60.14 64.13
+w/o beam-search 60.35 49.78 61.87
+w/o deductive-verifier 74.13 57.23 63.89
+Table 2: Ablation Studies of FiDeLiS using model gpt-3.5-turbo-0125. $\Delta$ refers to the performance gap between each component and the entire method.
+4.2 Ablation Study
+Table 2 demonstrates the ablation study of FiDeLiS using the gpt-3.5-turbo-0125 model, highlighting the contributions of individual components (Path-RAG and DVBS) to overall performance.
+We conduct the ablation of the Path-RAG by replacing it with either a vanilla retriever or ToG Sun et al.
+(2023) as retriever.
+We find that using ToG shows slight improvements over the vanilla retriever but remains below using Path-RAG.
+Ablating DVBS components also leads to performance declines, particularly when beam search is removed, causing Hits@1 on WebQSP to drop sharply to 60.35%.
+The deductive verifier and last-step reasoning show moderate but noticeable impacts on performance.
+The effects are less pronounced on CR-LT, suggesting it is more tolerant of simpler methods.
+Overall, the results confirm the critical roles of Path-RAG and DVBS, especially beam search, in ensuring robust and accurate reasoning across domains.
 Methods Backbones WebQSP CWQ CR-LT
 Hits@1 (%) Hits@1 (%) Acc (%)
-Vanilla Retriever w/ BM25 (Robertson et al., 2009) 58.31 48.39 50.73
-w/ SentenceBert (Reimers & Gurevych, 2019) 62.74 50.14 51.80
-w/ E5 (Wang et al., 2022a) 68.42 52.84 54.31
-w/ Openai-Embedding-Model 72.35 57.11 59.78
-Path-RAG w/ BM25 (Robertson et al., 2009) 70.34 56.11 58.77
-w/ SentenceBert (Reimers & Gurevych, 2019) 73.45 58.41 60.45
-w/ E5 (Wang et al., 2022a) 77.93 62.74 65.23
-w/ Openai-Embedding-Model 79.32 63.12 67.34
-Comparison using different searching widths and depths.
-To investigate how the search width and depth affect the performance of FiDeLiS, we have performed experiments with settings ranging from depths of 1 to 4 and widths of 1 to 4.
-The results, presented in Figure 2, show that performance generally improves as search depth and width increase.
-However, performance begins to decline when the search depth exceeds 3 for both the WebQSP and CWQ datasets.
-This decline is attributed to the fact that only a small fraction of questions in these datasets require reasoning at depths greater than 3.
-Additionally, FiDeLiS utilizes stricter rules for ending the search process, which can preemptively terminate misguided searches.
-Contrastingly, increasing the search width consistently shows potential for improved performance, suggesting benefits from broader exploration.
-However, considering the computational costs, which rise linearly with depth, we have chosen to set both the default beam width and depth to 4 for all experiments unless stated otherwise.
-This default setting aims to balance performance gains with computational efficiency.
-(a) RD over CWQ
-(b) RD over WebQSP
-(c) BW over CWQ
-(d) BW over WebQSP
-Figure 2: Performance (Hits@1) of FiDeLiS with different beam search widths (BW) and reasoning depths (RD) over WebQSP and CWQ.
-Comparison between Path-RAG and Vanilla Retriever.
-To verify the robustness of the Path-RAG compared to vanilla retrievers.
-We conduct the experiments as shown in Figure 3(a) and 3(b).
-We calculate the coverage ratio (CR) of the retrieved reasoning paths and the ground-truth reasoning paths as follows: $\text{CR}=\left(\frac{N_{\text{retrieved}}\cap N_{\text{ground-truth}}}{N_{\text{ground-truth}}}\right)\times 100\%$, where $N_{\text{retrieved}}$ is the number of retrieved reasoning paths candidates, and $N_{\text{ground-truth}}$ is the total number of ground-truth reasoning paths.
-We set the vanilla retriever as the baseline, specifically, we concatenate each entity with its corresponding relation to form a reasoning step.
-We then calculated the cosine similarity between the embeddings of each reasoning steps and the query embeddings to select candidate paths.
-We observe that compared to the vanilla retriever, Path-RAG achieves a higher CR value and shows better alignment with the ground-truth paths.
-It demonstrates that our method can better leverage the connections that are often missed by simpler retrieval models.
-This feature is critical in subsequent LLM to focus on the information of interest and arrive at a correct and reasonable answer.
-(a) Coverage Ratio over CWQ
-(b) Coverage Ratio over WebQSP
-(c) Error Analysis over WebQSP
-Figure 3: (a)-(b): Empirical study regarding the coverage ratio of the retrieved reasoning paths.
-(c): Error analysis of the validation of whole paths generated from RoG (Luo et al., 2024) over WebQSP.
-Error analysis regarding whole path generation.
-To verify the faithfulness of our stepwise method, we conduct the error analysis regarding the whole reasoning path generation.
-Specifically, we conduct an analysis of the validity of the whole reasoning path generation using the baseline methods RoG (Luo et al., 2024).
-The result is shown in Figure 3(c).
-The definition of the metric validity ratio (VR) is the ratio of reasoning steps that existed in the knowledge graph to the total number of the reasoning steps in the output reasoning path: $\text{VR}=\left(\frac{N_{\text{valid-steps}}}{N_{\text{all-steps}}}\right)\times 100\%$, where $N_{\text{valid-steps}}$ is the number of reasoning steps that existed in the knowledge graph, and $N_{\text{all-steps}}$ is the number of all the reasoning steps in the reasoning path.
-It show that only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG.
-This illustrates that the reasoning steps generated by language models offer few guarantees about feasibility especially when multiple consecutive steps are combined into a reasoning path.
-Table 4: Average depths of the returned reasoning paths. GT refers to ground-truth reasoning paths.
+Vanilla Retriever w/ BM25 58.31 48.39 50.73
+w/ SentenceBert 62.74 50.14 51.80
+w/ E5 68.42 52.84 54.31
+w/ Openai-Emb 72.35 57.11 59.78
+[…]
+Path-RAG w/ BM25 70.34 56.11 58.77
+w/ SentenceBert 73.45 58.41 60.45
+w/ E5 77.93 62.74 65.23
+w/ Openai-Emb 79.32 63.12 67.34
+Table 3: Performance of FiDeLiS with various embedding methods. * refers to text-embedding-3-small from OpenAI. We detail the tested embedding methods in §A.3.
+4.3 Robustness Analysis
+Robustness of Path-RAG.
+Table 3 presents the performance of FiDeLiS compared to a vanilla retriever with different embedding methods.
+The results consistently show that FiDeLiS outperforms the vanilla retriever irrespective of the underlying embedding strategy.
+For instance, with Openai-Emb, the vanilla retriever achieves 72.35% on WebQSP, whereas Path-RAG reaches 79.32%, indicating a notable improvement.
+Similar performance gains are observed with the other embeddings.
+These improvements suggest that integrating graph connections can enhance retrieval effectiveness by providing more informative and contextually relevant information, thereby bolstering the overall robustness and accuracy of the method.
+Effectiveness of Path-RAG.
+We verify the effectiveness of the retrieval module Path-RAG with two baselines: (1) a vanilla retriever and (2) KAPING Baek et al.
+(2023) method.
+The vanilla retriever concatenates each entity with its relation to form a reasoning step and selects candidates based on cosine similarity with the query embeddings.
+In contrast, KAPING Baek et al.
+(2023) converts each triple into text and retrieves the top-$K$ similar triples based on semantic similarity.
+We quantify the retrieval performance using the coverage ratio (CR), defined as the percentage of the ground-truth reasoning steps being retrieved throughout the reasoning path extension (i.e., $\mathrm{CR}=\frac{N_{\mathrm{retrieved}}\cap N_{\mathrm{ground-truth}}}{N_{\mathrm{ground-truth}}}$).
+Table 4 illustrate the experimental setup and corresponding results.
+We find that compared with the baselines, our Path-RAG achieves a higher CR value and aligns better with the ground-truth paths.
+It demonstrates superior ability to capture connections that simpler retrieval models may overlook.
+This advantage is critical for guiding subsequent LLM processing toward relevant information, ultimately yielding more accurate and coherent answers.
+Method Depth = 1 Depth = 2 Depth $>$ 3
+Vanilla Retriever 59.34 52.17 47.31
+KAPING Baek et al. (2023) 65.72 60.41 53.11
+Path-RAG w/ keywords 72.61 69.38 62.78
+Path-RAG w/o keywords 68.78 ($\downarrow$ 3.83) 65.27 ($\downarrow$ 4.11) 57.13 ($\downarrow$ 5.65)
+Table 4: Analysis of the CR of reasoning paths over CWQ.
+Methods WebQSP (hits@1) CWQ (hits@1)
+Deductive Verification 79.32 63.12
+Adequacy Verification (used in ToG) 74.13 57.23
+Logit-based Scoring 73.47 54.78
+Table 5: Analysis of different verification methods.
+Path Error Analysis.
+To verify the faithfulness of our step-wise method, we conduct an error analysis regarding the whole reasoning path generation using RoG Luo et al.
+(2024a).
+We quantify the validity of reasoning path using validity ratio (VR), which is defined as the ratio of reasoning steps that existed in the KG to the total number of the reasoning steps in the output reasoning path (i.e., $\mathrm{VR}=\frac{N_{\mathrm{valid-steps}}}{N_{\text{all-steps}}}$).
+As shown in Figure 3, only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG.
+This illustrates that the reasoning steps generated offer few guarantees about feasibility especially when multiple consecutive steps are combined into a reasoning path.
+While our method leverage step-wise verification to ensure that each of the reasoning step exist in the KG and logically connected.
+Figure 3: Analysis of reasoning errors in RoG Luo et al. (2024a) over WebQSP.
 Method WebQSP CWQ CR-LT
 GT 2.3 3.2 4.7
+[…]
 ToG 3.1 4.1 5.2
 FiDeLiS 2.4 2.8 4.6
+Table 6: Average depths of the generated reasoning paths. GT refers to ground-truth reasoning paths.
 Effectiveness of Deductive-Verification.
-To verify the effectiveness of deductive-verification, we calculate the average depths of the grounded reasoning path from ToG and our methods in Table 4.
-It shows that FiDeLiS consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to ToG.
-This implies that our method may offer more precise termination signals and potentially more accurate reasoning paths compared to baselines.
-Case Study.
-To verify whether FiDeLiS can conduct faithful reasoning on real cases, we also present a case study in Table 5.
-We can find that FiDeLiS illustrates better performance in identifying multiple possible governmental forms—Theocracy, Unitary state, and Islamic republic established in Iran in 1979, as compared to existing models such as GPT + CoT, RoG, and ToG.
-This observation highlights the importance of recognizing the complexity of Iran’s government system, which combines elements of religion and democracy.
-Different models handle this complexity with varying levels of detail.
-For example, GPT + CoT and RoG mainly identified Iran as an ‘Islamic Republic’, a change that occurred after the 1979 Iranian Revolution, but they missed other possible descriptions.
-The ToG model even gives mixed results, such as simply naming it ‘Iran’, which shows confusion in its analysis.
-However, our method shows a better understanding of the context and provides a more detailed and accurate set of answers.
-This illustrates FiDeLiS strength and efficiency in dealing with intricate questions.
-Table 5: Case study of FiDeLiS.
-The wrong answers are highlighted with red color, and the correct answers are highlighted with blue color.
-In this case, FiDeLiS scored an F1 of 0.857, outperforming the RoG method (0.4).
+To verify the effectiveness of deductive-verification mentioned in §3.2, we calculate the average depths of the generated reasoning paths as shown in Table 6.
+We find that by considering deductive verification, it consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to baseline.
+This implies that FiDeLiS may offer more precise termination signals and potentially more accurate reasoning paths.
+We also compare deductive-verification methods with other baselines in Table 5, like logit-based scoring that assign softmax probability scores to determine the endpoint of beam search process, and adequacy verification used in ToG Sun et al.
+(2023).
+Experiments show higher accuracy with deductive verification compared to adequacy verification and logit-based scoring, demonstrating its effectiveness in enhancing reasoning accuracy.
+Dataset Method Hits@1 Runtime Token ##\##
+WebQSP FiDeLiS (ours) 79.32 43.83 2,452 10.7
+   w/o Path-RAG using vanilla retriever 72.35 48.37 2,873 10.7
+   w/o Path-RAG using ToG 75.11 74.26 6,437 10.7
+FiDeLiS (ours) - GPT-4o 81.17 37.82 2,452 10.7
+FiDeLiS (ours) - GPT-4o-mini 76.48 24.31 2,452 10.7
+[…]
+CWQ FiDeLiS (ours) 63.12 74.59 2,741 15.2
+   w/o Path-RAG using vanilla retriever 57.11 78.41 3,093 15.2
+   w/o Path-RAG using ToG 59.47 132.59 5,372 15.2
+FiDeLiS (ours) - GPT-4o 65.33 50.12 2,741 15.2
+FiDeLiS (ours) - GPT-4o-mini 58.34 42.54 2,741 15.2
+Table 7: Runtime efficiency of FiDeLiS per question.
 Question What form of government is in the country that uses the Iranian Rail and was established in 1979?
@@
 So, the government form is the Islamic Republic of Iran, established in 1979 following the Iranian Revolution.
+[…]
 RoG location.country.form_of_government, location.location.containedby location.country.form_of_government’ location.location.containedby $\to$ government.form_of_government.countries The results are Islamic republic.
+[…]
 ToG Path-1: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic $\to$ government.form_of_government.countries $\to$ Iran Path-2: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy $\to$ government.form_of_government.countries $\to$ Iran Path-3: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state $\to$ government.form_of_government.countries $\to$ Iran Based on the reasoning paths, the result is Iran.
+[…]
 FiDeLiS Path-1: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic Path-2: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy Path-3: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state Based on the reasoning paths, the results are Theocracy, Unitary state, Islamic republic.
-3.4 RQ4: Efficiency of FiDeLiS
+Table 8: Case study of FiDeLiS. We highlight the wrong answers with red color, and correct answers with blue color.
+This finding is further supported by a case study regarding a complex question of Iran’s government system, which blends elements of religion and democracy as shown in Table 8.
+While baseline methods such as GPT + CoT and RoG predominantly identified Iran as an “Islamic Republic” and ToG produce mixed responses, our approach—enhanced by deductive verification—delivers a reasoning path that is both concise and context-aware.
+The proposed verification mechanism not only streamlines the reasoning process but also ensures comprehensive coverage of grounded answers, demonstrating FiDeLiS’s strength in handling intricate questions.
+4.4 Efficiency Analysis
 To investigate the runtime efficiency and cost efficiency of FiDeLiS, we present a comparison regarding the average runtime, average token usage, average times of LLM calling per question in Table 7.
-We found that (1) our method shows superior efficiency compared to the ToG (which is also training-free), by reducing approximately 1.7x runtime costs.
+We find that (1) our method shows superior efficiency compared to the ToG (which is also training-free), by reducing approximately 1.7x runtime costs.
 (2) Path-Rag component is critical in enhancing both the accuracy and efficiency of the model.
 Its ability to constrain potential path candidates effectively reduces unnecessary computational overhead, leading to quicker and more accurate results.
-Table 6: Runtime efficiency of FiDeLiS per question.
-Dataset Method Hits@1 (%) Avg Runtime (s) Avg Token Usage Avg LLM calling
-WebQSP FiDeLiS (ours) 79.32 43.83 2,452 10.7
-   w/o Path-RAG using vanilla retriever 72.35 48.37 2,873 10.7
-   w/o Path-RAG using ToG 75.11 74.26 6,437 10.7
-FiDeLiS (ours) - GPT-4o 81.17 37.82 2,452 10.7
-FiDeLiS (ours) - GPT-4o-mini 76.48 24.31 2,452 10.7
-CWQ FiDeLiS (ours) 63.12 74.59 2,741 15.2
-   w/o Path-RAG using vanilla retriever 57.11 78.41 3,093 15.2
-   w/o Path-RAG using ToG 59.47 132.59 5,372 15.2
-FiDeLiS (ours) - GPT-4o 65.33 50.12 2,741 15.2
-FiDeLiS (ours) - GPT-4o-mini 58.34 42.54 2,741 15.2
-To address concern regarding our method’s potential application in real-time scenarios, we also tested our method using faster and more advanced LLMs.
+To address concern regarding our method’s potential application in real-time scenarios, we also test our method using faster and more advanced LLMs.
 Table 7 shows that our method could be further accelerated with newer, faster models like GPT-4o or GPT-4-mini.
 The potential of the ongoing advancements in LLMs are expected to further enhance the scalability and efficiency of FiDeLiS, making it a practical development in challenging environments.
-More detailed analysis of bottleneck of computation of FiDeLiS can be further found in Appendix C.
+More detailed analysis of bottleneck of computation of FiDeLiS can be further found in Appendix B.
 5 Related Work
-KG-enhanced LLM.
-KGs have advantages in dynamic, explicit, and structured knowledge representation and techniques combining LLMs with KGs have been studied (Pan et al., 2024).
-Early studies (Luo et al., 2024; Yu et al., 2022) embed structured knowledge from KGs into the underlying neural networks during the pretraining or fine-tuning process.
-However, the reasoning steps generated by LMs are observed to be prone to errors, which could be non-existent in the knowledge graph, and can subsequently lead to incorrect reasoning during inference.
-Also, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating (Hu et al., 2023).
-Another solutions are to keep the reasoning model as an agent to explore the external structure knowledge source (Sun et al., 2023; Gu et al., 2023; Jiang et al., 2023b; Wang et al., 2023a) ToG (Sun et al., 2023) directly employs LLMs to output scores for candidate selections from the KG.
-This model operates at the relation-entity level of the KG, aiming to identify relevant triples that aid the LLM in making accurate and responsible final answer predictions.
-Knowledge Graph-based Question Answering.
-To integrate LLMs for KGQA, retrieval-augmented methods aim to retrieve the relative facts from the KGs to improve the reasoning performance (Li et al., 2023; Karpukhin et al., 2020).
-Recently, UniKGQA (Jiang et al., 2022) which unifies the graph retrieval and reasoning process into a single model with LLMs, achieves STOA performance.
-Semantic parsing methods convert the question into a structural query (e.g., SPARQL) by LLMs, which can be executed by a query engine to reason the answers on KGs (Sun et al., 2020; Lan & Jiang, 2020).
-However, these methods heavily rely on the quality of generated queries.
-If the query is not executable, no answers will be generated.
-DECAF (Yu et al., 2022) combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance.
+LLM Reasoning & Role of KGs.
+Large language models (LLMs) demonstrate impressive capabilities in reasoning tasks but often generate hallucinated or factually incorrect outputs, particularly in complex, multi-step scenarios Huang et al.
+(2025); Li et al.
+(2024).
+This unreliability reduces their effectiveness in knowledge-intensive applications.
+Knowledge graphs (KGs) have emerged as a solution by offering structured, verifiable data that supports transparent and multi-hop reasoning Sui et al.
+(2024).
+Unlike document-based retrieval-augmented generation approaches, KGs provide direct access to relational facts, enhancing both interpretability and traceability Chen et al.
+(2024).
+KG-enhanced LLM Reasoning.
+KG-enhanced reasoning methods are generally categorized into retrieval-based and agent-based models.
+Retrieval-based approaches, such as DeCAF Yu et al.
+(2023), rely on text-based retrieval to select relevant information from KGs and jointly generate answers and logical forms, but their performance can degrade without precise retrieval mechanisms.
+In contrast, agent-based models, like ToG Sun et al.
+(2023), iteratively explore reasoning paths but suffer from high computational overhead.
+To address these limitations, recent methods like RoG Luo et al.
+(2024a) and GCR Luo et al.
+(2024b) have sought to integrate KG structure into LLM training or decoding to improve reasoning fidelity and explanation generation.
+To improve the faithfulness of the LLM reasoning, KD-CoT Wang et al.
+(2023) verifies sub-reasoning steps through external KGs to prevent errors during inference, while NSM He et al.
+(2021) employs a teacher-student architecture to learn intermediate supervision signals that guide reasoning.
 6 Conclusion
@@
 The Path-RAG module and the use of deductive reasoning as a calibration tool effectively guide the reasoning process, leading to more accurate knowledge retrieval and prevention of misleading reasoning chains.
-Extensive experiments demonstrate that our method, being training-free, not only reduces computational costs but also offers superior generality, consistently outperforming established strong baselines across three distinct benchmarks.
+Extensive experiments demonstrate that our method, being training-free, not only reduces computational costs but also offers superior generality.
 We believe this study will significantly benefit the integration of LLMs and KGs, or serve as an auxiliary tool to enhance the interpretability and factual reliability of LLM outputs.
-Appendix A Definitions
-A.1 Knowledge Graph-based Question Answering (KGQA)
-In this work, we mainly focus on the question answering task based on the knowledge graph.
-Knowledge Graphs (KGs) contain abundant factual knowledge in the form of a set of triples: $\mathcal{G}=$$\left\{\left(e,r,e^{\prime}\right)\mid e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\right\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations, respectively.
-Knowledge Graph Question Answering (KGQA) is a typical reasoning task based on KGs.
-Given a natural language question $q$ and a $\mathrm{KG}$ $\mathcal{G}$, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ based on knowledge from $\mathcal{G}$, i.e., $a=f(q,\mathcal{G})$.
-Following previous works Luo et al.
-(2024; 2023), we assume the entities $e_{q}\in\mathcal{T}_{q}$ mentioned in $q$ and answers $a\in\mathcal{A}_{q}$ are labeled and linked to the corresponding entities in $\mathcal{G}$, i.e., $\mathcal{T}_{q},\mathcal{A}_{q}\subseteq\mathcal{E}$.
-Appendix B Potential Impacts and Limitations
-The proposed method holds the potential to significantly enhance the performance of large language models (LLMs) by tackling the issue of hallucinations, thereby fostering deep, responsible reasoning.
-By integrating KGs with LLMs, the approach not only facilitates more accurate knowledge retrieval but also leverages deductive reasoning capabilities to steer the reasoning process and circumvent logical fallacies.
-The method is characterized by its stepwise, generalizable approach, and the use of deductive verification as a stopping criterion, which together may reduce superfluous computation and curb misleading reasoning chains.
-Moreover, due to its training-free nature and lower computational demands, this method could seamlessly serve as a plug-in for other existing frameworks.
-However, it’s important to recognize certain limitations.
-Since the method is still in its early stages of development, it might encounter unanticipated challenges or drawbacks in real-world applications.
-Its dependency on external KGs means that the quality and comprehensiveness of these resources can affect its overall effectiveness.
-Computational constraints might also arise when dealing with very large or intricate graphs.
-Furthermore, although the method has demonstrated promise in benchmark tests, its performance across more diverse or specialized tasks remains untested.
-Appendix C Bottleneck of Beam Search Efficiency
+Limitations
+Our work demonstrates a promising advancement by integrating KGs with LLMs to reduce hallucinations and promote deep, faithful reasoning through deductive verification.
+However, the method exhibits certain limitations.
+Its reliance on external KGs means that the overall effectiveness is contingent on the quality and comprehensiveness of these resources, and challenges may arise when encountering incomplete, inconsistent or outdated information.
+Despite these limitations, the open-KGs like Wikidata and DBpedia used in our study are of high quality, benefiting from years of updates by an extensive community.
+For domain-specific KGs, although there may currently be gaps in quality, we are optimistic about future enhancements.
+Given the significant societal impact and the noticeable boosts in LLM performance facilitated by KGs, it is likely that community efforts will continue to refine and expand these resources.
+Appendix A Experiment Details
+A.1 Baselines
+We consider the following methods including training-free (highlighted with *) and training-based methods as baselines:
+  • NSM He et al.
+(2021) propose a teacher-student approach for KGQA task, where the student network aims to find the correct answer to the query, while the teacher network tries to learn intermediate supervision signals for improving the reasoning capacity of the student network.
+  • KD-CoT Wang et al. (2023) propose to verify the sub-reasoning process of LLMs through the external KGs to facilitate faithful reasoning.
+  • DeCAF Yu et al.
+(2023) use a text-based retrieval instead of entity linking to select question-related information from the KG, and generate logical forms and direct answers respectively.
+They combine the logical-form-executed answers and directly-generated answers to obtain the final output.
+  • KAPING Baek et al.
+(2023) proposes a zero-shot knowledge-augmented prompting method.
+It first retrieves triples related to the question from the graph, then prepends them to the input question in the form of a prompt, which is then forwarded to LLMs to generate the answer.
+  • ToG Sun et al.
+(2023): conduct the reasoning on KGs by iteratively exploring multiple potential reasoning paths and concludes the final answer by aggregating the evidence from retrieved reasoning paths.
+  • RoG Luo et al. (2024a): incorporate the underling structure of KGs into LLMs throught pre-training and fine-tuning to generate the reasoning path and explanation.
+  • GCR Luo et al. (2024b) propose to integrate KG structure into the LLM decoding process to conduct graph-constrained reasoning.
+A.2 Datasets & Metrics
+We consider three KGQA benchmark: WebQuestionSP (WebQSP) Yih et al.
+(2016), Complex WebQuestions (CWQ) Talmor and Berant (2018) and CR-LT-KGQA Guo et al.
+(2024) in this work.
+We follow previous work Luo et al.
+(2024a) to use the same training and testing splits for fair comparison over WebQSP and CWQ.
+The questions from both WebQSP and CWQ can be reasoned using Freebase KGs [^1] https://github.com/microsoft/FastRDFStore.
+To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we further test our method on CR-LT-KGQA (discussed in §A.2).
+We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples.
+Each of the question can be reasoned based on the Wikidata [^2] https://www.wikidata.org/wiki/Wikidata:Main_Page.
+The statistics of the datasets are given in Table 10 and Table 9.
+To streamline the KGs, we follow RoG Luo et al.
+(2024a) and utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ.
+Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA.
+We assess the performance of the methods by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
+The statistics of the datasets can be found in Table 9 and Table 10.
+Dataset 1 hop 2 hop $\geq$ 3 hop
+WebQSP 65.49 % 34.51% 0.00%
+CWQ 40.91 % 38.34% 20.75%
+CR-LT 5.31 % 43.22% 51.57%
+Table 9: Statistics of the question hops in WebQSP, CWQ and CR-LT-KGQA.
+Dataset #Ans = 1 2 $\geq$ #Ans $\leq$ 4 5 $\geq$ #Ans $\leq$ 9 #Ans $\geq$ 10
+WebQSP 51.2% 27.4% 8.3% 12.1%
+CWQ 70.6% 19.4% 6% 4%
+Table 10: Statistics of the number of answers for questions in WebQSP and CWQ.
+Motivation of CR-LT-KGQA.
+The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities.
+These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs.
+Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering.
+To this end, we utilize the CR-LT-KGQA benchmark, which features queries specifically crafted to target obscure and long-tail entities.
+Figure 4 illustrates the distribution of entity frequency and popularity in CR-LT, underscoring the inherent challenges of these queries.
+In such scenarios, knowledge graphs are indispensable as they offer a reliable, verifiable source of information, particularly for entities that are poorly represented in the training data of large language models.
+By testing our methods on CR-LT-KGQA, we investigate the extent to which integrating KGs can bolster LLM performance in less common knowledge domains, where their effectiveness typically declines.
+This evaluation not only demonstrates the potential synergy between LLMs and KGs but also clarifies the critical role that KGs continue to play in supporting LLMs across diverse query scenarios.
+Figure 4: Distribution of CR-LT-KGQA dataset.
+Figure 5: Parameter tuning for $\alpha$ for scoring function over WebQSP
+Backend Models WebQSP CWQ CR-LT
+Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
+Llama-2-13B 72.34 69.78 58.41 54.78 60.87
+Mistral-7B 74.11 70.23 60.71 56.87 63.12
+Table 11: Performance over Open-sourced LLMs.
+A.3 Backbone LLMs & Embedding Methods
+Backbone LLMs.
+We assess our approach on closed- and open-source LLMs.
+We consider closed-source models like GPT-4-turbo (between Feb, 2024 to July, 2024), GPT-3.5-turbo (between Feb, 2024 to July, 2024), GPT-4o, GPT-4o-mini (between Nov, 2024 to Jan 2025) from OpenAI, and open-sourced models like meta-llama-2-13B from Meta and mixtral-7B from Mixtral AI.
+The experiment results of open-source models can be found in Table 11.
+We set all the inference configs using temperature $T=0.3$ and $p=1.0$.
+Embedding Methods.
+We assess the robustness of the retrieval module Path-RAG on different embedding models.
+We consider probabilistic ranking function like BM25333https://en.wikipedia.org/wiki/Okapi_BM25, dense retrieval using smaller language models like SentenceBERT [^4] https://sbert.net and E5555https://huggingface.co/intfloat/e5-large, and more advanced embedding model like text-embedding-3-small from OpenAI [^6] https://platform.openai.com/docs/guides/embeddings.
+A.4 Implementation Details
+We set the default beam width as $4$ and depth as $4$ without specific annotation.
+We set the $\alpha$ in Eq 2 as $0.3$ to ensure reproducibility.
+For hyper-parameter tuning regarding $\alpha$ for Eq 2 and beam search width and length, we conduct experiments as shown in Figure 5 and Figure 6.
+Analysis of Beam Search.
+We investigate the effect of hyper-parameters like beam width and depth in beam search, as illustrated in Figure 6.
+By varying the width and depth from 1 to 4, we observe that overall performance improves as both parameters increase, peaking when the search depth exceeds 3 for the WebQSP and CWQ datasets.
+However, beyond a depth of 3, performance begins to decline, likely because only a small fraction of questions in these datasets require reasoning at greater depths.
+In contrast, increasing the beam width consistently enhances performance, highlighting the benefits of broader exploration in search.
+(a) RD over CWQ
+(b) RD over WebQSP
+(c) BW over CWQ
+(d) BW over WebQSP
+Figure 6: Analysis of various beam search width (BW) and reasoning depth (RD).
+A.5 Robustness Analysis Across Different Domains and KGs
+KGs vary in structure and domain-specific characteristics, so consistent performance across both general and specialized KGs can reflect a method’s adaptability to diverse real-world applications.
+To this end, we conduct robustness analysis of FiDeLiS across different domains and KGs to verify the generalizability.
+To perform this analysis, we introduced a new dataset, MedQA-USMIE, sourced from MedQA Jin et al.
+(2020), which is designed to require domain-specific biomedical and clinical knowledge.
+The dataset is a 4-way multiple-choice question-answering task, and we extracted 300 examples from its test set for evaluation.
+The corresponding biomedical KG is based on Disease Database and DrugBank Zhang et al.
+(2022).
+The results, presented in Table 2, indicate that our method exhibits consistent robustness across different types of KGs.
+Our scoring function, enhanced by incorporating next-hop neighbor information $S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}(S_{r}^{j}+S_{e}^{j})$, achieves higher performance gains in both WebQSP and MedQA-USMIE, particularly improving accuracy in the specialized biomedical domain.
+These findings validate that our method can effectively handle the challenges posed by both general and domain-specific knowledge graphs, indicating strong adaptability and robustness.
+Method WebQSP MedQA-USMIE
+ToG 81.84 42.37
+[…]
+Path-RAG w/ $S_{r}^{i}+S_{e}^{i}$ 83.15 44.31
+Path-RAG w/ $S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}(S_{r}^{j}+S_{e}^{j})$ 84.39 46.45
+Table 12: Robustness analysis of our method across different domains
+Appendix B Bottleneck of Beam Search Efficiency
 The bottleneck of computation is the beam search process, which contributes to $N*D$ times LLM calling, where $D$ is the depth (or equivalently length) of the reasoning path, and $N$ is the width of the beam-search (how many paths are remained in the pool in each iteration).
@@
 Sun et al.
-(2023) illustrates that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling.
+(2023) indicate that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling.
 In this way, we can reduce the number of LLM calling from $ND+D+C$ to $D+C$.
-However, Sun et al.
-(2023) shows that this optimization sacrifices the accuracy due to the weaker scoring model in decision making.
+However, this may sacrifices the accuracy due to the weaker scoring model in decision making Sun et al.
+(2023).
 We noted that $ND+D+C$ is the maximal computational complexity.
@@
 We can also see that this average number gets even smaller for dataset covering a lot of single-hop reasoning questions, such as WebQSP.
-Appendix D Experiment Details
-D.1 Baselines
-Baselines.
-(1) RoG (Luo et al., 2024): embed structure knowledge graph from KGs into the underlying neural networks during the pretraining and fine-tuning process to generate the reasoning path and explanation.
-(2) ToG (Sun et al., 2023): ask LLM to iteratively explore multiple possible reasoning paths on KGs until the LLM determines that the question can be answered based on the current reasoning paths.
-Our method FiDeLiS follows the similar paradigm.
-(3) NSM (He et al., 2021) utilizes the sequential model to mimic the multi-hop reasoning process.
-(4) KD-CoT (Wang et al., 2023a) retrieves relevant knowledge from KGs to generate faithful reasoning paths for LLMs.
-(5) DeCAF (Yu et al., 2022) combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance on KGQA tasks.
-Implementation Details.
-We set the default beam width as $4$ and depth as $4$ without specific annotation.
-We set the $\alpha$ in Eq 3 as $0.3$ to ensure reproducibility.
-For LLMs usage, we set all the inference using temperature $T=0.3$ and $p=1.0$.
-The hyperparameter tuning experiments for beam search and $\alpha$ can be found in Figure 2 and Table 9.
-Backboned LLMs.
-We assess our approach on closed- and open-source LLMs.
-For closed-source LLMs, we choose GPT-4-turbo, GPT-3.5-turbo [^1] We use the recent gpt-4-turbo model that is released in 2024-04-09 from https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 and gpt-3.5-turbo model notated as “gpt-3.5-turbo-0125” from https://platform.openai.com/docs/models/gpt-3-5-turbo.
-to report and compare the results on all datasets.
-We use Llama-2-13B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023a) as our open-source LLMs to conduct cost–performance analysis on different datasets.
-The experiment results of open-source models can be found in Table 10.
-Table 7: Statistics of the number of answers for questions in WebQSP and CWQ.
-Dataset #Ans = 1 2 $\geq$ #Ans $\leq$ 4 5 $\geq$ #Ans $\leq$ 9 #Ans $\geq$ 10
-WebQSP 51.2% 27.4% 8.3% 12.1%
-CWQ 70.6% 19.4% 6% 4%
-D.2 Datasets
-Datasets & Metrics.
-We adopt three benchmark KGQA datasets: WebQuestionSP (WebQSP) (Yih et al., 2016), Complex WebQuestions (CWQ) (Talmor & Berant, 2018) and CR-LT-KGQA (Guo et al., 2024) in this work.
-We follow previous work (Luo et al., 2024) to use the same training and testing splits for fair comparison over WebQSP and CWQ.
-The questins from both WebQSP and CWQ can be reasoned using Freebase KGs [^2] https://github.com/microsoft/FastRDFStore.
-To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we further test our method on CR-LT-KGQA (further discussed in Appendix Section 8).
-We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples.
-Each of the question can be reasoned based on the Wikidata [^3] https://www.wikidata.org/wiki/Wikidata:Main_Page.
-The statistics of the datasets are given in Table 7 and Table 8.
-To streamline the KGs, we utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ followsing RoG (Luo et al., 2024).
-Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA as well.
-We assess the performance of the methods by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
-Table 8: Statistics of the question hops in WebQSP, CWQ and CR-LT-KGQA.
-Dataset 1 hop 2 hop $\geq$ 3 hop
-WebQSP 65.49 % 34.51% 0.00%
-CWQ 40.91 % 38.34% 20.75%
-CR-LT 5.31 % 43.22% 51.57%
-Motivation of CR-LT-KGQA.
-The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities.
-These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs.
-Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering.
-Figure 4: Distribution of CR-LT-KGQA dataset.
-Against this backdrop, we specifically chose CR-LT-KGQA for our evaluation as it includes queries related to obscure or long-tail entities, the distribution of the frequency and entity popularity of CR-LT is shown as Figure 4.
-Such scenarios are where KGs play a critical role because they provide a reliable source of verifiable information, crucial when LLMs encounter entities not well covered in their training data.
-By testing our methods on CR-LT-KGQA, we aim to explore how effective LLMs can operate when combined with KGs, especially in contexts involving less common knowledge domains where LLM performance typically declines.
-This evaluation helps us understand the extent to which KGs remain necessary for supporting LLMs in a diverse range of query scenarios.
-D.3 Parameter Tuning for $\alpha$ for Scoring Function
-For hyperparameter tuning regarding $\alpha$ for Eq 3, we added an extra comparison in Table 9 which shows that $\alpha$ is actually not a very impact parameter for the whole system, however for the reproducibility of our method, we set the $\alpha$ as 0.3.
-Table 9: Parameter tuning for $\alpha$ for scoring function over WebQSP
-$\alpha$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
-Hits@1 83.76 84.15 84.39 83.98 83.55 84.12 83.79 82.25 83.60 83.73
-Appendix E Algorithm Description
+Appendix C Prompt List
+In this section, we show all the prompts that need to be used in the main experiments. The In-Context Few-shot refers to the few-shot examples we used for in-context learning.
+C.1 Plan-and-solve
+You are a helpful assistant designed to output JSON that aids in navigating a knowledge graph to answer a provided question. The response should include the following keys:
+(1) ’keywords’: an exhaustive list of keywords or relation names that you would use to find the reasoning path from the knowledge graph to answer the question.
+Aim for maximum coverage to ensure no potential reasoning paths will be overlooked;
+(2) ’planning_steps’: a list of detailed steps required to trace the reasoning path with. Each step should be a string instead of a dict.
+(3) ’declarative_statement’: a string of declarative statement that can be transformed from the given query, For example, convert the question ’What do Jamaican people speak?’ into the statement ’Jamaican people speak *placeholder*.’ leave the *placeholder* unchanged; Ensure the JSON object clearly separates these components.
+In-Context Few-shot
+Q: {Query}
+A:
+C.2 Deductive-verification
+You are asked to verify whether the reasoning step follows deductively from the question and the current reasoning path in a deductive manner. If yes return yes, if no, return no".
+In-Context Few-shot
+Whether the conclusion ’{declarative_statement}’ can be deduced from ’{parsed_reasoning_path}’, if yes, return yes, if no, return no.
+A:
+C.3 Adequacy-verification
+You are asked to verify whether it’s sufficient for you to answer the question with the following reasoning path.
+For each reasoning path, respond with ’Yes’ if it is sufficient, and ’No’ if it is not.
+Your response should be either ’Yes’ or ’No’.
+In-Context Few-shot
+Whether the reasoning path ’{reasoning_path}’ be sufficient to answer the query ‘{Query}’, if yes, return yes, if no, return no.
+A:
+C.4 Beam Search
+Given a question and the starting entity from a knowledge graph, you are asked to retrieve reasoning paths from the given reasoning paths that are useful for answering the question.
+In-Context Few-shot
+Considering the planning context {plan_context} and the given question {Query}, you are asked to choose the best {beam_width} reasoning paths from the following candidates with the highest probability to lead to a useful reasoning path for answering the question.
+{reasoning_paths}.
+Only return the index of the {beam_width} selected reasoning paths in a list.
+A:
+C.5 Reasoning
+Given a question and the associated retrieved reasoning path from a knowledge graph, you are asked to answer the following question based on the reasoning path and your knowledge.
+Only return the answer to the question.
+In-Context Few-shot
+Question: {Query}
+Reasoning path: {reasoning_path}
+Only return the answer to the question.
+A:
+C.6 Demonstration of Deductive Verification
+Deductive Verification Example Question: Who is the ex-wife of Justin Bieber’s father?
+After one round of beam searching, the current reasoning path is: Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber.
+The next step candidates are: 1.
+people.married_to.person $\to$ Erin Wagner 2.
+people.person.place_of_birth $\to$ US, $\ldots$ The deductive reasoning can be formulated as follows: Premises: - Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber (from the current reasoning path) - Jeremy_bieber $\to$ people.married_to.person $\to$ Erin Wagner (from the next step candidates) Conclusion: Erin Wagner is the ex-wife of Justin Bieber’s father.
+(Using a large language model (LLM) zero-shot approach to reformat the question into a cloze filling task, we use the last entity from the next step candidates, "Erin Wagner", to fill the cloze.) The prompt will ask whether the conclusion can be deduced from the given premises.
+If the answer is "yes", return “yes”, otherwise return “no.”
 Algorithm 1 Path-RAG Initialization and Retrieval Process
@@
 21:return the best hypothesis from $\mathcal{H}_{T}$.
-Appendix F Prompt List
-In this section, we show all the prompts that need to be used in the main experiments. The In-Context Few-shot refers to the few-shot examples we used for in-context learning.
-F.1 Plan-and-solve
-You are a helpful assistant designed to output JSON that aids in navigating a knowledge graph to answer a provided question. The response should include the following keys:
-(1) ’keywords’: an exhaustive list of keywords or relation names that you would use to find the reasoning path from the knowledge graph to answer the question.
-Aim for maximum coverage to ensure no potential reasoning paths will be overlooked;
-(2) ’planning_steps’: a list of detailed steps required to trace the reasoning path with. Each step should be a string instead of a dict.
-(3) ’declarative_statement’: a string of declarative statement that can be transformed from the given query, For example, convert the question ’What do Jamaican people speak?’ into the statement ’Jamaican people speak *placeholder*.’ leave the *placeholder* unchanged; Ensure the JSON object clearly separates these components.
-In-Context Few-shot
-Q: {Query}
-A:
-F.2 Deductive-verification
-You are asked to verify whether the reasoning step follows deductively from the question and the current reasoning path in a deductive manner. If yes return yes, if no, return no".
-In-Context Few-shot
-Whether the conclusion ’{declarative_statement}’ can be deduced from ’{parsed_reasoning_path}’, if yes, return yes, if no, return no.
-A:
-F.3 Adequacy-verification
-You are asked to verify whether it’s sufficient for you to answer the question with the following reasoning path.
-For each reasoning path, respond with ’Yes’ if it is sufficient, and ’No’ if it is not.
-Your response should be either ’Yes’ or ’No’.
-In-Context Few-shot
-Whether the reasoning path ’{reasoning_path}’ be sufficient to answer the query ‘{Query}’, if yes, return yes, if no, return no.
-A:
-F.4 Beam Search
-Given a question and the starting entity from a knowledge graph, you are asked to retrieve reasoning paths from the given reasoning paths that are useful for answering the question.
-In-Context Few-shot
-Considering the planning context {plan_context} and the given question {Query}, you are asked to choose the best {beam_width} reasoning paths from the following candidates with the highest probability to lead to a useful reasoning path for answering the question.
-{reasoning_paths}.
-Only return the index of the {beam_width} selected reasoning paths in a list.
-A:
-F.5 Reasoning
-Given a question and the associated retrieved reasoning path from a knowledge graph, you are asked to answer the following question based on the reasoning path and your knowledge.
-Only return the answer to the question.
-In-Context Few-shot
-Question: {Query}
-Reasoning path: {reasoning_path}
-Only return the answer to the question.
-A:
-F.6 Demonstration of Deductive Verification
-Deductive Verification Example Question: Who is the ex-wife of Justin Bieber’s father?
-After one round of beam searching, the current reasoning path is: Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber.
-The next step candidates are: 1.
-people.married_to.person $\to$ Erin Wagner 2.
-people.person.place_of_birth $\to$ US, $\ldots$ The deductive reasoning can be formulated as follows: Premises: - Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber (from the current reasoning path) - Jeremy_bieber $\to$ people.married_to.person $\to$ Erin Wagner (from the next step candidates) Conclusion: Erin Wagner is the ex-wife of Justin Bieber’s father.
-(Using a large language model (LLM) zero-shot approach to reformat the question into a cloze filling task, we use the last entity from the next step candidates, "Erin Wagner", to fill the cloze.) The prompt will ask whether the conclusion can be deduced from the given premises.
-If the answer is "yes", return “yes”, otherwise return “no.”
-Table 10: Comparison of FiDeLiS using different backbone LLMs.
-Backend Models Methods WebQSP CWQ CR-LT
-Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
-Llama-2-13B FiDeLiS 72.34 69.78 58.41 54.78 60.87
-Mistral-7B FiDeLiS 74.11 70.23 60.71 56.87 63.12
-Appendix G Related Works: Reasoning with LLM Prompting
-With the development of LLMs, many creative ways to leverage LLMs are proposed: End-to-End, Chain-of-Thought, and Semantic Parsing/Code Generation.
-End-to-End methods (Chen, 2022; Hoffmann et al., 2022) aims to leverage LLMs to generate final answers directly, often done by providing a task description and/or a few examples for in-context learning.
-While it offers convenience, however, it suffers from the un-interpretability of the generation, and the lack of explicit steps and reliance solely on the LLMs’ training data may result in un-robustness and exhibit sensitivity to slight input variations.
-Chain-of-Thought methods (Wei et al., 2022a; Kojima et al., 2022; Gong et al., 2020) emphasize breaking down complex reasoning into a series of intermediate steps to support complex reasoning.
-However, CoT suffers from unreliability and uncontrollability since the generated reasoning steps may not always align with the intended logical progression or may produce incorrect or inconsistent answers when multiple questions are posed consecutively.
-Plan-and-solve methods (Wang et al., 2023b) prompts LLMs to generate a plan and conduct reasoning based on it.
-DecomP He et al.
-(2021) prompts LLMs to decompose the reasoning task into a series of sub-tasks and solve them step by step.
-Semantic Parsing/Code Generation methods (Cheng et al., 2023; Ye et al., 2023; Gemmell & Dalton, 2023) leverages LLMs to convert natural language queries into executable code or structured representations.
-This approach enables more precise control over the output and facilitates better interpretability.
-However, it has its limitations, such as the limited coverage of programming language grammar and semantics.
-This restricts the model’s ability to handle complex programming tasks and may require additional techniques to handle out-of-domain queries effectively.
-Cheng et al.
-(2023) propose to leverage LLMs to bridge the out-of-domain queries, however, the form of knowledge injection still lacks a thorough exploration and the trigger functions are still naive for complex queries.
-However, the problem of hallucinations and lack of knowledge affect the faithfulness of LLMs’ reasoning.
-ReACT Yao et al.
-(2022) treats LLMs as agents, which interact with the environment to get the latest knowledge for reasoning.
-To explore faithful reasoning, FAME Hong et al.
-(2023) introduces the Monte-Carlo planning to generate faithful reasoning steps.
-RR He et al.
-(2022) and KD-CoT Wang et al.
-(2023a) further retrieve relevant knowledge from KGs to produce faithful reasoning plans for LLMs.

### v4

@@
 However, existing KG-enhanced methods, either retrieval-based or agent-based, encounter difficulties in accurately retrieving knowledge and efficiently traversing KGs at scale.
-In this paper, we propose a unified framework, FiDeLiS, designed to improve the factuality of LLM responses by anchoring answers to verifiable reasoning steps retrieved from a KG.
-To achieve this, we leverage step-wise beam search with a deductive scoring function, allowing the LLM to validate each reasoning step and halt the search once the question is deducible.
-In addition, our Path-rag module pre-selects a smaller candidate set for each beam search step, reducing computational costs by narrowing the search space.
-Extensive experiments show that our training-free and efficient approach outperforms strong baselines, enhancing both factuality and interpretability.
-Code is released at https://anonymous.4open.science/r/FiDELIS-E7FC.
+In this paper, we propose a unified framework, FiDeLiS [^1] Code is released at https://github.com/Y-Sui/FiDeLiS., designed to improve the factuality of LLM responses by anchoring answers to verifiable reasoning steps retrieved from KGs.
+To achieve this, we leverage step-wise beam search with a deductive scoring function, allowing the LLM to validate reasoning process step by step, and halt the search once the question is deducible.
+In addition, we propose a Path-RAG module to pre-select a smaller candidate set for each beam search step, reducing computational costs by narrowing the search space.
+Extensive experiments show that our method, as a training-free framework, not only improve the performance but also enhance the factuality and interpretability across different benchmarks.
 FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering
@@
 Large Language Models (LLMs) have shown impressive reasoning capabilities in tackling complex tasks Yu et al.
+(2024); Sui et al.
 (2025).
-However, the reasoning of LLMs is not always reliable and can be prone to generating outputs that are either inconsistent with real-world facts Xu et al.
+However, their reasoning processes are not always reliable, and can be prone to generating outputs that are either inconsistent with real-world facts Xu et al.
 (2024); Huang et al.
@@
 (2024); Sui et al.
-(2024), which greatly undermine the reliability of LLMs in real-world applications Kung et al.
+(2024).
+Such limitations significantly undermine the trustworthiness of LLMs in practical, real-world applications Kung et al.
 (2023); Zhang et al.
@@
 (2024); Luo et al.
-(2024a).
+(2024).
 Unlike traditional retrieval-augmented generation (RAG) that relies on web pages or documents Liu et al.
@@
 Existing KG-enhanced LLM reasoning methods face notable challenges and can be roughly categorized into two primary approaches: retrieval-based and agent-based paradigms Luo et al.
-(2024b).
+(2025).
 Retrieval-based methods Wang et al.
 (2023); Luo et al.
-(2024a); Baek et al.
+(2024); Baek et al.
 (2023) retrieve relevant KG facts to support LLM reasoning by either prompting Baek et al.
 (2023) or fine-tuning LLMs to learn the underlying structure of KG Luo et al.
-(2024a, b).
+(2024, 2025).
 These methods often suffer from incomplete or imprecise information extraction due to a lack of contextual understanding or an inability to fully capture the graph structure Luo et al.
-(2024b).Our error analysis of a strong retrieval-based method (i.e., Luo et al.
-(2024a)) in §4.3 reveals that only 67% of the generated reasoning steps are valid, with 33% containing format errors or referencing non-existent KG facts.
+(2025).
+Our error analysis (§4.3) of a strong retrieval-based method, RoG Luo et al.
+(2024) reveals that only 67% of the generated reasoning steps are valid, with 33% containing format errors or referencing non-existent KG facts.
 In contrast, agent-based methods Sun et al.
@@
 While this approach by nature can enhance reasoning accuracy, it is computationally expensive, resulting in high latency and scalability limitations.
-As illustrated in Figure 1, balancing faithfulness and efficiency remains a critical challenge for existing KG-enhanced reasoning methods.
-To this end, we propose FiDeLiS, a unified framework designed to improve the factual accuracy and reasoning efficiency of LLMs on KGQA task.
-FiDeLiS anchors LLM responses to verifiable reasoning steps derived from a KG by employing two core components: (1) Deductive-Verification Beam Search (DVBS) which systematically constructs and validates reasoning paths step-by-step, ensuring logical consistency and factual correctness (discussed in §3.2).
-This module also prevents premature reasoning termination and incorrect path extension to ensure the validity of the generated reasoning paths.
-(2) Path-RAG, a retrieval-augmented mechanism that pre-selects a constrained set of candidate entities and relations for each step to mitigate computational inefficiencies.
-It combines semantic similarity measures with graph-based connectivity analysis to optimize the search space, significantly reducing latency without sacrificing recall or accuracy (discussed in §3.1).
-Extensive experiments show that our method outperform strong baselines in both accuracy and efficiency, offering a scalable, training-free solution for KG-enhanced LLM reasoning.
-Overall, our main contributions include:
-  • We propose FiDeLiS, a unified framework designed to improve the factual accuracy of LLMs by grounding reasoning paths in structured KG efficiently.
-  • We enable efficient, verifiable reasoning by deductively validating reasoning steps and narrowing the search space with high-quality retrieval mechanism.
-  • FiDeLiS performs robustly across different experiments without the need for model fine-tuning, demonstrating adaptability and scalability with improved performance on multiple benchmarks.
-  • By anchoring responses in verifiable reasoning paths, FiDeLiS enhances interpretability, enabling users to verify and understand each reasoning step.
+As illustrated in Figure 1, how to balance the trade-off between faithfulness and efficiency remains a critical challenge for existing KG-enhanced reasoning methods.
+To this end, we propose FiDeLiS, a unified framework designed to enhance both the factual accuracy and reasoning efficiency of LLMs.
+FiDeLiS grounds LLM responses in verifiable reasoning steps derived from a KG through two core components: (1) Deductive-Verification Beam Search (DVBS), which systematically constructs and validates reasoning paths step-by-step to ensure logical consistency and factual correctness (detailed in §3.2).
+This module also prevents premature termination and incorrect path extensions, thereby guaranteeing the validity of generated reasoning chains.
+(2) Path-RAG, a retrieval-augmented mechanism that pre-selects a constrained set of candidate entities and relations at each reasoning step to mitigate computational inefficiencies.
+By combining semantic similarity with graph-based connectivity analysis, Path-RAG effectively narrows the search space, significantly reducing latency without compromising recall or accuracy (discussed in §3.1).
+Extensive experiments demonstrate that FiDeLiS outperforms strong baselines in both accuracy and efficiency, providing a scalable, training-free solution for KG-enhanced LLM reasoning.
+Our main contributions are summarized as follows:
+  • We propose FiDeLiS, a unified framework that efficiently grounds LLM reasoning in structured knowledge graphs to improve factual accuracy.
+  • We enable verifiable and efficient reasoning by combining deductive validation of reasoning steps with a high-quality retrieval mechanism that constrains the search space.
+  • We demonstrate FiDeLiS’s robustness and scalability across multiple benchmarks without requiring any model fine-tuning.
+  • By anchoring responses in verifiable reasoning paths, FiDeLiS enhances interpretability and allows users to verify and understand each reasoning step from LLM reasoning.
 2 Preliminary
@@
 […]
+Figure 2: An illustration of FiDeLiS.
+Top: The workflow of Path-RAG.
+An LLM first extracts key terms and generates dense embeddings that feed into the Path-RAG module.
+Then, Path-RAG rapidly retrieves relevant entities and relations from a pre-embedded KG and constructs candidate reasoning steps by combining semantic similarity with graph connectivity.
+Bottom: The workflow of DVBS.
+Next, the DVBS module uses LLM-generated planning to guide a beam search that builds reasoning paths step-by-step over candidates constructed by Path-RAG, with deductive verification ensuring each step logically follows the previous steps and support the user question.
 Task definition.
@@
 $P(a|q,\mathcal{G})=\sum_{\mathcal{P}}P_{\theta}(a|q,\mathcal{P})\prod_{k=1}^{n}P_{\phi}(t_{k}|q,t_{<k},\mathcal{G})$ (1)
-To acquire valid reasoning paths, most prior studies follow the retrieval-based Li et al.
+To acquire valid reasoning paths, most prior studies follow either retrieval-based Li et al.
 (2023); Luo et al.
-(2024a) or agent-based Sun et al.
+(2024, 2025) or agent-based Sun et al.
+(2023); Ma et al.
 (2024) paradigm.
 As indicated in Luo et al.
-(2024b), retrieval-based methods rely on precise additional retrievers, while agent-based methods are computationally intensive and lead to high latency.
-To address these issues, we propose our method, FiDeLiS, to enable both efficient and faithful reasoning over KGs.
-Figure 2: An illustration of FiDeLiS.
-Top: The workflow of Path-RAG.
-An LLM first extracts key terms and generates dense embeddings that feed into the Path-RAG module.
-Path-RAG rapidly retrieves relevant entities and relations from a pre-embedded KG and constructs candidate reasoning steps by combining semantic similarity with graph connectivity.
-Bottom: The workflow of DVBS.
-Then, the DVBS module uses an LLM-generated planning outline to guide a beam search that builds reasoning paths step-by-step over candidates constructed by Path-RAG, with deductive verification ensuring each step logically follows the previous steps and support the user question.
+(2025), retrieval-based methods rely on additional precise retrievers, while agent-based methods are computationally intensive and may lead to high latency.
+To balance this trade-off, we propose our method, FiDeLiS, to enable both efficient and faithful reasoning over KGs.
 3 Method
@@
 The overall framework of FiDeLiS is illustrated in Figure 2, which consists of two main components: (1) Reasoning Path Retrieval-Augmented Generation (Path-RAG, Algorithm 1) and (1) Deductive-verification Beam Search (DVBS, Algorithm 2).
-Given a complex question $q$, we first use an LLM to extract key terms from $q$ and generate dense embeddings that capture the question’s core concepts.
-These embeddings are input into Path-RAG module, which rapidly retrieves relevant entities and relations from a pre-embedded KG to select a smaller candidate sets for further beam search step, addressing the latency and computational burden of traditional agent-based methods.
+Given a complex question $q$, we first employ an LLM to extract key terms and generate dense embeddings that capture the question’s core concepts.
+These embeddings serve as input to the Path-RAG module, which efficiently retrieves relevant entities and relations from a pre-embedded KG, narrowing down the candidate set for subsequent beam search.
+This approach addresses the latency and computational overhead typical of traditional agent-based methods.
 Path-RAG then constructs candidate reasoning steps by combining immediate semantic similarity with the structural connectivity of the graph, overcoming the dependence on highly precise retrievers in standard retrieval-based approaches.
-Next, the DVBS module employs an LLM-generated planning outline to guide a beam search that builds reasoning paths step-by-step.
-At each step, deductive verification checks that the accumulated reasoning steps logically supports the user question, ensuring the final reasoning path is both verifiable and accurate.
+Next, the DVBS module employs an LLM-generated planning to guide the beam search that builds reasoning paths step-by-step.
+At each step, deductive verification checks whether the accumulated reasoning steps logically supports the user question, ensuring the final reasoning path is both verifiable and accurate.
 3.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
@@
 (2023) treat LLMs as agents that iteratively interact with KGs to find reasoning paths and answer, which necessitate multiple rounds of interaction between agents and KGs and lead to high computational costs and latency.
-We instead propose a module, Path-RAG which iteratively pre-select a smaller candidate set to reduces the search space for exploring the potential reasoning paths from KGs.
+We instead propose a module, Path-RAG, which iteratively pre-select a smaller candidate set to reduces the search space for exploring the potential reasoning paths from KGs.
 It consists of three steps and we detail the workflow as follows:
@@
 KD-CoT Wang et al. (2023) 73.7 50.2 50.5 - -
-RoG Luo et al. (2024a) 83.15 69.81 61.39 56.17 60.32
+RoG Luo et al. (2024) 83.15 69.81 61.39 56.17 60.32
 […]
@@
 To verify the faithfulness of our step-wise method, we conduct an error analysis regarding the whole reasoning path generation using RoG Luo et al.
-(2024a).
+(2024).
 We quantify the validity of reasoning path using validity ratio (VR), which is defined as the ratio of reasoning steps that existed in the KG to the total number of the reasoning steps in the output reasoning path (i.e., $\mathrm{VR}=\frac{N_{\mathrm{valid-steps}}}{N_{\text{all-steps}}}$).
@@
 While our method leverage step-wise verification to ensure that each of the reasoning step exist in the KG and logically connected.
-Figure 3: Analysis of reasoning errors in RoG Luo et al. (2024a) over WebQSP.
+Figure 3: Analysis of reasoning errors in RoG Luo et al. (2024) over WebQSP.
 Method WebQSP CWQ CR-LT
@@
 (2025); Li et al.
+(2024); Chen et al.
+(2025); He et al.
 (2025).
@@
 To address these limitations, recent methods like RoG Luo et al.
-(2024a) and GCR Luo et al.
-(2024b) have sought to integrate KG structure into LLM training or decoding to improve reasoning fidelity and explanation generation.
+(2024) and GCR Luo et al.
+(2025) have sought to integrate KG structure into LLM training or decoding to improve reasoning fidelity and explanation generation.
 To improve the faithfulness of the LLM reasoning, KD-CoT Wang et al.
@@
 (2023): conduct the reasoning on KGs by iteratively exploring multiple potential reasoning paths and concludes the final answer by aggregating the evidence from retrieved reasoning paths.
-  • RoG Luo et al. (2024a): incorporate the underling structure of KGs into LLMs throught pre-training and fine-tuning to generate the reasoning path and explanation.
-  • GCR Luo et al. (2024b) propose to integrate KG structure into the LLM decoding process to conduct graph-constrained reasoning.
+  • RoG Luo et al. (2024): incorporate the underling structure of KGs into LLMs throught pre-training and fine-tuning to generate the reasoning path and explanation.
+  • GCR Luo et al. (2025) propose to integrate KG structure into the LLM decoding process to conduct graph-constrained reasoning.
 A.2 Datasets & Metrics
@@
 We follow previous work Luo et al.
-(2024a) to use the same training and testing splits for fair comparison over WebQSP and CWQ.
+(2024) to use the same training and testing splits for fair comparison over WebQSP and CWQ.
 The questions from both WebQSP and CWQ can be reasoned using Freebase KGs [^2] https://github.com/microsoft/FastRDFStore.
@@
 To streamline the KGs, we follow RoG Luo et al.
-(2024a) and utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ.
+(2024) and utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ.
 Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA.
@@
 We assess the robustness of the retrieval module Path-RAG on different embedding models.
-We consider probabilistic ranking function like BM25333https://en.wikipedia.org/wiki/Okapi_BM25, dense retrieval using smaller language models like SentenceBERT [^4] https://sbert.net and E5555https://huggingface.co/intfloat/e5-large, and more advanced embedding model like text-embedding-3-small from OpenAI [^6] https://platform.openai.com/docs/guides/embeddings.
+We consider probabilistic ranking function like BM25444https://en.wikipedia.org/wiki/Okapi_BM25, dense retrieval using smaller language models like SentenceBERT [^5] https://sbert.net and E5666https://huggingface.co/intfloat/e5-large, and more advanced embedding model like text-embedding-3-small from OpenAI [^7] https://platform.openai.com/docs/guides/embeddings.
 A.4 Implementation Details